### v1

# Graph-Theoretic Analysis of $n$-Replica Time Evolution in the Brownian Gaussian Unitary Ensemble

Abstract
In this paper, we investigate the $n$-replica time evolution operator $\mathcal{U}_{n}(t)\equiv e^{\mathcal{L}_{n}t}$ for the Brownian Gaussian Unitary Ensemble (BGUE) using a graph-theoretic approach. We examine the moments of the generating operator $\mathcal{L}_{n}$, which governs the Euclidean time evolution within an auxiliary $D^{2n}$-dimensional Hilbert space, where $D$ represents the dimension of the Hilbert space for the original system. Explicit representations for the cases of $n=2$ and $n=3$ are derived, emphasizing the role of graph categorization in simplifying calculations. Furthermore, we present a general approach to streamline the calculation of time evolution for arbitrary $n$, supported by a detailed example of $n=4$. Our results demonstrate that the $n$-replica framework not only facilitates the evaluation of various observables but also provides valuable insights into the relationship between Brownian disordered systems and quantum information theory.
preprint: USTC-ICTS/PCFT-25-05
1 Introduction
Brownian disordered quantum systems have garnered significant interest across various fields, including many-body quantum chaos Chen_2019 ; S_nderhauf_2019 ; kalsi2024hierarchicalanalyticalapproachuniversal , quantum gravity Brady_2013 ; pelliconi2024gravitymesoscopic , and quantum information theory Brady_2013 ; Bentsen_2021 ; liu2023signaturescrambloneffectivefield ; Zhang_2023 ; Tang:2024kpv . Unlike systems with static (quenched) disorder, Brownian disorder is characterized by time-dependent randomness, allowing ensemble averaging to yield analytically tractable descriptions of non-equilibrium dynamics. A paradigmatic example is the Brownian Sachdev-Ye-Kitaev (SYK) model Saad:2018bqo , which is significantly trackable than the regular SYK Maldacena_2016 ; Maldacena:2016hyu . This has led to its widespread use as a toy model for quantum gravity Narovlansky:2023lfz ; Milekhin:2023bjv ; Stanford:2023npy ; Hayden_2007 , quantum error correction (QEC) Balasubramanian:2023xdp , out-of-time-ordered correlators (OTOCs) Sunderhauf:2019djv , and other fundamental aspects of quantum chaos Jian:2020krd ; Agarwal:2020yky ; Roberts_2017 ; Cotler_2017 ; jian2022lineargrowthcircuitcomplexity ; Peng:2022pfa ; fava2023designsfreeprobability .
Building on these developments, the recently proposed Brownian Gaussian Unitary Ensemble (BGUE) Tang:2024kpv offers an alternative setting to explore Brownian disorder in quantum systems. Unlike the standard Gaussian Unitary Ensemble (GUE), where the Hamiltonian is randomly drawn but remains fixed in time, BGUE describes a system evolving under a time-dependent noisy Hamiltonian,
$H_{ij}(t)=\eta_{ij}(t),\eta_{ij}(t)=\eta_{ij}^{\ast}(t)\in\mathbb{C},$ (1)
where $\eta_{ij}(t)$ are independent Brownian Gaussian random variables with vanishing mean and nonzero variance
$\mathbb{E}\left(\eta_{ij}(t)\right)=0,\ \mathbb{E}\left(\eta_{ij}(t)\eta_{kl}(t^{\prime})\right)={J\over D}\delta_{il}\delta_{jk}\delta(t-t^{\prime})\ .$ (2)
The time evolution in this setting can be regarded as the limiting case of an infinitely long piecewise stochastic evolution
$U=e^{-\text{i}H_{1}t_{1}}e^{-\text{i}H_{2}t_{2}}\cdots e^{-\text{i}H_{n}t_{n}},$ (3)
where each $H_{i}$ is an independent realization of a GUE matrix. This model exhibits several remarkable properties, including rapid convergence to unitary $k$-designs dankert2005efficientsimulationrandomquantum ; Gross_2007 ; ambainis2007quantumtdesignstwiseindependence ; Dankert_2009 ; Cotler_2017 ; Roberts_2017 ; chen2024efficientunitarytdesignsrandom , hyperfast scrambling and the emergence of temperature in de Sitter space lin2022infinitetemperatureshot . Compared to conventional chaotic systems, the Brownian nature of BGUE ensures that ensemble-averaged quantities remain well-defined and analytically trackable, making it a powerful tool for exploring non-equilibrium dynamics.
One of the key motivations for studying BGUE lies in its connection to quantum information theory PhysRevADankert ; Gross2007 ; 4262758 ; Scott_2008 ; Roy_2009 ; Harrow_2009 ; Brand_o_2016 ; Nakata_2017 ; hunterjones2019unitarydesignsstatisticalmechanics . Random matrix models have long served as powerful tools for understanding quantum thermalization, entanglement growth, and operator spreading. In particular, BGUE has been shown to efficiently generate unitary designs, which play a central role in quantum information processing tasks such as classical shadow tomography Aaronson:2017qlb and quantum state reconstruction Tang:2024kpv . Furthermore, an essential feature of disordered quantum systems is their behavior under replica dynamics, a technique in which multiple copies of the system are introduced to probe statistical correlations. Many important quantum information measures—including entanglement entropy, operator norms, Rényi divergences, and multi-contour OTOC Haehl:2017pak —require analyzing multi-replica observables, making the study of $n$-replica time evolution in BGUE a crucial problem. However, existing studies have focused primarily on the cases $n=1,2$, leaving the general $n$-replica case largely unexplored. This serves as one of our motivations to study the general $n$-replica case and develop a systematic way to solve this question.
However, a fundamental challenge in extending the analysis to arbitrary $n$ is the rapid growth in computational complexity. The effective time evolution of the $n$-replica
$\mathcal{U}_{n}(t)\equiv\mathbb{E}\left(U^{\otimes n}\otimes U^{*\otimes n}\right)\equiv e^{\mathcal{L}_{n}t}$ (4)
evolves within a Hilbert space of dimension $D^{2n}$, making direct matrix representations infeasible for large $n$. To simplify this question, we should note that the calculation of the $n$-replica time evolution operator $\mathcal{U}_{n}(t)$ can be reduced to finding the general expression for the moments of $\mathcal{L}_{n}$, allowing for a straightforward expansion:
$\mathcal{U}_{n}(t)=\sum_{r=0}^{\infty}{(\mathcal{L}_{n})^{r}t^{r}\over r!}\ .$ (5)
We find that the linearly independent terms, which we refer to as “graphs”, that appear in the moments of $\mathcal{L}_{n}$ are finite in number, allowing for a natural matrix representation. In the case of $n=2$, the author of Tang:2024kpv categorized 24 individual graphs into 8 groups, reducing the matrix dimension from 24 to 8. However, as $n$ increases, the number of individual graphs that emerge in the moments of $\mathcal{L}_{n}$ grows rapidly (see Table 1), necessitating a systematic method to group these graphs into appropriate categories to achieve a significantly smaller-dimensional representation.
In this work, we develop a graph-theoretic framework for analyzing the time evolution of $n$-replica observables in BGUE. Our approach generalizes and extends the graph-based categorization introduced for $n=2$, introducing a more structured notation that enables an algebraic treatment of replica dynamics. We demonstrate that the moments of $\mathcal{L}_{n}$ can be mapped onto a finite set of graph categories, significantly reducing computational complexity.
For $n=3$, we have 26 graph categories, so that $\mathcal{L}_{n}$ is a $26$-dimensional matrix. This enables us to obtain an explicit analytical expression for
$e^{\mathcal{L}_{3}t}=\sum_{a=1}^{26}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\ ,$ (6)
where the analytical expressions of $f_{a}(t)$ are listed in Appendix A. Furthermore, we discuss the generalization of our framework to arbitrary $n$ and present explicit results for $n=4$.
Compared to representation-theoretic approaches, our graph-based method offers a lower-dimensional irreducible representation, reducing numerical computation time by handling a lower-dimensional matrix. Previous approaches, such as those in Guo:2024zmr , relied on
$[\mathcal{L}_{n},V^{\otimes n}\otimes V^{*\otimes n}]=0,\forall V\in U(D)\ ,$ (7)
which allows for the determination of eigenvalues and wavefunctions for $\mathcal{L}_{n}$ by decomposing direct products of SU$(d)$ into irreducible representations. Moreover, these representations are constructed systematically using Young tableaux Hamermesh1962GroupTA . We will show the detailed result for $n=2$ at the end of Sec. 3.
The graph representation also presents a straightforward way to compute the expectation values of observables. In cases where symmetries like (7) do not exist, the methodology discussed here can serve as a valuable toolbox. Beyond its application to BGUE, our framework has broader implications to other Brownian disordered systems, like random unitary circuit Nahum_2017 ; Nahum_2018 ; von_Keyserlingk_2018 ; Zhou_2020 ; Fisher_2023 and random tensor networks Hayden_2016 ; Hosur:2015ylk . The techniques developed here could provide new insights into the replica method in many-body quantum chaos, as well as potential applications to problems in wormholes and quantum gravity Maldacena_2003 ; saad2019latetimecorrelationfunctions ; stanford2020quantumnoisewormholes ; Peng:2021vhs .
Structure of the Paper
The remainder of this paper is structured as follows. In Sec. 2, we introduce the notation and graph-based formalism for representing the time evolution operator $\mathcal{U}_{n}(t)$. In Sec. 3, we apply our framework to the case $n=2$ as a warm-up example, treated as a double check with the results from Tang:2024kpv . Sec. 4 extends this analysis to the non-trivial case $n=3$, where we show how to find the graph categories and assign the graphs into them. In Sec. 5, we present a general algorithm for arbitrary $n$ and illustrate its application to $n=4$. Finally, in Sec. 6, we discuss potential future directions and broader implications of our findings.
2 Analysis on Graphs
In this section, we introduce a concise notation to represent arbitrary graphs, which facilitates the evaluation of graph contractions and enables a more transparent classification of graphs. We first establish the fundamental concepts, then discuss the inherent redundancies in this notation, and finally analyze the action of the group $\mathcal{L}_{n}$ on individual graphs.
At first, we introduce some notations for the graph representation of $\mathcal{U}_{n}$ and $\mathcal{L}_{n}$. We treat $\mathcal{U}_{n},\mathcal{L}_{n}$ as a $D^{2n}$-dimensional matrix, with row and column indices labeled by two $n$-length lists: $\{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}\}$ and $\{I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}\}$. We then adapt this to the graph representation (taking $n=2$ as an example):
$\left(\mathcal{U}_{2}(t)\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ $\equiv\mathbb{E}\left(U_{I_{1}I_{1}^{\prime}}U_{J_{1}J_{1}^{\prime}}^{*}U_{I_{2}I_{2}^{\prime}}U_{J_{2}J_{2}^{\prime}}^{*}\right)=\mathbb{E}\left[\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\right]=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\ .$ (8)
We label the indices in graphs using uppercase letters $I,J,I^{\prime},J^{\prime}$, where $I/I^{\prime}$ represent the left/right-hand side indices of a contour $H$, and $J/J^{\prime}$ denote the dual contour $H^{*}$. For simplicity, we label the contours (and dual contours) with lowercase letters $i,j,k,l$ (and their corresponding bars $\bar{i},\bar{j},\bar{k},\bar{l}$). Since we take $n=2$, we can also represent $\mathcal{U}_{2}$ as
$\mathcal{U}_{2}(t)=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}$ (9)
By a similar derivation as shown in Tang:2024kpv , it is straightforward to obtain
$\mathcal{L}_{n}=\mathsf{w}\mathbb{I}+{J\over D}\sum_{i\bar{j}}^{n}P_{i\bar{j}}-{J\over D}\sum_{ij}^{n}(X_{ij}+X_{\bar{i}\bar{j}})\quad,$ (10)
where $J$ is the strength of noise, $\mathbb{I}$ denotes identity operator. Other two kinds of operators are represented as below:
$P_{i\bar{j}}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$},~{}X_{ij}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\quad,$ (11)
where each line with two indices $i,j$ at the endpoints of the graph represents a “propagator” between the sites on the contour and the dual contour, i.e., $\delta_{ij}$. Other trivial propagators connecting sites on the same contour, which are represented as straight lines, are given by $\delta_{I_{i}I_{i}^{\prime}}$ or $\delta_{J_{i}J_{i}^{\prime}}$. Taking $n=2$ as an example, we have
$\left(P_{1\bar{2}}\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ =[$\right]_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}=\delta_{I_{1}J_{2}}\delta_{I_{1}^{\prime}J_{2}^{\prime}}\delta_{J_{1}J_{1}^{\prime}}\delta_{I_{2}I_{2}^{\prime}}\ ,$
$\left(X_{12}\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ =[$\right]_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}=\delta_{I_{1}I_{2}^{\prime}}\delta_{I_{2}I_{1}^{\prime}}\delta_{J_{1}J_{1}^{\prime}}\delta_{J_{2}J_{2}^{\prime}}\ .$ (12)
And $\mathsf{w}$ in (10) is a factor that carries the indices of a graph
$\mathsf{w}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}=-\text{i}\sum_{j=1}^{n}(E_{I_{j}}-E_{J_{j}})-nJ,~{}\text{i}=\sqrt{-1}\ .$ (13)
Notice that the factor $\mathsf{w}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}$ commutes with all of the terms in $\mathcal{L}_{n}$. For instance,
$\mathsf{w}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}\left(P_{a\bar{b}}\right)_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}=\left(P_{a\bar{b}}\right)_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}\mathsf{w}_{I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}\ .$ (14)
Thus, from now on, we can treat $\mathsf{w}$ as an ordinary constant and simply write it as $w$. For later use, we define the spectrum factor
$\mathsf{E}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}=-\text{i}\sum_{j=1}^{n}(E_{I_{j}}-E_{J_{j}})\ .$ (15)
Since $\mathsf{w}=\mathsf{E}-nJ\mathbb{I}$, the factor $\mathsf{E}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}$ also commutes with all terms in $\mathcal{L}_{n}$.
2.1 Notations of a Graph
In this work, we employ a concise notation $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$ to represent an arbitrary graph. Here, $0\leq p\leq n$ denotes the number of paired vertices, and $\sigma\equiv\sigma_{H}\otimes\sigma_{H^{*}}\in S_{n}\otimes S_{n}$ encodes the permutation acting on the indices appearing on the right-hand side of the graph. The element of a $D^{2n}$-dimensional matrix $\mathsf{F}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}$ would be:
$\left(\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]\right)_{\{IJ\};\{I^{\prime}J^{\prime}\}}=\prod_{i\not\in\{a\},\bar{j}\not\in\{\bar{b}\}}\delta_{I_{i}I^{\prime}_{\sigma_{1}\left(i\right)}}\delta_{J_{\bar{j}}J^{\prime}_{\sigma\left(\bar{j}\right)}}\prod_{i=1}^{p}\delta_{I_{a_{i}}J_{\bar{b}_{i}}}\delta_{I^{\prime}_{\sigma_{1}\left(a_{i}\right)}J^{\prime}_{\sigma_{2}\left(\bar{b}_{i}\right)}}\ .$ (16)
Here, $c[\sigma]$ is the permutation on the contour labels on the right-hand side, $c_{i\bar{j}}$ represents the pair between the left sites $i,\bar{j}$ and the right sites $\sigma(i)^{\prime},\sigma(\bar{j})^{\prime}$. To make the complicated notations clear, we can take a simple example $c_{1\bar{1}}c[X_{12}\otimes X_{\bar{1}\bar{2}}]=\delta_{I_{1}J_{1}}\delta_{I_{2}^{\prime}J_{2}^{\prime}}\delta_{I_{2}I_{1}^{\prime}}\delta_{J_{2}J_{1}^{\prime}}$ and show how the graphs work as follows:
[…]
The filled dots are the sites on the contour $H$ and the hollow circles are those on the dual contour $H^{*}$. And we can directly find how the swapping and connecting operations work in the first two steps. The graph is finally denoted as the R.H.S. of (17).
The notations presented in (16) for the graphs make it much easier to calculate the contraction (product) of two graphs and study the action of $\mathcal{L}_{n}$ on a certain graph, as we will see later in this section.
Symmetries of a Graph
Before examining the action of $P_{i\bar{j}}$ and $X_{ij},X_{\bar{i}\bar{j}}$ on a graph $\mathsf{F}$, we note that the notations presented above have some redundancy. For a graph $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$, the vertices for each connected pair $(a_{i},\bar{b}_{i})$ are identical. To simplify it, we define the action of a permutation on the graph as
$\sigma^{\prime}*\mathsf{F}$ $\equiv\sigma^{\prime}(\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}})c[\sigma\sigma^{\prime}]=\prod_{i=1}^{p}c_{\sigma^{\prime}_{1}(a_{i})\sigma^{\prime}_{2}(\bar{b}_{i})}c[\sigma\sigma^{\prime}]\ .$ (18)
If a permutation $\sigma^{\prime}=\sigma^{\prime}_{H}\otimes\sigma^{\prime}_{H^{*}}$ is an element of the group $G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})$ generated by $\{X_{a_{i}a_{j}}\otimes X_{\bar{b}_{i}\bar{b}_{j}}\}_{i,j=1,2,\ldots,p}$, then its action is trivial: $\sigma^{\prime}*\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b_{i}}}c[\sigma\sigma^{\prime}]=\mathsf{F}$. It can be observed that $G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})$ resembles the gauge group in quantum field theory and is isomorphic to the permutation group $S_{p}$.
Number of Individual Graphs
$n$ $N_{n}$#\## Graph Categories
1 2 2
2 24 8
3 720 26
4 40320 /
5 3628800 /
Table 1: Number of individual graphs and graph categories.
For a given $n$, the task is to count the number of individual graphs $N_{n}$ generated by $\mathcal{L}_{n}$. A simple counting yields:
$N_{1}=2,\quad N_{2}=24.$
For general $n$, using $|G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})|=p!$ and $|S_{n}\otimes S_{n}|=(n!)^{2}$, we have
$N_{n}$ $=\sum_{\boldsymbol{a},\bar{\boldsymbol{b}}}{|S_{n}\otimes S_{n}|\over|G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})|}=\sum_{p=0}^{n}\binom{n}{p}^{2}p!\frac{\left(n!\right)^{2}}{p!}=\sum_{p=0}^{n}\left(n!\binom{n}{p}\right)^{2}\equiv\sum_{p=0}^{n}N_{n,p}\ .$ (19)
One can utilize the individual graphs as a basis, then $\mathcal{L}_{n}\mathsf{F}_{i}=\sum_{j=1}^{N_{n}}\widetilde{M}_{ji}\mathsf{F}_{j}$. However, as indicated in (19), the dimension of $\widetilde{M}$ grows too quickly with $n$. A clever approach is to group these individual graphs into categories $\mathsf{F}_{\alpha}=\sum_{i=1}^{N_{n;\alpha}}\mathsf{F}_{\alpha,i}$ such that
$\mathcal{L}_{n}\mathsf{F}_{\alpha}=\sum_{\beta=1}^{N_{n}^{\text{ca}}}M_{\beta\alpha}\mathsf{F}_{\beta}\ .$ (20)
In this paper, we further require that $\mathsf{F}_{\alpha}\cap\mathsf{F}_{\beta}=\emptyset$ for all $\alpha\neq\beta$. The dimension of the transfer matrix $M$, denoted as $N_{n}^{\text{ca}}$, can be significantly smaller than $N_{n}$, as illustrated in Table 1.
2.2 Operating $\mathcal{L}_{n}$ on an Individual Graph
Figure 1: Four non-trivial cases of $P_{i\bar{j}}\mathsf{F}$.
Since $\mathcal{L}_{n}$ consists of $P_{i\bar{j}}$ and $X_{ij},X_{\bar{i}\bar{j}}$, we first evaluate their action on an individual graph $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$. We denote the set of paired sites as $C(\mathsf{F})=\{(a_{i},\bar{b}_{i})\}_{i=1}^{p}$. For later reference, we also define the set of all possible paired sites between different contours as $\mathcal{C}$ and those on the same contour as $\mathcal{E}$ and $\bar{\mathcal{E}}$, respectively:
[…]
Operation of $P_{i\bar{j}}$
There are five cases:
$P_{i\bar{j}}\mathsf{F}=\begin{cases}c_{i\bar{j}}\mathsf{F},&\text{if}\ (i,\bar{j})\cap C(\mathsf{F})=\emptyset\\ X_{\bar{j}\bar{b}_{1}}*\mathsf{F},&\text{if}\ (i,\bar{j})\cap C(\mathsf{F})=\{a_{1}\}\\ X_{ia_{1}}*\mathsf{F},&\text{if}\ (i,\bar{j})\cap C(\mathsf{F})=\{\bar{b}_{1}\}\\ D\mathsf{F},&\text{if}\ (i,\bar{j})\cap C(\mathsf{F})=\{a_{1},\bar{b}_{1}\}\\ X_{a_{1}a_{2}}*\mathsf{F}=X_{\bar{b}_{1}\bar{b}_{2}}*\mathsf{F},&\text{if}\ (i,\bar{j})\cap C(\mathsf{F})=\{a_{1},\bar{b}_{2}\}\end{cases}\ .$ (22)
Since the first case is trivial, we illustrate the other four cases in Figure 1.
Operation of $X_{ij},{X}_{\bar{i}\bar{j}}$
There are five cases depending on the result of $(i,j)\cap C(\mathsf{F})$. We depict two nontrivial cases in Figure 2. It is observed that we always have
$X_{ij}\mathsf{F}=X_{ij}*\mathsf{F},$
and the action of $X_{\bar{i}\bar{j}}$ is given by
$X_{\bar{i}\bar{j}}\mathsf{F}=X_{\bar{i}\bar{j}}*\mathsf{F}.$
Operation of $\mathcal{L}_{n}$
Next, we can study the operation of $\mathcal{L}_{n}$. Notice that we have the condition $\bar{u}\ \cap\ C[\mathsf{F}]=\emptyset$. Hence the permutations $X_{u}$ and $X_{\bar{u}}$ act trivially on $C(\mathsf{F})$. Thus, we ultimately find a simple relation
$\mathcal{L}_{n}\mathsf{F}=\left(w+|C[\mathsf{F}]|J\right)\mathsf{F}+\frac{J}{D}\sum_{u\cap C[\mathsf{F}]=\emptyset}c_{u}\mathsf{F}-\frac{J}{D}\sum_{u\cap C[\mathsf{F}]=\emptyset}\mathsf{F}[\sigma X_{u}]-\frac{J}{D}\sum_{\bar{u}\cap C[\mathsf{F}]=\emptyset}\mathsf{F}[\sigma X_{\bar{u}}]$ (23)
where $|C[\mathsf{F}]|$ counts the number of pairs, and $\mathsf{F}[\sigma\sigma^{\prime}]\equiv\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma\sigma^{\prime}]$. The definitions of $u$ and $\bar{u}$ are encoded in the summation expressions. For example, in the term $\sum_{u\cap C[\mathsf{F}]=\emptyset}c_{u}\mathsf{F}$, $u$ can be considered as a pair in $\mathcal{C}$ that is excluded by $C(\mathsf{F})$. Meanwhile, in the term $\sum_{u\cap C[\mathsf{F}]=\emptyset}\mathsf{F}[\sigma X_{u}]$, $u$ can be regarded as an element of $\mathcal{E}$ that is excluded by $C(\mathsf{F})$. Here we choose the set of $u$ to meet those conditions of $u\cap C[\mathsf{F}]=\emptyset$ from all the summations, which requires flattening $C(\mathsf{F})$ into a list of site labels and find those $u$ that are available. For example, taking $u=\{(1,\bar{1})\},C(\mathsf{F})=\{(1,\bar{2}),(2,\bar{3})\}$, then
$u\cap C(\mathsf{F})=\{1,\bar{1}\}\cap\{1,\bar{2},2,\bar{3}\}=\{1\}\not=\emptyset\ ,$ (24)
which shows that $u=\{(1,\bar{1})\}$ is not available.
We always define three graph categories, as they play a crucial role in the definition of $\mathcal{L}_{n}$:
$\mathsf{F}_{0}\equiv c[I],\mathsf{F}_{1}\equiv\sum_{u}c_{u}c[I],\mathsf{F}_{0,X+\bar{X}}\equiv\sum_{e}c[X_{e}]+\sum_{\bar{e}}c[X_{\bar{e}}]\ .$ (25)
Then we apply $\mathcal{L}_{n}$ to the three initial categories and their descendants until a complete basis is obtained. After that, we obtain the expression for the transfer matrix $M$ in (20), which allows us to solve $\mathcal{U}_{n}(t)$ using (5). We will illustrate the basic idea with examples for $n=2$ and $n=3$ in the next two sections.
Figure 2: Two non-trivial cases of $X_{ij}\mathsf{F}$.
3 Time Evolution for $n=2$
As a warm-up, we consider $n=2$, which gives us four sites: $\{1,\bar{1},2,\bar{2}\}$. The permutation group is $S_{2}\otimes S_{2}$, where $S_{2}=\{I,X_{12}\}$. Then
$\mathcal{C}$ =$=\left\{\left(1,\bar{1}\right),\left(1,\bar{2}\right),\left(2,\bar{1}\right),\left(2,\bar{2}\right)\right\}\ ,$ (26)
[…]
So we have
$\mathcal{L}_{2}=w\mathbb{I}+{J\over D}\left(P_{1\bar{1}}+P_{2\bar{2}}+P_{1\bar{2}}+P_{2\bar{1}}\right)-{J\over D}(X_{12}+X_{\bar{1}\bar{2}})\ .$ (27)
The number of individual graphs is $24$, while the number of operators ($P_{i\bar{j}},X_{ij},X_{\bar{i}\bar{j}}$) is $6$. Therefore, in principle, one needs to calculate $6\times 24=168$ graph contractions when calculating $M$ as defined earlier. Fortunately, (23) enables us to perform a simple algebraic calculation without drawing any graphs.
3.1 Counting Graph Categories
As we have discussed earlier, we will find the appropriate graph groups by applying $\mathcal{L}_{n}$ to the three initial categories $\mathsf{F}_{0},\mathsf{F}_{1},\mathsf{F}_{0,X+\bar{X}}$ and their descendants until no new graph category is generated. In the case of $n=2$, we will perform $\mathcal{L}_{2}$ on the initial sets as follows.
$\mathcal{L}_{2}$ acts on $\mathsf{F}_{0},\mathsf{F}_{1},\mathsf{F}_{0,X+\bar{X}}$
  • $\mathcal{L}_{2}\mathsf{F}_{0}$: The operation of $\mathcal{L}_{2}$ on $\mathsf{F}_{0}$ is trivial, since $\mathsf{F}_{0}$ is a trivial set
    $\mathcal{L}_{2}\mathsf{F}_{0}=w\mathsf{F}_{0}+{J\over D}\mathsf{F}_{1}-{J\over D}\mathsf{F}_{0,X+\bar{X}}\ .$ (28)
  • $\mathcal{L}_{2}\mathsf{F}_{1}$: By definition we have
    $\mathcal{L}_{2}\mathsf{F}_{1}=\left(w+J\right)\mathsf{F}_{1}+\frac{J}{D}\sum_{u\not=v}c_{u}c_{v}c[I]-\frac{J}{D}\left(\sum_{e\cap u=\emptyset}c_{u}c[X_{e}]+\sum_{\bar{e}\cap u=\emptyset}c_{u}c[X_{\bar{e}}]\right)\ .$ (29)
    Notice that the condition $e\cap u=\emptyset$, where $e\in\mathcal{E}$ and $u\in\mathcal{C}$, cannot be satisfied for $n=2$. This is an important fact that simplifies the calculations for $n=2$. Therefore, we have
    $\mathcal{L}_{2}\mathsf{F}_{1}=\left(w+J\right)\mathsf{F}_{1}+\frac{2J}{D}\mathsf{F}_{2}\ ,$ (30)
    where we have defined
    $\mathsf{F}_{2}\equiv\frac{1}{2!}\sum_{u\not=v}c_{u}c_{v}c[I]=c_{1\bar{1}}c_{2\bar{2}}c[I]+c_{1\bar{2}}c_{2\bar{1}}c[I]\ .$ (31)
  • $\mathcal{L}_{2}\mathsf{F}_{0,X+\bar{X}}$: It is direct to obtain
    $\mathcal{L}_{2}\mathsf{F}_{0,X+\bar{X}}$ $=w\mathsf{F}_{0,X+\bar{X}}+\frac{J}{D}\sum_{u}c_{u}\mathsf{F}_{0,X+\bar{X}}$
    $\hskip 15.0pt-\frac{J}{D}\sum_{e,u}\left(c[X_{e}X_{u}]+c[X_{\bar{e}}X_{u}]+c[X_{e}X_{\bar{u}}]+c[X_{\bar{e}}X_{\bar{u}}]\right)\ ,$ (32)
    then we can define $\mathsf{F}_{1,X+\bar{X}}\equiv\sum_{u}c_{u}\mathsf{F}_{0,X+\bar{X}}$. Notice that
    $\sum_{e,u}\left(c[X_{e}X_{u}]+c[X_{\bar{e}}X_{u}]+c[X_{e}X_{\bar{u}}]+c[X_{\bar{e}}X_{\bar{u}}]\right)=2c[I]+2c[X_{12}X_{\bar{1}\bar{2}}]\ .$ (33)
    So we can define $\mathsf{F}_{0,X\bar{X}}\equiv c[X_{12}X_{\bar{1}\bar{2}}]$, then
    $\mathcal{L}_{2}\mathsf{F}_{0,X+\bar{X}}=w\mathsf{F}_{0,X+\bar{X}}+\frac{J}{D}\mathsf{F}_{1,X+\bar{X}}-\frac{2J}{D}\left(\mathsf{F}_{0}+\mathsf{F}_{0,X\bar{X}}\right)\ .$ (34)
We find that three new graph categories, $\mathsf{F}_{2},\mathsf{F}_{1,X+\bar{X}},\mathsf{F}_{0,X\bar{X}}$, emerge after applying $\mathcal{L}_{2}$ on initial sets, which is the first order calculations. Next, we consider the second order, which means the operation of $\mathcal{L}_{2}$ on these new categories. For clarity, we restate the definitions here
$\mathsf{F}_{2}=c_{1\bar{1}}c_{2\bar{2}}c[I]+c_{1\bar{2}}c_{2\bar{1}}c[I]\ ,$ (35)
$\mathsf{F}_{1,X+\bar{X}}=\sum_{u}c_{u}\mathsf{F}_{0,X+\bar{X}}\ ,$ (36)
$\mathsf{F}_{0,X\bar{X}}=c[X_{12}X_{\bar{1}\bar{2}}]\ .$ (37)
$\mathcal{L}_{2}$ acts on $\mathsf{F}_{2},\mathsf{F}_{1,X+\bar{X}},\mathsf{F}_{0,X\bar{X}}$
  • $\mathcal{L}_{2}\mathsf{F}_{2}$: Notice that only the first term in (23) survives. The operation of $\mathcal{L}_{2}$ on $\mathsf{F}_{2}$ is straightforward to obtain
    $\mathcal{L}_{2}\mathsf{F}_{2}$ $=\left(w+2J\right)\mathsf{F}_{2}\ .$ (38)
  • $\mathcal{L}_{2}\mathsf{F}_{1,X+\bar{X}}$: Using (23), we have
    $\mathcal{L}_{2}\mathsf{F}_{1,X+\bar{X}}$ $=\left(w+J\right)\mathsf{F}_{1,X+\bar{X}}+\frac{J}{D}\left(\sum_{e}\sum_{g\cap u=\emptyset}c_{g}c_{u}c[X_{e}]+\left(e\leftrightarrow\bar{e}\right)\right)\ ,$ (39)
    so we can define
    $\mathsf{F}_{2,X+\bar{X}}\equiv\frac{1}{4}\sum_{e}\sum_{g\cap u=\emptyset}c_{g}c_{u}c[X_{e}]+\left(e\leftrightarrow\bar{e}\right)=c_{1\bar{1}}c_{2\bar{2}}c[X_{12}]+c_{1\bar{2}}c_{2\bar{1}}c[X_{12}]\ .$ (40)
    Finally, we have
    $\mathcal{L}_{2}\mathsf{F}_{1,X+\bar{X}}=\left(w+J\right)\mathsf{F}_{1,X+\bar{X}}+{4J\over D}\mathsf{F}_{2,X+\bar{X}}\ .$ (41)
  • $\mathcal{L}_{2}\mathsf{F}_{0,X\bar{X}}$: Using (23), we have
    $\mathcal{L}_{2}\mathsf{F}_{0,X\bar{X}}$ $=w\mathsf{F}_{0,X\bar{X}}+\frac{J}{D}\sum_{u}c_{u}c[X_{12}X_{\bar{1}\bar{2}}]-\frac{J}{D}\left(\sum_{u}c[X_{12}X_{\bar{1}\bar{2}}X_{u}]+\sum_{u}c[X_{12}X_{\bar{1}\bar{2}}X_{\bar{u}}]\right)$
    $=w\mathsf{F}_{0,X\bar{X}}+\frac{J}{D}\mathsf{F}_{1,X\bar{X}}-\frac{J}{D}\left(c[X_{\bar{1}\bar{2}}]+c[X_{12}]\right)$
    $=w\mathsf{F}_{0,X\bar{X}}+\frac{J}{D}\mathsf{F}_{1,X\bar{X}}-\frac{J}{D}\mathsf{F}_{0,X+\bar{X}}$ (42)
    where we have used $X_{u}=\{X_{12}\}$, so that $X_{12}X_{\bar{1}\bar{2}}X_{u}=X_{\bar{1}\bar{2}},X_{12}X_{\bar{1}\bar{2}}X_{\bar{u}}=X_{12}$.
Here, we find two new graph categories, $\mathsf{F}_{2,X+\bar{X}}$ and $\mathsf{F}_{1,X\bar{X}}$. We will then study the third order calculations by operating $\mathcal{L}_{2}$ on these categories
$\mathsf{F}_{2,X+\bar{X}}=c_{1\bar{1}}c_{2\bar{2}}c[X_{12}]+c_{1\bar{2}}c_{2\bar{1}}c[X_{12}]\ ,$ (43)
$\mathsf{F}_{1,X\bar{X}}=\sum_{u}c_{u}c[X_{12}X_{\bar{1}\bar{2}}]\ .$ (44)
$\mathcal{L}_{2}$ acts on $\mathsf{F}_{2,X+\bar{X}},\mathsf{F}_{1,X\bar{X}}$
The calculations are straightforward by using (23)
  • $\mathcal{L}_{2}\mathsf{F}_{2,X+\bar{X}}$:
    $\mathcal{L}_{2}\mathsf{F}_{2,X+\bar{X}}=\left(w+2J\right)\mathsf{F}_{2,X+\bar{X}}\ .$ (45)
  • $\mathcal{L}_{2}\mathsf{F}_{1,X\bar{X}}$:
    $\mathcal{L}_{2}\mathsf{F}_{1,X\bar{X}}$ $=\left(w+J\right)\mathsf{F}_{1,X\bar{X}}+\frac{J}{D}\sum_{v\cap u=\emptyset}\sum_{u}c_{v}c_{u}c[X_{12}X_{\bar{1}\bar{2}}]$ (46)
    $=\left(w+J\right)\mathsf{F}_{1,X\bar{X}}+\frac{J}{D}\sum_{v\cap u=\emptyset}\sum_{u}c_{v}c_{u}c[I]$
    $=\left(w+J\right)\mathsf{F}_{1,X\bar{X}}+\frac{2J}{D}\mathsf{F}_{2}\ ,$
    where we have used the gauge redundancy: $c_{u}c_{v}c[\sigma X_{12}X_{\bar{1}\bar{2}}]=c_{u}c_{v}c[\sigma]$.
Rank $r$ New Categories #$\#\mathsf{F}$
1 $\mathsf{F}_{0},\mathsf{F}_{1},\mathsf{F}_{0,X+\bar{X}}$ 1+4+2=7
2 $\mathsf{F}_{2},\mathsf{F}_{1,X+\bar{X}},\mathsf{F}_{0,X\bar{X}}$ 2+8+1=11
3 $\mathsf{F}_{2,X+\bar{X}},\mathsf{F}_{1,X\bar{X}}$ 2+4=6
4 $\emptyset$ 0
Table 2: 8 categories for $n=2$.
So far, we find that no new categories emerge. In summary, we list the total of 8 categories in Table 2.
3.2 Observables
Comparing with the categories in Tang:2024kpv , we can choose
$\{\mathsf{F}_{a}\}_{a=1}^{8}=\{\mathsf{F}_{0,I},\mathsf{F}_{1,I},\mathsf{F}_{0,X+\bar{X}},\mathsf{F}_{1,X+\bar{X}},\mathsf{F}_{2,I},\mathsf{F}_{0,X\bar{X}},\mathsf{F}_{2,X\bar{X}},\mathsf{F}_{1,X\bar{X}}\}\ ,$ (47)
then we obtain the matrix representation of $\mathcal{L}_{2}$:
$M=\left(\begin{array}[]{cccccccc}w&0&-\frac{2J}{D}&0&0&0&0&0\\ \frac{J}{D}&J+w&0&0&0&0&0&0\\ -\frac{J}{D}&0&w&0&0&-\frac{J}{D}&0&0\\ 0&0&\frac{J}{D}&J+w&0&0&0&0\\ 0&\frac{2J}{D}&0&0&2J+w&0&0&\frac{2J}{D}\\ 0&0&-\frac{2J}{D}&0&0&w&0&0\\ 0&0&0&\frac{4J}{D}&0&0&2J+w&0\\ 0&0&0&0&0&\frac{J}{D}&0&J+w\\ \end{array}\right)\ .$ (56)
We can use the same method as in the reference Tang:2024kpv to solve $\mathcal{U}_{2}$:
$\mathcal{U}_{2}(t)\equiv\sum_{i=1}^{8}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\ .$ (57)
where $f_{a}(t)$ are given by
$\begin{bmatrix}f_{1}(t)\\ f_{2}(t)\\ f_{3}(t)\\ f_{4}(t)\\ f_{5}(t)\\ f_{6}(t)\\ f_{7}(t)\\ f_{8}(t)\end{bmatrix}=\begin{bmatrix}0&0&\frac{1}{4}&\frac{1}{2}&\frac{1}{4}\\ 0&\frac{D^{2}-2}{D\left(D^{2}-4\right)}&\frac{-1}{4(D-2)}&\frac{-1}{2D}&\frac{-1}{4(D+2)}\\ 0&0&-\frac{1}{4}&0&\frac{1}{4}\\ 0&\frac{-1}{D^{2}-4}&\frac{1}{4(D-2)}&0&\frac{-1}{4(D+2)}\\ \frac{1}{D^{2}-1}&\frac{-2}{D^{2}-4}&\frac{1}{2(D-1)(D-2)}&0&\frac{1}{2(D+1)(D+2)}\\ 0&0&\frac{1}{4}&-\frac{1}{2}&\frac{1}{4}\\ \frac{-1}{D^{3}-D}&\frac{4}{D(D^{2}-4)}&\frac{-1}{2(D-1)(D-2)}&0&\frac{1}{2(D+1)(D+2)}\\ 0&\frac{2}{D(D^{2}-4)}&\frac{-1}{4(D-2)}&\frac{1}{2D}&\frac{-1}{4(D+2)}\end{bmatrix}\times\begin{bmatrix}1\\ e^{-Jt}\\ e^{-(2-2D)^{-1}Jt}\\ e^{-2Jt}\\ e^{-(2+2D)^{-1}Jt}\end{bmatrix}\ .$ (58)
Generally, we can calculate OTOC by using the expression of (57)
$\text{Tr}\left(O_{1}(t)O_{2}O_{3}(t)O_{4}\right)=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\equiv\sum_{a=1}^{8}f_{a}(t)\times\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\right)_{O_{1},O_{2},O_{3},O_{4}}\ .$ (59)
A special case would be the noise-free case, which means that the noise strength $J$ vanishes. We expect that the time evolution will be simplified into a unitary one. Note that from (58) we can directly find that all $f_{a}(t)$ vanishes except $f_{1}(t)=1$. Hence
$\text{Tr}\left(O_{1}(t)O_{2}O_{3}(t)O_{4}\right)=\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{0}\right)_{O_{1},O_{2},O_{3},O_{4}}\ .$ (60)
which is just the unitary time evolution for a closed system and finishes our consistency check. Thus, we will be more confident to use (59) to calculate the fluctuations of two-point functions, spectral form factor (SFF) and other observables.
To end this section, we would mention that there is an alternative method Guo:2024zmr to evaluate $\mathcal{U}_{2}(t)$ by observing that
$[\mathcal{L}_{n},V^{\otimes n}\otimes V^{*\otimes n}]=0,~{}\forall V\in U(D)\ .$ (61)
We can always decompose a wavefunction $\psi=\psi_{i_{1}i_{2}}^{j_{1}j_{2}}\ket{i_{1},j_{1},i_{2},j_{2}}_{1\bar{1}2\bar{2}}$ into two singlets $\mathbf{1}$, four adjoints $\mathbf{D}^{2}-\mathbf{1}$, an antisymmetric traceless $\mathbf{A}$, two mixed $\mathbf{M}$ and a symmetric traceless $\mathbf{S}$ rank $(2,2)$ irreducible representations of $\text{SU}(D)$, with all the coefficients. Thus, we can also do the exponentiation on this basis. However, notice that the dimension of the representation is $10$, which is larger than the number of independent graph categories $8$, indicating that there are still some redundancies in the Young tableaux method. This might cause a longer processing time in numerical calculations.
4 Time Evolution for $n=3$
Then we consider $n=3$, so we have six sites: $\{1,\bar{1},2,\bar{2},3,\bar{3}\}$. The permutation group is $S_{3}\otimes S_{3}$, $S_{3}=\{I,X_{12},X_{13},X_{23},X_{123},X_{132}\}$, then
[…]
$p$ $\mathsf{F}_{a}$$\#\mathsf{F}$ $I$ $X+\bar{X}$ $X\bar{X}$ $XX+\bar{X}\bar{X}$ $XX\bar{X}+\bar{X}\bar{X}X$ $XX\bar{X}\bar{X}$
0 1 6 9 4 12 4
1 9 $18^{(e)},36^{(i)}$ $9^{(ee)},36^{(ei)},36^{(ii)}$ 36 $36^{(e)},72^{(i)}$ 36
2 18 $72^{(ei)},18^{(ii)}$ $36^{(ei,ei)},36^{(ei,ie)},72^{(ei,ii)}$ 0 72 0
3 6 18 0 12 0 0
Table 3: The graph categories for $n=3$.
4.1 Grouping Graphs into Categories
Here we first define some notations:
$X\equiv\{X_{12},X_{13},X_{23}\},~{}[XX]\equiv\{X_{123},X_{132}\}\ .$ (63)
As discussed earlier, the number of individual graphs with $p$ connected pairs is given by $N_{n,p}=(n!C_{n}^{p})^{2}$. For $n=3$, it is
$N_{3,p}=\{36,324,324,36\}\ ,\quad\text{for }p=\{0,1,2,3\}\ \ .$ (64)
Next, we discuss the graph categories $\mathsf{F}_{p,\alpha}$ with $p$ connected pairs. For a graph with $p$ pairs, notice that the second term in (23) ($\sum_{u\cap C[\mathsf{F}]=\emptyset}c_{u}\mathsf{F}$) represents graphs with $p+1$ pairs, while the last two terms indicate applying additional 2 swaps on $\mathsf{F}$, which are excluded by $C(\mathsf{F})$.
  • For $p=0$, there are no restrictions on $X$.
  • For $p=1$, with $C[\mathsf{F}]=c_{u}$, we should require $X\cap u=\emptyset$. We can define the allowed set $X^{(0)}=X^{(ex)}[u]\equiv\{X|X\cap u\neq\emptyset\}$ and its complement as $X^{(1)}$. For example, if $u=(1\bar{1})$, then we have $X^{(0)}=\{X_{23}\}$ and $X^{(1)}=X-X^{(0)}=\{X_{12},X_{13}\}$.
  • For $p>1$, the condition $X\cap u=\emptyset$ cannot be satisfied when $n=3$.
While the discussion of graph categories by connected pairs is complete for $n=3$, we could generalize the definitions of $X^{(0)}$ and $X^{(1)}$ to $X^{(\alpha)}$ for generic $n$ and $p$. We define the conditions set as follows
$\mathsf{Cons}\equiv\otimes^{p}\{ex,in\}\ .$ (65)
For simplicity, we just denote $ex=0,in=1$, and define
$\text{Con}(0,A,u)\equiv(A\cap u=\emptyset),~{}\text{Con}(1,A,u)\equiv(A\cap u\not=\emptyset)\ ,$ (66)
showing that whether $u$ intersects with $A$ or not. Hence for a series of certain connected pairs $u_{1},u_{2},\ldots,u_{p}$, $X^{(\alpha)}$ is a set of 2-swap operators:
$X^{(\alpha)}=\{X_{e}|\cap_{j=1}^{p}\text{Con}(\alpha[j],e,u_{j})\}\ .$ (67)
In our case, for instance, we take $n=3,p=2$, then the condition set is $\textsf{Cons}=\{00,01,10,11\}$. And if we further choose $u_{1}=(1\bar{1}),u_{2}=(2\bar{2})$, we have
$X^{(00)}=\emptyset,X^{(01)}=\{X_{23}\},X^{(10)}=\{X_{13}\},X^{(11)}=\{X_{12}\}\ .$ (68)
Equipped with the new categories $\mathsf{F}_{0,I}$, we could calculate the operations of $\mathcal{L}_{n}^{r}$ on those categories. We will find that $\mathcal{L}_{n}^{r}\mathsf{F}_{0,I}$ can be expressed as a linear combination of graphs represented by permutations
$\sum_{u_{1},u_{2},\ldots,u_{p}}c(u_{1},u_{2},\ldots,u_{p})\ c\left[\prod_{j=1}^{m}(X+\bar{X})^{(\alpha_{j})}\right]$ (69)
where the product should be understood as the product between sets, which means the repeated terms should be eliminated. And the summation runs over all possible permutations. We will primarily discuss how to assign these blocks (69) into the appropriate graph categories in the followings.
Zero Pair: $C[\mathsf{F}]=0$, #$\#\mathsf{F}=36$
For $p=0$, (69) becomes
$c\left[\prod_{j=1}^{m}(X+\bar{X})\right],m=0,1,2,3,\ldots\ .$ (70)
Notice that for $n=3$, we have the following relation for permutation sets
$XX=3I+3[XX],~{}[XX]X=2X\ .$ (71)
Hence all terms appearing in $(X+\bar{X})^{m}$ are
$I,X+\bar{X},X\bar{X},[XX]+[\bar{X}\bar{X}],[XX]\bar{X}+[\bar{X}\bar{X}]X,[XX][\bar{X}\bar{X}]\ .$ (72)
Since the gauge group vanishes, there is no redundancy. Hence we obtain $6$ categories
$\mathsf{F}_{0,I},\mathsf{F}_{0,X+\bar{X}},\mathsf{F}_{0,X\bar{X}},\mathsf{F}_{0,[XX]+[\bar{X}\bar{X}]},\mathsf{F}_{0,[XX]\bar{X}+[\bar{X}\bar{X}]X},\mathsf{F}_{0,[XX][\bar{X}\bar{X}]}\ ,$ (73)
where $F_{0,A}\equiv c[A]=\cup_{x\in A}c[x]$, or explicitly:
$\mathsf{F}_{0,I}\equiv c[I],\quad\mathsf{F}_{0,X+\bar{X}}\equiv\sum_{e}c[X_{e}]+\sum_{\bar{e}}c[X_{\bar{e}}]+(X\leftrightarrow\bar{X})\ ,$ (74)
$\mathsf{F}_{0,X\bar{X}}\equiv\sum_{u,\bar{e}}c[X_{\bar{e}}X_{u}]+(X\leftrightarrow\bar{X})\ ,$
$\mathsf{F}_{0,[XX]+[\bar{X}\bar{X}]}\equiv c[X_{123}]+c[X_{132}]+c[X_{\bar{1}\bar{2}\bar{3}}]+c[X_{\bar{1}\bar{3}\bar{2}}]+(X\leftrightarrow\bar{X})\ ,$
$\mathsf{F}_{0,[XX]\bar{X}+[\bar{X}\bar{X}]X}\equiv\sum_{e}c\left[X_{\bar{1}\bar{2}\bar{3}}X_{e}+X_{\bar{1}\bar{3}\bar{2}}X_{e}\right]+\sum_{\bar{e}}c\left[X_{123}X_{\bar{e}}+X_{132}X_{\bar{e}}\right]+(X\leftrightarrow\bar{X})\ ,$
$\mathsf{F}_{0,[XX][\bar{X}\bar{X}]}\equiv c\left[X_{123}X_{\bar{1}\bar{2}\bar{3}}\right]+c\left[X_{123}X_{\bar{1}\bar{3}\bar{2}}\right]+c\left[X_{132}X_{\bar{1}\bar{2}\bar{3}}\right]+c\left[X_{132}X_{\bar{1}\bar{3}\bar{2}}\right]+(X\leftrightarrow\bar{X})\ .$
It is straightforward to count their sizes, which are $\{1,6,9,4,12,4\}$ respectively. Thus, we have $1+6+9+4+12+4=36$ graphs, as we have calculated in (64).
One Pair: $C[\mathsf{F}]=1$, #$\#\mathsf{F}=324$
For $p=1$, (69) becomes
$\sum_{u}c_{u}\left[\prod_{j=1}^{m}(X+\bar{X})^{(\alpha_{j})}\right],m=0,1,2,3,\ldots,\alpha_{j}\in\{(0),(1)\}\ .$ (75)
The size of the gauge group is just $p!=1$, so there is no redundancy. For arbitrary $u=(i\bar{j})$, we find that
#$\#X^{(0)}=1,~{}\#X^{(1)}=2,~{}X=X^{(0)}+X^{(1)}\ .$ (76)
Hence we have
$X^{(0)}X^{(0)}=\{I\},~{}X^{(0)}X^{(1)}=X^{(1)}X^{(0)}=[XX],~{}X^{(1)}X^{(1)}=2I+2[XX]\ .$ (77)
All terms which can appear are
$I,X^{(0)}+\bar{X}^{(0)},X^{(1)}+\bar{X}^{(1)},X^{(0)}\bar{X}^{(0)},X^{(0)}\bar{X}^{(1)}+\bar{X}^{(0)}X^{(1)},X^{(1)}\bar{X}^{(1)},$ (78)
$[XX]+[\bar{X}\bar{X}],[XX]\bar{X}^{(0)}+[\bar{X}\bar{X}]X^{(0)},[XX]\bar{X}^{(1)}+[\bar{X}\bar{X}]X^{(1)},[XX][\bar{X}\bar{X}]\ .$
So there are $10$ categories
$\mathsf{F}_{1,I},\mathsf{F}_{1,X}^{(0)},\mathsf{F}_{1,X}^{(1)},\mathsf{F}_{1,X\bar{X}}^{(00)},\mathsf{F}_{1,X\bar{X}}^{(01)},\mathsf{F}_{1,X\bar{X}}^{(11)},\mathsf{F}_{1,XX},\mathsf{F}_{1,XX\bar{X}}^{(0)},\mathsf{F}_{1,XX\bar{X}}^{(1)},\mathsf{F}_{1,XX\bar{X}\bar{X}}$ (79)
with there sizes to be $\{1,18,36,9,36,36,36,36,72,36\}$.
Two Pairs: $C[\mathsf{F}]=2$,#$\#\mathsf{F}=324$
For $p=2$, (69) becomes
$\sum_{u,v}c_{u}c_{v}\left[\prod_{j=1}^{m}(X+\bar{X})^{(\alpha_{j})}\right],m=0,1,2,3,\ldots,\alpha_{j}\in\{(00),(01),(10),(11)\}\ .$ (80)
$\forall u,v,\ u\cap v=\emptyset$, we have
$X^{(00)}=\emptyset,X=\{X^{(01)},X^{(10)},X^{(11)}\}\ .$ (81)
So all terms which can appear are
$I,X^{(01)+(10)}+\bar{X}^{(01)+(10)},X^{(11)}+\bar{X}^{(11)},X^{(01)}\bar{X}^{(01)}+X^{(10)}\bar{X}^{(10)},$ (82)
$X^{(01)}\bar{X}^{(10)}+X^{(10)}\bar{X}^{(01)},X^{(01)+(10)}\bar{X}^{(11)}+X^{(11)}\bar{X}^{(01)+(10)},X^{(11)}\bar{X}^{(11)},[XX]+[\bar{X}\bar{X}],$
$[XX]\bar{X}^{(01)+(10)}+[\bar{X}\bar{X}]X^{(01)+(10)},[XX]\bar{X}^{(11)}+[\bar{X}\bar{X}]X^{(11)},[XX][\bar{X}\bar{X}]\ .$
But for $p=2$, the gauge group generated by $X^{(11)}\otimes\bar{X}^{(11)}$ has size $p!=2$. We have the equivalent relation
$c_{u}c_{v}c[\sigma]=c_{u}c_{v}c\left[\sigma\cdot X^{(11)}\bar{X}^{(11)}\right]\ ,$ (83)
which means $X^{(11)}\bar{X}^{(11)}\sim I$, and furthermore
$\sim[XX]X^{(11)}\bar{X}^{(11)}=X^{(10)+(01)}\bar{X}^{(11)}\ ,$ (84)
$[XX][\bar{X}\bar{X}]$ $\sim[XX]X^{(11)}[\bar{X}\bar{X}]\bar{X}^{(11)}=X^{(10)+(01)}\bar{X}^{(10)+(01)}\ ,$
$[XX]\bar{X}^{(11)}$ $\sim[XX]X^{(11)}\bar{X}^{(11)}\bar{X}^{(11)}=X^{(10)+(01)}\ .$
So we only have $7$ categories
$\mathsf{F}_{2,I},\mathsf{F}_{2,X}^{(01)},\mathsf{F}_{2,X}^{(11)},\mathsf{F}_{2,X\bar{X}}^{(01,01)},\mathsf{F}_{2,X\bar{X}}^{(01,10)},\mathsf{F}_{2,X\bar{X}}^{(01,11)},\mathsf{F}_{2,XX\bar{X}}\ .$ (85)
Their sizes are $\{18,72,18,36,36,72,72\}$.
Three Pairs: $C[\mathsf{F}]=3$,#$\#\mathsf{F}=36$
For $p=2$, (69) becomes
$\sum_{u_{1},u_{2},u_{3}}c_{u_{1}}c_{u_{2}}c_{u_{3}}\left[\prod_{j=1}^{m}(X+\bar{X})^{(\alpha_{j})}\right],m=0,1,2,3,\ldots$ (86)
For three pairs $u_{1},u_{2},u_{3}$, we can only take three conditions
$\alpha_{j}\in\{(011),(101),(110)\}\ .$ (87)
Any permutation $X_{e}X_{\bar{e}}$ is considered as a redundancy, allowing us to fix the order of $\bar{1},\bar{2},\bar{3}$. This means we only need to consider permutations of $1,2,3$. Utilizing the fact that $X=X^{(011)}+X^{(101)}+X^{(110)}$, we identify three building blocks: $I,X,[XX]$. The respective graph categories are:
$\mathsf{F}_{3,I}=\cup_{u_{1},u_{2},u_{3}}c[I],~{}\mathsf{F}_{3,X}=\cup_{u}c_{u_{1}}c_{u_{2}}c_{u_{3}}c[X],~{}\mathsf{F}_{3,XX}=\cup_{u_{1},u_{2},u_{3}}c_{u_{1}}c_{u_{2}}c_{u_{3}}c[[XX]]\ .$ (88)
They have sizes of $\{6,18,12\}$, respectively.
4.2 Operation of $\mathcal{L}_{3}$
Using the relations for $n=3$
$XX$ $=3I+3[XX],~{}[XX]X=2X,$ (89)
$XX_{e}$ $=X_{e}+[XX],~{}[XX]X_{e}=X_{e}[XX]=X-X_{e}\ ,$
and the gauge symmetry discussed earlier, it is straightforward to calculate the operation of $\mathcal{L}_{3}$ on the graph categories defined in the previous section. For simplicity, we define $\mathcal{L}_{n}^{\prime}$ as the non-trivial part of $\mathcal{L}_{n}$:
$\mathcal{L}_{n}\mathsf{F}_{p,\alpha}=(w+pJ)\mathsf{F}_{p,\alpha}+\mathcal{L}_{n}^{\prime}\mathsf{F}_{p,\alpha}\ .$ (90)
With some calculations one can get:
Zero Pair: $\mathsf{F}_{0,I},\mathsf{F}_{0,X},\mathsf{F}_{0,X\bar{X}},\mathsf{F}_{0,XX},\mathsf{F}_{0,XX\bar{X}},\mathsf{F}_{0,XX\bar{X}\bar{X}}$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{0,I}$ $={J\over D}\mathsf{F}_{1,I}-{J\over D}\mathsf{F}_{0,X}\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{0,X}$ $=\frac{J}{D}\left(\mathsf{F}^{(0)}_{1,X}+\mathsf{F}^{(1)}_{1,X}\right)-\frac{J}{D}\left(6\mathsf{F}_{0,I}+3\mathsf{F}_{0,XX}+2\mathsf{F}_{0,X\bar{X}}\right)\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{0,X\bar{X}}$ $=\frac{J}{D}\left(\mathsf{F}_{1,X\bar{X}}^{(00)}+\mathsf{F}_{1,X\bar{X}}^{(11)}+\mathsf{F}_{1,X\bar{X}}^{(01)}\right)-\frac{J}{D}\left(3\mathsf{F}_{0,X}+3\mathsf{F}_{0,XX\bar{X}}\right)\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{0,XX}$ $=\frac{J}{D}\mathsf{F}_{1,XX}-\frac{J}{D}\left(2\mathsf{F}_{0,X}+\mathsf{F}_{0,XX\bar{X}}\right)\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{0,XX\bar{X}}$ $=\frac{J}{D}\left(\mathsf{F}_{1,XX\bar{X}}^{(0)}+\mathsf{F}_{1,XX\bar{X}}^{(1)}\right)-\frac{J}{D}\left(4\mathsf{F}_{0,X\bar{X}}+3\mathsf{F}_{0,XX}+6\mathsf{F}_{0,XX\bar{X}\bar{X}}\right)\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{0,XX\bar{X}\bar{X}}$ $=\frac{J}{D}\mathsf{F}_{1,XX\bar{X}\bar{X}}-\frac{2J}{D}\mathsf{F}_{0,XX\bar{X}}\ .$ (91)
One Pair: $\mathsf{F}_{1,I},\mathsf{F}_{1,X}^{(0)},\mathsf{F}_{1,X}^{(1)},\mathsf{F}_{1,X\bar{X}}^{(00)},\mathsf{F}_{1,X\bar{X}}^{(01)},\mathsf{F}_{1,X\bar{X}}^{(11)},\mathsf{F}_{1,XX},\mathsf{F}_{1,XX\bar{X}}^{(0)},\mathsf{F}_{1,XX\bar{X}}^{(1)},\mathsf{F}_{1,XX\bar{X}\bar{X}}$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{1,I}$ $=\frac{2J}{D}\mathsf{F}_{2,I}-\frac{J}{D}\mathsf{F}_{1,X}^{(0)}\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{1,X}^{(0)}$ $=\frac{J}{D}\mathsf{F}_{2,X}^{(01)}-\frac{J}{D}\left(2\mathsf{F}_{1,I}+2\mathsf{F}_{1,X\bar{X}}^{(00)}\right)\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{1,X}^{(1)}$ $=\frac{J}{D}\left(\mathsf{F}_{2,X}^{(01)}+4\mathsf{F}_{2,X}^{(11)}\right)-\frac{J}{D}\left(\mathsf{F}_{1,XX}+\mathsf{F}_{1,X\bar{X}}^{(01)}\right)\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{1,XX}$ $=\frac{2J}{D}\mathsf{F}_{2,X\bar{X}}^{(01,11)}-\frac{J}{D}\left(\mathsf{F}_{1,X}^{(1)}+\mathsf{F}_{1,XX\bar{X}}^{(0)}\right)\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{1,X\bar{X}}^{(00)}$ $=\frac{J}{D}\mathsf{F}_{2,X\bar{X}}^{(01,01)}-\frac{J}{D}\mathsf{F}_{1,X}^{(0)}\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{1,X\bar{X}}^{(01)}$ $=\frac{J}{D}\left(2\mathsf{F}_{2,X\bar{X}}^{(01,10)}+\mathsf{F}_{2,X\bar{X}}^{(01,11)}\right)-\frac{J}{D}\left(\mathsf{F}_{1,X}^{(1)}+\mathsf{F}_{1,XX\bar{X}}^{(0)}\right)\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{1,X\bar{X}}^{(11)}$ $=\frac{J}{D}\left(\mathsf{F}_{2,X\bar{X}}^{(01,01)}+\mathsf{F}_{2,X\bar{X}}^{(01,11)}+2\mathsf{F}_{2,I}\right)-\frac{J}{D}\mathsf{F}_{1,XX\bar{X}}^{(1)}\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{1,XX\bar{X}}^{(0)}$ $=\frac{2J}{D}\mathsf{F}_{2,XX\bar{X}}-\frac{J}{D}\left(\mathsf{F}_{1,XX}+\mathsf{F}_{1,X\bar{X}}^{(01)}\right)\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{1,XX\bar{X}}^{(1)}$ $=\frac{J}{D}\left(\mathsf{F}_{2,XX\bar{X}}^{(01)}+2\mathsf{F}_{2,X+\bar{X}}^{(01)}\right)-\frac{2J}{D}\left(\mathsf{F}_{1,XX\bar{X}\bar{X}}+\mathsf{F}_{1,X\bar{X}}^{(11)}\right)\ ,$
$\mathcal{L}_{3}^{\prime}\mathsf{F}_{1,XX\bar{X}\bar{X}}$ $=\frac{J}{D}\left(\mathsf{F}_{2,X\bar{X}}^{(01;01)}+2\mathsf{F}_{2,X\bar{X}}^{(01;10)}\right)-\frac{J}{D}\mathsf{F}_{1,XX\bar{X}}^{(1)}\ .$ (92)
Two Pairs: $\mathsf{F}_{2,I},\mathsf{F}_{2,X}^{(01)},\mathsf{F}_{2,X}^{(11)},\mathsf{F}_{2,X\bar{X}}^{(01,01)},\mathsf{F}_{2,X\bar{X}}^{(01,10)},\mathsf{F}_{2,X\bar{X}}^{(01,11)},\mathsf{F}_{2,XX\bar{X}}$
$\mathcal{L}_{3}\mathsf{F}_{2,I}$ $=\left(w+2J\right)\mathsf{F}_{2,I}+\frac{3J}{D}\mathsf{F}_{3,I}\ ,$
$\mathcal{L}_{3}\mathsf{F}_{2,X}^{(01)}$ $=\left(w+2J\right)\mathsf{F}_{2,X}^{(01)}+\frac{6J}{D}\mathsf{F}_{3,X}\ ,$
$\mathcal{L}_{3}\mathsf{F}_{2,X}^{(11)}$ $=\left(w+2J\right)\mathsf{F}_{2,X}^{(11)}+\frac{2J}{D}\mathsf{F}_{3,X}\ ,$
$\mathcal{L}_{3}\mathsf{F}_{2,X\bar{X}}^{(01,01)}$ $=\left(w+2J\right)\mathsf{F}_{2,X\bar{X}}^{(01,01)}+\frac{3J}{D}\mathsf{F}_{3,I}\ ,$
$\mathcal{L}_{3}\mathsf{F}_{2,X\bar{X}}^{(01,10)}$ $=\left(w+2J\right)\mathsf{F}_{2,X\bar{X}}^{(01,01)}+\frac{3J}{D}\mathsf{F}_{3,XX}\ ,$
$\mathcal{L}_{3}\mathsf{F}_{2,X\bar{X}}^{(01,11)}$ $=\left(w+2J\right)\mathsf{F}_{2,X\bar{X}}^{(01,11)}+\frac{3J}{D}\mathsf{F}_{3,XX}\ ,$
$\mathcal{L}_{3}\mathsf{F}_{2,XX\bar{X}}^{(01)}$ $=\left(w+2J\right)\mathsf{F}_{2,XX\bar{X}}^{(01)}+\frac{6J}{D}\mathsf{F}_{3,X}\ .$ (93)
Three Pairs: $\mathsf{F}_{3,I},\mathsf{F}_{3,X},\mathsf{F}_{3,XX}$
$\mathcal{L}_{3}\mathsf{F}_{3,\beta}=(w+3J)\mathsf{F}_{3,\beta},\quad\beta=I,X,XX\ .$ (94)
4.3 Observables
Recall the factor $\mathsf{w}=-\text{i}\mathsf{E}-nJ$ from (13), then we have $M=-\text{i}\mathsf{E}\mathbf{1}+M_{J}$. Here $\mathbf{1}$ denotes identity matrix. $M_{J}$ has the explicit form
$\left(\begin{array}[]{cccccccccccccccccccccccccc}-3J&-\frac{6J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ -\frac{J}{D}&-3J&-\frac{3J}{D}&-\frac{2J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&-\frac{2J}{D}&-3J&0&-\frac{4J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&-\frac{3J}{D}&0&-3J&-\frac{3J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&-\frac{3J}{D}&-\frac{J}{D}&-3J&-\frac{2J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&-\frac{6J}{D}&-3J&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ \frac{J}{D}&0&0&0&0&0&-2J&-\frac{2J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&\frac{J}{D}&0&0&0&0&-\frac{J}{D}&-2J&0&-\frac{J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&\frac{J}{D}&0&0&0&0&0&0&-2J&0&-\frac{J}{D}&0&-\frac{J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&\frac{J}{D}&0&0&0&0&-\frac{2J}{D}&0&-2J&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&\frac{J}{D}&0&0&0&0&0&-\frac{J}{D}&0&-2J&0&0&-\frac{J}{D}&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&\frac{J}{D}&0&0&0&0&0&0&0&0&-2J&0&0&-\frac{2J}{D}&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&\frac{J}{D}&0&0&0&0&-\frac{J}{D}&0&0&0&-2J&-\frac{J}{D}&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&\frac{J}{D}&0&0&0&0&0&-\frac{J}{D}&0&-\frac{J}{D}&-2J&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&\frac{J}{D}&0&0&0&0&0&0&-\frac{J}{D}&0&0&-2J&-\frac{J}{D}&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&\frac{J}{D}&0&0&0&0&0&0&0&0&-\frac{2J}{D}&-2J&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&\frac{2J}{D}&0&0&0&0&\frac{2J}{D}&0&0&0&0&-J&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&\frac{J}{D}&\frac{J}{D}&0&0&0&0&0&\frac{2J}{D}&0&0&-J&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&\frac{4J}{D}&0&0&0&0&0&0&0&0&0&-J&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&\frac{J}{D}&0&\frac{J}{D}&0&0&0&\frac{2J}{D}&0&0&0&-J&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&\frac{2J}{D}&0&0&0&0&\frac{2J}{D}&0&0&0&0&-J&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&\frac{J}{D}&\frac{J}{D}&\frac{2J}{D}&0&0&0&0&0&0&0&0&-J&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&\frac{2J}{D}&\frac{2J}{D}&0&0&0&0&0&0&0&-J&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&\frac{3J}{D}&0&0&\frac{6J}{D}&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&\frac{4J}{D}&\frac{J}{D}&0&0&0&\frac{4J}{D}&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&\frac{3J}{D}&\frac{6J}{D}&0&0&0&0\\ \end{array}\right)\ .$ (121)
It is straightforward to obtain the spectrum of $M_{J}$:
[…]
$-3J,-3J,-2J,-2J,-2J,-2J,-J,-J,-J,-J,-J,-J,-J,$
$-\frac{3(D+1)J}{D},-\frac{2(D+1)J}{D},-\frac{2(D+1)J}{D},-\frac{2(D+1)J}{D},-\frac{3(D+2)J}{D}\Bigg{\}}\ .$
Then we can find the analytical expression of $\mathcal{U}_{3}(t)$ as
$\mathcal{U}_{3}(t)=e^{\mathcal{L}_{3}t}=e^{-\text{i}\mathsf{E}t}\sum_{a=1}^{26}f_{a}(t)\mathsf{F}_{a}\ .$ (123)
We plot $f_{a}(t)$ in Fig.3, while the explicit expressions are given in Appendix A.
Figure 3: The pattern coefficient $f_{a}(t)$ and its log-plot, where we set $D=10$ and $J=1$. We depict the categories of $p=0,1,2,3$ with red, orange, green, and blue colors, respectively. One can see that the three categories with $p=3$ have energy 00, so they (the blue lines) remain finite at time $t=\infty$.
Generally, we can calculate the correlation functions as
$\text{Tr}\left[O_{1}(t)O_{2}O_{3}(t)O_{4}O_{5}(t)O_{6}\right]=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\equiv\sum_{a=1}^{26}f_{a}(t)\times\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\right)_{O_{1},O_{2},O_{3},O_{4},O_{5},O_{6}}$ (124)
where we have used $O(t)\equiv U^{\dagger}OU$ and the graph representation for an operator $A$
$A_{ij}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\ .$ (125)
To calculate $\mathsf{F}_{a}[O_{1},O_{2},O_{3},O_{4},O_{5},O_{6}]$, in principle, we need to contract each individual graph in the category with the operator tensor in the way as shown in (124). The expression $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$ has told us the connection between points in the graph. We treat the connections in the graph as free propagators and the operators as bridges between points $i,\bar{i}$ and $\bar{i}^{\prime},\overline{(i+1)}^{\prime}$, where the prime means that the point is on the right-hand side of the graph. So there is no need to draw the graph. We can just start with any points following the connections in the graph and form a closed loop. Hence we will have a trace of operators
$\text{Tr}(O_{i}O_{j}O^{T}_{k}\cdots)\ .$ (126)
If there are still points that we have not passed, we then start with those points, finding another closed loop until all points in the graph have been visited. The process may be a bit tedious, but it can be accomplished with some code. Alternatively, one can simply write down the explicit expression of a graph using the Kronecker delta function as defined in (16).
Moreover, with the expression of $\mathcal{U}_{1,2,3}(t)$, one can calculate the correlation function with general time order $\text{Tr}\left[O_{1}(t_{1})O_{2}O_{3}(t_{2})O_{4}O_{5}(t_{3})O_{6}\right]$. For example, we consider the case $t_{1}<t_{2}<t_{3}$, then the ensemble-averaged time-evolution $\mathcal{U}_{3}(t_{1},t_{2},t_{3})$ is
$\mathbb{E}\left(\mathbb{I}_{D}^{\otimes 2}\otimes\mathbb{I}_{D}^{\otimes 2}\otimes U(t_{32})\otimes U^{*}(t_{32})\right)\ \mathbb{E}\left(\mathbb{I}_{D}\otimes\mathbb{I}_{D}\otimes U(t_{21})^{\otimes 2}\otimes U^{*}(t_{21})^{\otimes 2}\right)\ \mathbb{E}\left(U(t_{1})^{\otimes 3}\otimes U^{*}(t_{1})^{\otimes 3}\right)$
$=$ $\left[\mathbb{I}_{D}^{\otimes 2}\otimes\mathbb{I}_{D}^{\otimes 2}\otimes\mathcal{U}_{1}(t_{32})\right]\times\left[\mathbb{I}_{D}\otimes\mathbb{I}_{D}\otimes\mathcal{U}_{2}(t_{21})\right]\times\mathcal{U}_{3}(t_{1})\ ,$ (127)
where $\mathbb{I}_{D}$ denotes the $D$-dimensional identity matrix. Finally, we simply need to contract $\mathcal{U}_{3}(t_{1},t_{2},t_{3})$ with the operator tensor as demonstrated in (124). In this paper, we do not provide explicit expressions for the correlation functions.
5 General Cases
For general $n$, we have $2n$ sites $\{1,\bar{1},2,\bar{2},3,\bar{3},\ldots,n,\bar{n}\}$. The permutation group is $S_{n}\otimes S_{n}$. As discussed in $n=3$ case, we should define
$X^{(\boldsymbol{\alpha}_{j})}\equiv\left\{X_{e}|\cap_{k=1}^{p}\text{Con}(\boldsymbol{\alpha}_{j}[k],e,u_{k})\right\}\ .$ (128)
$X^{(\boldsymbol{\alpha}_{j})}$ is a set of 2-swap operators submitting to the condition $\boldsymbol{\alpha}_{j}$. Our primary goal is to deal with the following graph expression
$\sum_{u_{1},u_{2},\ldots,u_{p}}c(u_{1},u_{2},\ldots,u_{p})c\left[\prod_{j=1}^{m}(X+\bar{X})^{(\boldsymbol{\alpha}_{j})}\right]\ ,$ (129)
and assign them into appropriate categories.
Since it is free to change the order of $u_{1},u_{2},\ldots,u_{p}$, there is a $S_{p}$ permutation redundancy for the conditions:
$\{\boldsymbol{\alpha}_{j}\}_{j=1,2,\ldots,p}\sim\{\sigma\boldsymbol{\alpha}_{j}\}_{j=1,2,\ldots,p},~{}\forall\sigma\in S_{p}\ .$ (130)
In practice, we can always find a permutation $\sigma\in S_{n}$ as a product of cycles $\{\sigma_{i}\}_{i=1,2,\ldots,g}$ with length $1<l_{1}\leq l_{2}\ldots\leq l_{g}$ such that:
$X_{[l_{1},l_{2},\ldots,l_{g}]}\equiv\left\{X_{\sigma_{1}\sigma_{2}\cdots\sigma_{g}}||\sigma_{i}|=l_{i},\sigma_{i}\cap\sigma_{j}=0,i\not=j\right\}\ .$ (131)
Then for a certain $p$, we first analyze all permutation categories appearing in the products $\prod_{j=1}^{m}X^{(\alpha_{j})},m=0,1,2,\ldots$, which gives us a division of the permutation group $S_{n}$
$S_{n}=\cup_{\alpha=1}^{N_{X}}X_{\alpha},X_{\alpha}\cap X_{\beta}=\emptyset,\forall\alpha\not=\beta\ .$ (132)
So naively we have ${N_{X}(N_{X}+1)\over 2}$ graph categories:
$\mathsf{F}_{p,X_{\alpha}\bar{X}_{\beta}}\equiv\cup_{x\in X_{\alpha}\bar{X}_{\beta}\cup X_{\beta}\bar{X}_{\alpha}}\cup_{u_{1},u_{2},\ldots,u_{p}}c(u_{1},u_{2},\ldots,u_{p})c[x]\ .$ (133)
Finally when $p\geq 2$, we need to consider removing the permutation redundancy of the conditions and the gauge symmetry of the graph representation. In other words, we need to module some equivalent relations.
5.1 Example: $n=4$
Here we take $n=4$ as an example. For different number of connected pairs, the number of graphs should be $N_{4,p}={\{576,9216,20736,9216,576\}}$. We then define
$X_{[2]}\equiv X=\{X_{12},X_{13},X_{14},X_{23},X_{24},X_{34}\}\ .$ (134)
Like what we do for $n=3$ case, we discuss the graph categories according to the number of pairs.
Zero Pair $p=0,\#\mathsf{F}=576$
For $p=0$, there is no restriction on $X$, so we have the division of $S_{4}$ as
$S_{4}=I+X_{[2]}+X_{[3]}+X_{[2,2]}+X_{[4]},~{}\#S_{4}=1+6+8+3+6=24\ .$ (135)
Since there is no permutation redundancy or gauge redundancy, we have ${N_{X}(N_{X}+1)\over 2}=15$ independent graph categories:
$\mathsf{F}_{0,I},\mathsf{F}_{0,X},\mathsf{F}_{0,X_{[3]}},\mathsf{F}_{0,X_{[2,2]}},\mathsf{F}_{0,X_{[4]}},\mathsf{F}_{0,X\bar{X}},\mathsf{F}_{0,X\bar{X}_{[3]}},\mathsf{F}_{0,X\bar{X}_{[2,2]}},\mathsf{F}_{0,X\bar{X}_{[4]}},\mathsf{F}_{0,X_{[3]}\bar{X}_{[3]}},$
$\mathsf{F}_{0,X_{[3]}\bar{X}_{[2,2]}},\mathsf{F}_{0,X_{[3]}\bar{X}_{[4]}},\mathsf{F}_{0,X_{[2,2]}\bar{X}_{[2,2]}},\mathsf{F}_{0,X_{[2,2]}\bar{X}_{[4]}},\mathsf{F}_{0,X_{[4]}\bar{X}_{[4]}}\ .$ (136)
One Pair $p=1,\#\mathsf{F}=9216$
For an arbitrary pair $u$, we have two conditions for $X$ and $X_{[3]}$, so that we have the division
$S_{4}=I+X^{(0)}+X^{(1)}+X_{[3]}^{(0)}+X_{[3]}^{(1)}+X_{[2,2]}+X_{[4]}\equiv\cup_{\alpha=1}^{7}X_{\alpha}\ .$ (137)
Since $p<2$, there is no permutation or gauge redundancy, so we have ${7\times 8\over 2}=28$ categories:
$F_{1,X_{\alpha}\bar{X}_{\beta}}\equiv\sum_{u}\sum_{x\in X_{\alpha}\bar{X}_{\beta}\cup X_{\beta}\bar{X}_{\alpha}}c_{u}c[x].$ (138)
Two Pairs $p=2,\#\mathsf{F}=20736$
For any two pairs $u_{1},u_{2}$, we find
$X=X^{(00)}+X^{(01)}+X^{(10)}+X^{(11)},~{}X_{[3]}^{(00)}=\emptyset,~{}X_{[3]}=X_{[3]}^{(01)}+X^{(10)}_{[3]}+X_{[3]}^{(11)}\ ,$ (139)
so we obtain the division to be
$S_{4}=I+X^{(00)}+X^{(01)}+X^{(10)}+X^{(11)}+X_{[3]}^{(01)}+X_{[3]}^{(10)}+X_{[3]}^{(11)}+X_{[2,2]}+X_{[4]}\ .$ (140)
Naively, we have ${10\times 11\over 2}=55$ categories. But here we need to remove the permutation redundancy. For an example, changing the order of conditions $\alpha_{j}$, we have
$\mathsf{F}_{2,X}^{(01)}=\mathsf{F}_{2,X}^{(10)},~{}\mathsf{F}_{2,X\bar{X}}^{(00,01)}=\mathsf{F}_{2,X\bar{X}}^{(00,10)}\ .$ (141)
And notice the symmetry between $X$ and $\bar{X}$, we have
$\mathsf{F}_{2,X\bar{X}}^{(00,01)}=\mathsf{F}_{2,X\bar{X}}^{(01,00)}\ .$ (142)
As for the gauge redundancy, here we impose the equivalence $X^{(11)}\bar{X}^{(11)}\sim I$, so immediately we have
$\mathsf{F}_{2,X\bar{X}}^{(11,11)}=\mathsf{F}_{2,I}\ .$ (143)
There are other equivalence between categories, we will not shown them here. After removing all redundancy, we will obtain the proper graph categories for $p=2$.
Three pairs
For three pairs, we have
$X$ $=X^{(001)}+X^{(010)}+X^{(100)}+X^{(011)}+X^{(101)}+X^{(110)},$
$X_{[3]}$ $=X_{[3]}^{(011)}+X_{[3]}^{(101)}+X_{[3]}^{(110)}\ ,$ (144)
so we obtain the division as
$S_{4}$ $=I+X^{(001)}+X^{(010)}+X^{(100)}+X^{(011)}+X^{(101)}+X^{(110)}$
$\hskip 15.0pt+X_{[3]}^{(011)}+X_{[3]}^{(101)}+X_{[3]}^{(110)}+X_{[2,2]}+X_{[4]}\ .$ (145)
Naively, there are ${12\times 13\over 2}=96$ categories. The permutation redundancy is similar to $p=2$ case and the gauge group is generated by
$X^{(011)}\bar{X}^{(011)},X^{(101)}\bar{X}^{(101)},X^{(110)}\bar{X}^{(110)}\ .$ (146)
After removing the all redundancy, the number of categories will be much smaller.
Four Pairs
We notice that
$X$ $=X^{(0011)}+X^{(0101)}+X^{(1001)}+X^{(0110)}+X^{(1010)}+X^{(1100)}\ ,$
$X_{[3]}$ $=X_{[3]}^{(0111)}+X_{[3]}^{(1011)}+X_{[3]}^{(1101)}+X_{[3]}^{(1110)}\ ,$ (147)
which gives the division as
$S_{4}$ $=I+X^{(0011)}+X^{(0101)}+X^{(1001)}+X^{(0110)}+X^{(1010)}+X^{(1100)}$
$\hskip 15.0pt+X_{[3]}^{(011)}+X_{[3]}^{(101)}+X_{[3]}^{(110)}+X_{[2,2]}+X_{[4]}\ .$ (148)
Notice the gauge group is generated by each permutation like $X_{e}\bar{X}_{e}$. So we can fix the order of $\bar{i}$ to be $\bar{1}\bar{2}\bar{3}\bar{4}$, and we just need to consider the categories
$\mathsf{F}_{\alpha}\equiv\cup_{u_{1},u_{2},u_{3},u_{4}}c(u_{1},u_{2},u_{3},u_{4})c[X_{\alpha}]\ .$ (149)
Notice the permutation redundancy, we have
$X^{(0011)}\sim X^{(0101)}\sim X^{(1001)}\sim X^{(0110)}\sim X^{(1010)}\sim X^{(1100)},X_{[3]}^{(011)}\sim X_{[3]}^{(101)}\sim X_{[3]}^{(110)}\ ,$ (150)
so we only have five categories left
$\mathsf{F}_{4,I},\mathsf{F}_{4,X},\mathsf{F}_{4,X_{[3]}},\mathsf{F}_{4,X_{[2,2]}},\mathsf{F}_{4,X_{[4]}}\ .$ (151)
6 Discussions and Outlook
In this paper we analyze the time evolution of the $n$-replica physics of the BGUE model in a graph-based framework. This approach proves to be particularly effective for lower values of $n$, as explicitly demonstrated in the $n=2$ and $n=3$ cases. We also provide a systematic strategy that extends to generic $n$. Our construction highlights how the increasing complexity of graph structures can be managed by organizing them into well-defined categories, which will simplify the calculations of two-point functions, OTOC, and other relevant observables.
The general formulation outlined in Sec. 5 suggests that this method can, in principle, be extended to arbitrary $n$. However, as $n$ increases, the growing number of graph structures presents significant challenges. Addressing these challenges will require the development of more efficient techniques for classifying, encoding, and manipulating graphs at large $n$. An important direction for future research is to explore algorithmic approaches to systematically handle this complexity. Moreover, further investigation into the implications of this framework for Brownian disordered systems and quantum information theory may uncover deeper structural insights into the role of replica dynamics in random matrix models.
A key insight is that the factor $w$, which encodes the dependence on the system spectrum, commutes with the graph-based operators. This allows us to analyze the system with infinite-temperature GUE noise without introducing additional complexity.
It is worth noting that the techniques developed in this paper can also be applied to other models, such as the Brownian Gaussian Orthogonal Ensemble (BGOE) and the Brownian Gaussian Symplectic Ensemble (BGSE). In particular, for BGOE, we have:
$\mathcal{L}_{n}=\frac{-JD}{D+1}\left[n\mathbb{I}+\sum_{i<j}\left(P_{ij}+P_{\bar{i}\bar{j}}\right)-\sum_{i,\bar{j}}P_{i\bar{j}}\right]+\frac{-J}{D+1}\left[n+\sum_{i<j}X_{ij}+X_{\bar{i}\bar{j}}-\sum_{i\bar{j}}X_{i\bar{j}}\right]\ ,$ (152)
and for BGSE:
$\mathcal{L}_{n}=\frac{-J}{D-1}\left[2n(D-1)\mathbb{I}-D\sum_{i<j}\left(P_{ij}+P_{\bar{i}\bar{j}}\right)-2D\sum_{i,\bar{j}}P_{i\bar{j}}+4\sum_{i<j}\left(X_{ij}+X_{\bar{i}\bar{j}}\right)+2\sum_{i,\bar{j}}X_{i\bar{j}}\right]\ .$ (153)
We find there are new operators such as $P_{ij},P_{\bar{i}\bar{j}},X_{i\bar{j}}$. Hence an individual graph appearing in $\mathcal{L}_{n}^{r}$ can be written as
$\mathsf{F}=\prod_{i=1}^{p_{1}}c_{a_{i}a^{\prime}_{i}}\prod_{i=1}^{p_{2}}c_{\bar{b}_{i}\bar{b}^{\prime}_{i}}\prod_{i=1}^{p_{3}}c_{r_{i}\bar{s}_{i}}c\left[\sigma\right]\ .$ (154)
As a result, when $n$ increases, the introduction of new operators in BGOE and BGSE further complicates the graph structures, leading to a greater number of individual graphs that become increasingly challenging to systematically organize them into appropriate categories. While this classification seems to be complicated, it is, at its core, a well-structured problem. As discussed in the introduction, the method presented in this paper is fundamentally equivalent to the approach outlined in Guo:2024zmr . Moreover, the inherent symmetry of $\mathcal{L}_{n}$ in the BGUE, BGOE, and BGSE cases simplifies its diagonalization, which is closely related to a well-studied mathematical problem: the decomposition of the direct product of fundamental representations. Given this, we expect that the techniques developed in this work can also be successfully extended to models without obvious symmetry, further demonstrating the power and versatility of the proposed framework.
Acknowledgment
The authors would like to thank Haifeng Tang, Yingyu Yang, and Yanyuan Li for their valuable discussions and comments during the preparation of this work. CP, JY and TL are supported by NSFC NO. 12175237, CP is also supported in part by NSFC NO. 12447108, NSFC NO. 12247103, the Fundamental Research Funds for the Central Universities, and funds from the Chinese Academy of Sciences.
Appendix A Expression of $\mathcal{U}_{3}(t)$
As discussed in Sec.4, we write $\mathcal{U}_{3}=\sum_{a=1}^{26}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}$. Here we list the explicit expressions for $f_{a}$. For simplicity, we denote
$f_{a}(t)=\boldsymbol{v}_{a}\cdot\boldsymbol{\xi}(t)\ ,$ (155)
where $\boldsymbol{\xi}(t)$ are $10$ wavefunctions.
$\boldsymbol{\xi}(t)=\left\{e^{-Jt},e^{-2Jt},e^{-\frac{2(D+1)Jt}{D}},e^{-\frac{2(D-1)Jt}{D}},e^{-3Jt},e^{-\frac{3(D+1)Jt}{D}},e^{-\frac{3(D-1)Jt}{D}},e^{-\frac{3(D+2)Jt}{D}},e^{-\frac{3(D-2)Jt}{D}},1\right\}\ ,$ (156)
and the $26$ vectors are
$\boldsymbol{v}_{1}$ =$=\left\{0,0,0,0,\frac{1}{2},\frac{2}{9},\frac{2}{9},\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{2}$ =$=\left\{0,0,0,0,0,\frac{1}{9},-\frac{1}{9},\frac{1}{36},-\frac{1}{36},0\right\},$
$\boldsymbol{v}_{3}$ =$=\left\{0,0,0,0,-\frac{1}{18},0,0,\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{4}$ =$=\left\{0,0,0,0,-\frac{1}{6},\frac{1}{18},\frac{1}{18},\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{5}$ =$=\left\{0,0,0,0,0,-\frac{1}{18},\frac{1}{18},\frac{1}{36},-\frac{1}{36},0\right\},$
$\boldsymbol{v}_{6}$ =$=\left\{0,0,0,0,\frac{1}{6},-\frac{1}{9},-\frac{1}{9},\frac{1}{36},\frac{1}{36},0\right\},$
$\boldsymbol{v}_{7}$ =$=\Bigg{\{}0,\frac{5-D^{2}}{18D-2D^{3}},\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{D^{2}-3D-2}{4(D-4)(D-1)(D+2)},$
$\hskip 15.0pt-\frac{20-9D^{2}}{72D-18D^{3}},-\frac{2(D+2)}{9(D+1)(D+3)},-\frac{2(D-2)}{9\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg{\}},$
$\boldsymbol{v}_{8}$ =$=\Bigg{\{}0,0,\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{-D^{2}+3D+2}{4\left(D^{3}-3D^{2}-6D+8\right)},\frac{4}{36-9D^{2}},$
$\hskip 15.0pt-\frac{1}{9D+9},\frac{1}{9(D-1)},-\frac{1}{36(D+4)},\frac{1}{36(D-4)},0\Bigg{\}},$
$\boldsymbol{v}_{9}$ =$=\Bigg{\{}0,\frac{1}{18-2D^{2}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},-\frac{D-2}{4(D-4)(D-1)(D+2)},$
$\hskip 15.0pt\frac{2}{9\left(D^{2}-4\right)},-\frac{2D+3}{18(D+1)(D+3)},\frac{2D-3}{18(D-3)(D-1)},-\frac{1}{36(D+4)},\frac{1}{36(D-4)},0\Bigg{\}},$
$\boldsymbol{v}_{10}$ =$=\Bigg{\{}0,-\frac{D^{2}-5}{2D\left(D^{2}-9\right)},\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{D^{2}-3D-2}{4(D-4)(D-1)(D+2)},\frac{20-D^{2}}{72D-18D^{3}},$
$\hskip 15.0pt-\frac{2}{9(D+1)(D+3)},\frac{2}{9\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg{\}},$
$\boldsymbol{v}_{11}$ =$=\Bigg{\{}0,\frac{1}{18D-2D^{3}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D-2}{4(D-4)(D-1)(D+2)},-\frac{D^{2}+4}{72D-18D^{3}},$
$\hskip 15.0pt-\frac{1}{18(D+1)(D+3)},\frac{1}{18\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg{\}},$
$\boldsymbol{v}_{12}$ =$=\Bigg{\{}0,\frac{1}{D\left(D^{2}-9\right)},\frac{1}{2(D-2)(D+1)(D+4)},\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{8-D^{2}}{72D-18D^{3}},$
$\hskip 15.0pt\frac{1}{9\left(D^{2}+4D+3\right)},-\frac{1}{9\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg{\}},$
$\boldsymbol{v}_{13}$ =$=\Bigg{\{}0,-\frac{1}{18D-2D^{3}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D-2}{4(D-4)(D-1)(D+2)},\frac{4-3D^{2}}{72D-18D^{3}},$
$\hskip 15.0pt-\frac{D+2}{18(D+1)(D+3)},-\frac{D-2}{18\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg{\}},$
$\boldsymbol{v}_{14}$ =$=\Bigg{\{}0,\frac{1}{2\left(D^{2}-9\right)},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},-\frac{D-2}{4(D-4)(D-1)(D+2)},\frac{2}{9\left(D^{2}-4\right)},$
$\hskip 15.0pt\frac{D}{18\left(D^{2}+4D+3\right)},-\frac{D}{18\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{36(D-4)},0\Bigg{\}},$
$\boldsymbol{v}_{15}$ =$=\Bigg{\{}0,0,\frac{1}{2(D-2)(D+1)(D+4)},-\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{1}{36-9D^{2}},\frac{1}{18D+18},$
$\hskip 15.0pt\frac{1}{18-18D},-\frac{1}{36(D+4)},\frac{1}{36(D-4)},0\Bigg{\}},$
$\boldsymbol{v}_{16}$ =$=\Bigg{\{}0,\frac{1}{9D-D^{3}},\frac{1}{2(D-2)(D+1)(D+4)},\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},-\frac{8-3D^{2}}{72D-18D^{3}},$
$\hskip 15.0pt\frac{D+2}{9\left(D^{2}+4D+3\right)},\frac{D-2}{9(D-3)(D-1)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg{\}},$
$\boldsymbol{v}_{17}$ =$=\Bigg{\{}\frac{D^{4}-8D^{2}+6}{D^{6}-13D^{4}+36D^{2}},\frac{3-D^{2}}{D^{2}\left(D^{2}-9\right)},-\frac{D(D+3)}{2\left(D^{4}+5D^{3}-20D-16\right)},-\frac{(D-3)D}{2\left(D^{4}-5D^{3}+20D-16\right)},$
$\hskip 15.0pt\frac{6-4D^{2}}{36D^{2}-9D^{4}},\frac{2}{9(D+1)(D+3)},\frac{2}{9\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},\frac{1}{18\left(D^{2}-7D+12\right)},0\Bigg{\}},$
$\boldsymbol{v}_{18}$ =$=\Bigg{\{}\frac{1}{9D-D^{3}},-\frac{1}{18D-2D^{3}},-\frac{D}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D}{4\left(D^{3}-3D^{2}-6D+8\right)},-\frac{2}{36D-9D^{3}},$
$\hskip 15.0pt\frac{1}{18\left(D^{2}+4D+3\right)},-\frac{1}{18\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},-\frac{1}{18(D-4)(D-3)},0\Bigg{\}},$
$\boldsymbol{v}_{19}$ =$=\Bigg{\{}\frac{1}{9D-D^{3}},\frac{2}{D\left(D^{2}-9\right)},\frac{1}{(D-2)(D+1)(D+4)},\frac{1}{D^{3}-3D^{2}-6D+8},\frac{4}{36D-9D^{3}},$
$\hskip 15.0pt\frac{2}{9(D+1)(D+3)},-\frac{2}{9\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},-\frac{1}{18(D-4)(D-3)},0\Bigg{\}},$
$\boldsymbol{v}_{20}$ =$=\Bigg{\{}\frac{2D^{2}-3}{D^{2}\left(D^{4}-13D^{2}+36\right)},\frac{D^{2}-3}{2D^{2}\left(D^{2}-9\right)},\frac{D^{2}+3D+4}{-4D^{4}-20D^{3}+80D+64},$
$\hskip 15.0pt\frac{D^{2}-3D+4}{-4D^{4}+20D^{3}-80D+64},\frac{D^{2}+3}{9D^{2}\left(D^{2}-4\right)},-\frac{1}{9(D+1)(D+3)},-\frac{1}{9\left(D^{2}-4D+3\right)},$
$\hskip 15.0pt\frac{1}{18(D+3)(D+4)},\frac{1}{18\left(D^{2}-7D+12\right)},0\Bigg{\}},$
$\boldsymbol{v}_{21}$ =$=\Bigg{\{}\frac{D^{2}+6}{D^{6}-13D^{4}+36D^{2}},\frac{3}{D^{2}\left(D^{2}-9\right)},\frac{D}{2\left(D^{4}+5D^{3}-20D-16\right)},\frac{D}{-2D^{4}+10D^{3}-40D+32},$
$\hskip 15.0pt\frac{D^{2}-6}{9D^{2}\left(D^{2}-4\right)},-\frac{1}{9(D+1)(D+3)},-\frac{1}{9\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},\frac{1}{18\left(D^{2}-7D+12\right)},0\Bigg{\}},$
$\boldsymbol{v}_{22}$ =$=\Bigg{\{}\frac{2D^{2}-3}{D^{2}\left(D^{4}-13D^{2}+36\right)},\frac{3}{18D^{2}-2D^{4}},\frac{3D+4}{4\left(D^{4}+5D^{3}-20D-16\right)},\frac{4-3D}{4\left(D^{4}-5D^{3}+20D-16\right)},$
$\hskip 15.0pt\frac{3-2D^{2}}{9D^{2}\left(D^{2}-4\right)},\frac{1}{18\left(D^{2}+4D+3\right)},\frac{1}{18\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},\frac{1}{18\left(D^{2}-7D+12\right)},0\Bigg{\}},$
$\boldsymbol{v}_{23}$ =$=\Bigg{\{}-\frac{5}{D^{5}-13D^{3}+36D},\frac{1}{9D-D^{3}},\frac{D}{2\left(D^{4}+5D^{3}-20D-16\right)},\frac{D}{2\left(D^{4}-5D^{3}+20D-16\right)},$
$\hskip 15.0pt\frac{1}{36D-9D^{3}},-\frac{1}{9(D+1)(D+3)},\frac{1}{9\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},-\frac{1}{18(D-4)(D-3)},0\Bigg{\}},$
$\boldsymbol{v}_{24}$ =$=\Bigg{\{}\frac{3}{9D-D^{3}},0,\frac{3}{2(D-2)(D+1)(D+4)},\frac{3}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{2}{12D-3D^{3}},0,0,$
$\hskip 15.0pt-\frac{1}{6(D+2)(D+3)(D+4)},-\frac{1}{6\left(D^{3}-9D^{2}+26D-24\right)},\frac{2-D^{2}}{D^{5}-5D^{3}+4D}\Bigg{\}},$
$\boldsymbol{v}_{25}$ =$=\Bigg{\{}\frac{5}{D^{4}-13D^{2}+36},0,\frac{1}{2\left(D^{3}+7D^{2}+14D+8\right)},\frac{1}{-2D^{3}+14D^{2}-28D+16},0,0,0,$
$\hskip 15.0pt-\frac{1}{6(D+2)(D+3)(D+4)},\frac{1}{6(D-4)(D-3)(D-2)},\frac{1}{D^{4}-5D^{2}+4}\Bigg{\}},$
$\boldsymbol{v}_{26}$ =$=\Bigg{\{}-\frac{15}{D^{5}-13D^{3}+36D},0,-\frac{3}{D^{4}+5D^{3}-20D-16},\frac{3}{D^{4}-5D^{3}+20D-16},-\frac{1}{12D-3D^{3}},$
$\hskip 15.0pt0,0,-\frac{1}{6(D+2)(D+3)(D+4)},-\frac{1}{6\left(D^{3}-9D^{2}+26D-24\right)},-\frac{2}{D^{5}-5D^{3}+4D}\Bigg{\}}\ .$

### v2

@@
-Graph-Theoretic Analysis of $n$-Replica Time Evolution in the Brownian Gaussian Unitary Ensemble
+
 Abstract
-In this paper, we investigate the $n$-replica time evolution operator $\mathcal{U}_{n}(t)\equiv e^{\mathcal{L}_{n}t}$ for the Brownian Gaussian Unitary Ensemble (BGUE) using a graph-theoretic approach.
-We examine the moments of the generating operator $\mathcal{L}_{n}$, which governs the Euclidean time evolution within an auxiliary $D^{2n}$-dimensional Hilbert space, where $D$ represents the dimension of the Hilbert space for the original system.
-Explicit representations for the cases of $n=2$ and $n=3$ are derived, emphasizing the role of graph categorization in simplifying calculations.
-Furthermore, we present a general approach to streamline the calculation of time evolution for arbitrary $n$, supported by a detailed example of $n=4$.
-Our results demonstrate that the $n$-replica framework not only facilitates the evaluation of various observables but also provides valuable insights into the relationship between Brownian disordered systems and quantum information theory.
-preprint: USTC-ICTS/PCFT-25-05
+In this paper, we investigate the $n$-replica time evolution operator $\mathcal{U}_{n}(t)\equiv e^{\mathcal{L}_{n}t}$ for the Brownian Gaussian Unitary Ensemble (BGUE) through a graph-theoretic approach.
+We analyze the moments of the generator operator $\mathcal{L}_{n}$, which governs the Euclidean time evolution within an auxiliary $D^{2n}$-dimensional Hilbert space, where $D$ represents the dimension of the Hilbert space for the original system.
+Explicit representations for the cases of $n=2$ and $n=3$ are derived, highlighting the significance of graph categorization in simplifying complex calculations.
+Furthermore, we discuss a general formulation applicable to arbitrary $n$ and provide detailed examples, including the case of $n=4$.
+Our findings reveal that the $n$-replica framework not only facilitates the evaluation of various observables but also offers insights into the interplay between Brownian disordered systems and quantum information theory.
+Keywords:
+Brownian, GUE, Graph Representation
 1 Introduction
-Brownian disordered quantum systems have garnered significant interest across various fields, including many-body quantum chaos Chen_2019 ; S_nderhauf_2019 ; kalsi2024hierarchicalanalyticalapproachuniversal , quantum gravity Brady_2013 ; pelliconi2024gravitymesoscopic , and quantum information theory Brady_2013 ; Bentsen_2021 ; liu2023signaturescrambloneffectivefield ; Zhang_2023 ; Tang:2024kpv .
-Unlike systems with static (quenched) disorder, Brownian disorder is characterized by time-dependent randomness, allowing ensemble averaging to yield analytically tractable descriptions of non-equilibrium dynamics.
-A paradigmatic example is the Brownian Sachdev-Ye-Kitaev (SYK) model Saad:2018bqo , which is significantly trackable than the regular SYK Maldacena_2016 ; Maldacena:2016hyu .
-This has led to its widespread use as a toy model for quantum gravity Narovlansky:2023lfz ; Milekhin:2023bjv ; Stanford:2023npy ; Hayden_2007 , quantum error correction (QEC) Balasubramanian:2023xdp , out-of-time-ordered correlators (OTOCs) Sunderhauf:2019djv , and other fundamental aspects of quantum chaos Jian:2020krd ; Agarwal:2020yky ; Roberts_2017 ; Cotler_2017 ; jian2022lineargrowthcircuitcomplexity ; Peng:2022pfa ; fava2023designsfreeprobability .
-Building on these developments, the recently proposed Brownian Gaussian Unitary Ensemble (BGUE) Tang:2024kpv offers an alternative setting to explore Brownian disorder in quantum systems.
-Unlike the standard Gaussian Unitary Ensemble (GUE), where the Hamiltonian is randomly drawn but remains fixed in time, BGUE describes a system evolving under a time-dependent noisy Hamiltonian,
-$H_{ij}(t)=\eta_{ij}(t),\eta_{ij}(t)=\eta_{ij}^{\ast}(t)\in\mathbb{C},$ (1)
-where $\eta_{ij}(t)$ are independent Brownian Gaussian random variables with vanishing mean and nonzero variance
-$\mathbb{E}\left(\eta_{ij}(t)\right)=0,\ \mathbb{E}\left(\eta_{ij}(t)\eta_{kl}(t^{\prime})\right)={J\over D}\delta_{il}\delta_{jk}\delta(t-t^{\prime})\ .$ (2)
-The time evolution in this setting can be regarded as the limiting case of an infinitely long piecewise stochastic evolution
-$U=e^{-\text{i}H_{1}t_{1}}e^{-\text{i}H_{2}t_{2}}\cdots e^{-\text{i}H_{n}t_{n}},$ (3)
-where each $H_{i}$ is an independent realization of a GUE matrix.
-This model exhibits several remarkable properties, including rapid convergence to unitary $k$-designs dankert2005efficientsimulationrandomquantum ; Gross_2007 ; ambainis2007quantumtdesignstwiseindependence ; Dankert_2009 ; Cotler_2017 ; Roberts_2017 ; chen2024efficientunitarytdesignsrandom , hyperfast scrambling and the emergence of temperature in de Sitter space lin2022infinitetemperatureshot .
-Compared to conventional chaotic systems, the Brownian nature of BGUE ensures that ensemble-averaged quantities remain well-defined and analytically trackable, making it a powerful tool for exploring non-equilibrium dynamics.
-One of the key motivations for studying BGUE lies in its connection to quantum information theory PhysRevADankert ; Gross2007 ; 4262758 ; Scott_2008 ; Roy_2009 ; Harrow_2009 ; Brand_o_2016 ; Nakata_2017 ; hunterjones2019unitarydesignsstatisticalmechanics .
-Random matrix models have long served as powerful tools for understanding quantum thermalization, entanglement growth, and operator spreading.
-In particular, BGUE has been shown to efficiently generate unitary designs, which play a central role in quantum information processing tasks such as classical shadow tomography Aaronson:2017qlb and quantum state reconstruction Tang:2024kpv .
-Furthermore, an essential feature of disordered quantum systems is their behavior under replica dynamics, a technique in which multiple copies of the system are introduced to probe statistical correlations.
-Many important quantum information measures—including entanglement entropy, operator norms, Rényi divergences, and multi-contour OTOC Haehl:2017pak —require analyzing multi-replica observables, making the study of $n$-replica time evolution in BGUE a crucial problem.
-However, existing studies have focused primarily on the cases $n=1,2$, leaving the general $n$-replica case largely unexplored.
-This serves as one of our motivations to study the general $n$-replica case and develop a systematic way to solve this question.
-However, a fundamental challenge in extending the analysis to arbitrary $n$ is the rapid growth in computational complexity. The effective time evolution of the $n$-replica
+Motivation
+Brownian (disordered) systems hold significant importance in various fields, including quantum gravity and quantum information.
+Unlike the quenched disorder found in typical disordered systems, Brownian disorder results in a local theory after taking the ensemble average, making these systems considerably easier to solve.
+For example, the Brownian SYK model, a crucial toy model in quantum gravity, is much more manageable than its regular counterpart Maldacena_2016 ; Maldacena:2016hyu .
+Furthermore, the results obtained from the Brownian SYK can be utilized to evaluate the spectral form factor (SFF) of the standard SYK model Saad:2018bqo .
+The connection between the Brownian SYK and de Sitter spacetime has also been explored Narovlansky:2023lfz ; Milekhin:2023bjv ; Stanford:2023npy .
+Additionally, Brownian systems can be employed to achieve unitary $k$-designs dankert2005efficientsimulationrandomquantum ; Gross_2007 ; ambainis2007quantumtdesignstwiseindependence ; Dankert_2009 , which are distributions of unitaries that reproduce the first $k$ moments of the Haar measure.
+Unitary designs play a prominent role in various areas of quantum information theory PhysRevADankert ; Gross2007 ; 4262758 ; Scott_2008 ; Roy_2009 ; Harrow_2009 ; Brand_o_2016 ; Nakata_2017 ; hunterjones2019unitarydesignsstatisticalmechanics , many-body quantum chaos Roberts_2017 ; Cotler_2017 ; jian2022lineargrowthcircuitcomplexity ; fava2023designsfreeprobability , and models of black holes Hayden_2007 .
+Recently, a model known as the Brownian Gaussian Unitary Ensemble (BGUE) was explored in Tang:2024kpv , alongside other similar models discussed in Guo:2024zmr .
+These studies consider a closed quantum system characterized by a noisy Hamiltonian
+$H_{ij}(t)=\eta_{ij}(t),\eta_{ij}(t)=\eta_{ji}^{*}(t)\in\mathbb{C}$ (1)
+where $1\leq i,j\leq D$. $D$ is the dimension of the Hilbert space. And $\eta_{ij}(t)$ are independent Brownian Gaussian random variables with vanishing mean and non-vanishing variances
+$\mathbb{E}\left(\eta_{ij}(t)\right)=0,~\mathbb{E}\left(\eta_{ij}(t)\eta_{kl}(t^{\prime})\right)={J\over D}\delta_{il}\delta_{jk}\delta(t-t^{\prime})\ .$ (2)
+Here, $\mathbb{E}$ denotes the disorder average.
+The model can be regarded as an infinite version of the evolution $U=e^{-\text{i}H_{1}t_{1}}e^{-\text{i}H_{2}t_{2}}\cdots e^{-\text{i}H_{n}t_{n}},\text{i}=\sqrt{-1}$, where each $H_{i}$ is independently drawn from the Gaussian Unitary Ensemble (GUE).
+Researchers consider this piecewise evolution and expect it to approach the Haar measure much faster than the pure evolution $e^{-iHt}$ Cotler_2017 ; Roberts_2017 ; chen2024efficientunitarytdesignsrandom on a timescale independent of $D$.
+Similar to the Brownian SYK model, it was found that the BGUE exhibits hyperfast scrambling and the emergence of temperature lin2022infinitetemperatureshot .
+In contrast to systems without disorder, when dealing with general $n$-replica observables in this model, we ultimately obtain an effective time evolution on $2n$-contours:
 $\mathcal{U}_{n}(t)\equiv\mathbb{E}\left(U^{\otimes n}\otimes U^{*\otimes n}\right)\equiv e^{\mathcal{L}_{n}t}$ (3)
-evolves within a Hilbert space of dimension $D^{2n}$, making direct matrix representations infeasible for large $n$.
-To simplify this question, we should note that the calculation of the $n$-replica time evolution operator $\mathcal{U}_{n}(t)$ can be reduced to finding the general expression for the moments of $\mathcal{L}_{n}$, allowing for a straightforward expansion:
+where $U(t)\equiv e^{-\text{i}Ht}$, $\mathcal{L}_{n}$ is the generator operator of $\mathcal{U}_{n}(t)$, functioning as the “Hamiltonian” for the Euclidean time evolution in the $D^{2n}$-dimensional Hilbert space $\mathcal{H}^{\otimes 2n}$.
+The calculation method bears similarities to random tensor network models Hayden_2016 and random unitary circuit models Nahum_2017 ; Nahum_2018 ; von_Keyserlingk_2018 ; Zhou_2020 ; Fisher_2023 .
+In Tang:2024kpv , the author has derived explicit analytical results for one-replica ($n=1$) and two-replica ($n=2$) time evolutions, enabling the evaluation of two-point functions, OTOCs, entropy, and other quantities in the model.
+For instance, we find a non-decaying two-point function as $t\to\infty$ Maldacena_2003 ; saad2019latetimecorrelationfunctions and non-vanishing fluctuations of the two-point function with a zero mean value stanford2020quantumnoisewormholes .
+Moreover, the application of the model to $n$-design and classical shadow tomography is also discussed.
+As argued in the paper, many interesting observables require three (or more) replicas, such as the max-version shadow norm.
+This serves as one of our motivations to study the general $n$-replica case.
+Note that the calculation of the $n$-replica time evolution operator $\mathcal{U}_{n}(t)$ can be reduced to finding the general expression for the moments of $\mathcal{L}_{n}$, allowing for a straightforward construction of $\mathcal{U}_{n}(t)$
 $\mathcal{U}_{n}(t)=\sum_{r=0}^{\infty}{(\mathcal{L}_{n})^{r}t^{r}\over r!}\ .$ (4)
-We find that the linearly independent terms, which we refer to as “graphs”, that appear in the moments of $\mathcal{L}_{n}$ are finite in number, allowing for a natural matrix representation.
-In the case of $n=2$, the author of Tang:2024kpv categorized 24 individual graphs into 8 groups, reducing the matrix dimension from 24 to 8.
-However, as $n$ increases, the number of individual graphs that emerge in the moments of $\mathcal{L}_{n}$ grows rapidly (see Table 1), necessitating a systematic method to group these graphs into appropriate categories to achieve a significantly smaller-dimensional representation.
-In this work, we develop a graph-theoretic framework for analyzing the time evolution of $n$-replica observables in BGUE.
-Our approach generalizes and extends the graph-based categorization introduced for $n=2$, introducing a more structured notation that enables an algebraic treatment of replica dynamics.
-We demonstrate that the moments of $\mathcal{L}_{n}$ can be mapped onto a finite set of graph categories, significantly reducing computational complexity.
-For $n=3$, we have 26 graph categories, so that $\mathcal{L}_{n}$ is a $26$-dimensional matrix. This enables us to obtain an explicit analytical expression for
-$e^{\mathcal{L}_{3}t}=\sum_{a=1}^{26}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\ ,$ (6)
-where the analytical expressions of $f_{a}(t)$ are listed in Appendix A. Furthermore, we discuss the generalization of our framework to arbitrary $n$ and present explicit results for $n=4$.
-Compared to representation-theoretic approaches, our graph-based method offers a lower-dimensional irreducible representation, reducing numerical computation time by handling a lower-dimensional matrix.
-Previous approaches, such as those in Guo:2024zmr , relied on
+We find that the linearly independent terms, which we refer to as “graphs,” that appear in the moments of $\mathcal{L}_{n}$ are finite in number, allowing for a natural matrix representation.
+For the case of $n=2$, author in Tang:2024kpv previously addressed this problem by analyzing the features of the 24 individual graphs and categorizing them into 8 groups, effectively reducing the dimension of the matrix representation from 24 to 8.
+However, as $n$ increases, the number of individual graphs that emerge in the moments of $\mathcal{L}_{n}$ grows rapidly (see Table 1), creating the challenge of how to effectively group these graphs into appropriate categories to achieve a significantly smaller-dimensional matrix representation.
+In this paper, we introduce a systematic notation for the graphs in the model, which facilitates a purely algebraic approach to the graph classification problem.
+This method enables us to derive a compact matrix representation for $\mathcal{L}_{n}$, making it a valuable tool for investigating the complex behaviors governed by this operator.
+We anticipate that the methodology presented in this paper can be applied to other models involving Brownian disorder.
+The model
+Here, we can generalize the BGUE by introducing an intrinsic spectral term
+$H_{ij}=E_{i}\delta_{ij}+\eta_{ij}(t)\ .$ (5)
+Summary of results
+In this work, we establish a general framework to address the $n$-replica time evolution problem.
+We provide detailed calculations for the cases of $n=2$ and $n=3$ as illustrative examples, presenting the analytical expressions for the corresponding evolution operators $\mathcal{U}_{2}$ and $\mathcal{U}_{3}$.
+For $n=3$, we have 26 graph categories, so that $\mathcal{L}_{n}$ is a $26$-dimensional matrix, then we obtain $e^{\mathcal{L}_{3}t}=\sum_{a=1}^{26}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}$, the analytical expressions of $f_{a}(t)$ are listed in Appendix A.
+Note Added
+After completing the paper, we found an alternative approach discussed in Guo:2024zmr . Noticing
 $[\mathcal{L}_{n},V^{\otimes n}\otimes V^{*\otimes n}]=0,\forall V\in U(D)\ ,$ (6)
-which allows for the determination of eigenvalues and wavefunctions for $\mathcal{L}_{n}$ by decomposing direct products of SU$(d)$ into irreducible representations.
+which allows for the determination of eigenvalues and wavefunctions for $\mathcal{L}_{n}$ by decomposing direct products of SU(d)(d) into irreducible representations.
 Moreover, these representations are constructed systematically using Young tableaux Hamermesh1962GroupTA .
-We will show the detailed result for $n=2$ at the end of Sec.
+Thus, they have provided a systematic method for calculating any observable in the Brownian GUE.
+[^1] Similarly, one can solve Brownian GOE and Brownian GSE.
+We believe that the method presented in this paper is fundamentally similar to that in Guo:2024zmr , albeit expressed in a different form.
+In cases where symmetries like Eq.(6) do not exist, the methodology discussed here can serve as a valuable toolbox.
+Structure of the paper
+We first introduce the denotation for graphs in Sec.
+2.
+Then we study the $n=2$ case as a warm-up in Sec.
 3.
-The graph representation also presents a straightforward way to compute the expectation values of observables.
-In cases where symmetries like (7) do not exist, the methodology discussed here can serve as a valuable toolbox.
-Beyond its application to BGUE, our framework has broader implications to other Brownian disordered systems, like random unitary circuit Nahum_2017 ; Nahum_2018 ; von_Keyserlingk_2018 ; Zhou_2020 ; Fisher_2023 and random tensor networks Hayden_2016 ; Hosur:2015ylk .
-The techniques developed here could provide new insights into the replica method in many-body quantum chaos, as well as potential applications to problems in wormholes and quantum gravity Maldacena_2003 ; saad2019latetimecorrelationfunctions ; stanford2020quantumnoisewormholes ; Peng:2021vhs .
-Structure of the Paper
-The remainder of this paper is structured as follows.
-In Sec.
-2, we introduce the notation and graph-based formalism for representing the time evolution operator $\mathcal{U}_{n}(t)$.
-In Sec.
-3, we apply our framework to the case $n=2$ as a warm-up example, treated as a double check with the results from Tang:2024kpv .
-Sec.
-4 extends this analysis to the non-trivial case $n=3$, where we show how to find the graph categories and assign the graphs into them.
-In Sec.
-5, we present a general algorithm for arbitrary $n$ and illustrate its application to $n=4$.
-Finally, in Sec.
-6, we discuss potential future directions and broader implications of our findings.
-2 Analysis on Graphs
+Next, we evaluate the non-trivial $n=3$ case in Sec.
+4.
+Finally, we present a universal algorithm for the general $n$ case in Sec.
+5, where the $n=4$ case is also discussed.
+Lastly, we provide some comments on future research in Sec.
+6.
+2 Denotation for graphs
 In this section, we introduce a concise notation to represent arbitrary graphs, which facilitates the evaluation of graph contractions and enables a more transparent classification of graphs.
@@
 At first, we introduce some notations for the graph representation of $\mathcal{U}_{n}$ and $\mathcal{L}_{n}$.
-We treat $\mathcal{U}_{n},\mathcal{L}_{n}$ as a $D^{2n}$-dimensional matrix, with row and column indices labeled by two $n$-length lists: $\{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}\}$ and $\{I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}\}$.
+We treat $\mathcal{U}_{n}/\mathcal{L}_{n}$ as a $D^{2n}$-dimensional matrix, with row and column indices labeled by two $n$-length lists: $\{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}\}$ and $\{I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}\}$.
 We then adapt this to the graph representation (taking $n=2$ as an example):
-$\left(\mathcal{U}_{2}(t)\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ $\equiv\mathbb{E}\left(U_{I_{1}I_{1}^{\prime}}U_{J_{1}J_{1}^{\prime}}^{*}U_{I_{2}I_{2}^{\prime}}U_{J_{2}J_{2}^{\prime}}^{*}\right)=\mathbb{E}\left[\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\right]=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\ .$ (8)
+$\left(\mathcal{U}_{2}(t)\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ $\equiv\mathbb{E}\left(U_{I_{1}I_{1}^{\prime}}U_{J_{1}J_{1}^{\prime}}^{*}U_{I_{2}I_{2}^{\prime}}U_{J_{2}J_{2}^{\prime}}^{*}\right)=\mathbb{E}\left[\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\right]=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\ .$ (7)
 We label the indices in graphs using uppercase letters $I,J,I^{\prime},J^{\prime}$, where $I/I^{\prime}$ represent the left/right-hand side indices of a contour $H$, and $J/J^{\prime}$ denote the dual contour $H^{*}$.
 For simplicity, we label the contours (and dual contours) with lowercase letters $i,j,k,l$ (and their corresponding bars $\bar{i},\bar{j},\bar{k},\bar{l}$).
-Since we take $n=2$, we can also represent $\mathcal{U}_{2}$ as
-$\mathcal{U}_{2}(t)=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}$ (9)
+Therefore, we also denote
+$\mathcal{U}_{2}(t)=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}$ (8)
 By a similar derivation as shown in Tang:2024kpv , it is straightforward to obtain
-$\mathcal{L}_{n}=\mathsf{w}\mathbb{I}+{J\over D}\sum_{i\bar{j}}^{n}P_{i\bar{j}}-{J\over D}\sum_{ij}^{n}(X_{ij}+X_{\bar{i}\bar{j}})\quad,$ (10)
-where $J$ is the strength of noise, $\mathbb{I}$ denotes identity operator. Other two kinds of operators are represented as below:
-$P_{i\bar{j}}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$},~{}X_{ij}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\quad,$ (11)
-where each line with two indices $i,j$ at the endpoints of the graph represents a “propagator” between the sites on the contour and the dual contour, i.e., $\delta_{ij}$.
-Other trivial propagators connecting sites on the same contour, which are represented as straight lines, are given by $\delta_{I_{i}I_{i}^{\prime}}$ or $\delta_{J_{i}J_{i}^{\prime}}$.
+$\mathcal{L}_{n}=\mathsf{w}\mathbb{I}+{J\over D}\sum_{i\bar{j}}^{n}P_{i\bar{j}}-{J\over D}\sum_{ij}^{n}(X_{ij}+X_{\bar{i}\bar{j}})\ .$ (9)
+Here, $\mathbb{I}$ denotes the identity operator, and we sometimes simply denote it as $I$. The other two kinds of operators are represented below:
+$P_{i\bar{j}}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$},~X_{ij}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\ .$ (10)
+where each line with two indices $i,j$ at the endpoints of the graph represents a ”propagator,” i.e., $\delta_{ij}$.
+Other trivial contours, represented by straight lines, are given by $\delta_{I_{i}I_{i}^{\prime}}$ or $\delta_{J_{i}J_{i}^{\prime}}$.
 Taking $n=2$ as an example, we have
-$\left(P_{1\bar{2}}\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ =[$\right]_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}=\delta_{I_{1}J_{2}}\delta_{I_{1}^{\prime}J_{2}^{\prime}}\delta_{J_{1}J_{1}^{\prime}}\delta_{I_{2}I_{2}^{\prime}}\ ,$
-$\left(X_{12}\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ =[$\right]_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}=\delta_{I_{1}I_{2}^{\prime}}\delta_{I_{2}I_{1}^{\prime}}\delta_{J_{1}J_{1}^{\prime}}\delta_{J_{2}J_{2}^{\prime}}\ .$ (12)
-And $\mathsf{w}$ in (10) is a factor that carries the indices of a graph
+$\left(P_{1\bar{2}}\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ $=\left[\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\right]_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}=\delta_{I_{1}J_{2}}\delta_{I_{1}^{\prime}J_{2}^{\prime}}\delta_{J_{1}J_{1}^{\prime}}\delta_{I_{2}I_{2}^{\prime}}\ ,$
+$\left(X_{12}\right)_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}$ $=\left[\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\right]_{I_{1}J_{1}I_{2}J_{2};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}}=\delta_{I_{1}I_{2}^{\prime}}\delta_{I_{2}I_{1}^{\prime}}\delta_{J_{1}J_{1}^{\prime}}\delta_{J_{2}J_{2}^{\prime}}\ .$ (11)
+And $\mathsf{w}$ in Eq. (9) is a factor that carries the indices of a graph
 $\mathsf{w}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}=-\text{i}\sum_{j=1}^{n}(E_{I_{j}}-E_{J_{j}})-nJ,~\text{i}=\sqrt{-1}\ .$ (12)
-Notice that the factor $\mathsf{w}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}$ commutes with all of the terms in $\mathcal{L}_{n}$. For instance,
+Notice that the factor $\mathsf{w}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}$ commutes with all of the terms in $\mathcal{L}_{n}$, e.g.,
 $\mathsf{w}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}\left(P_{a\bar{b}}\right)_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}=\left(P_{a\bar{b}}\right)_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}\mathsf{w}_{I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}\ .$ (13)
@@
 $\mathsf{E}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}=-\text{i}\sum_{j=1}^{n}(E_{I_{j}}-E_{J_{j}})\ .$ (14)
-Since $\mathsf{w}=\mathsf{E}-nJ\mathbb{I}$, the factor $\mathsf{E}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}$ also commutes with all terms in $\mathcal{L}_{n}$.
-2.1 Notations of a Graph
-In this work, we employ a concise notation $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$ to represent an arbitrary graph.
-Here, $0\leq p\leq n$ denotes the number of paired vertices, and $\sigma\equiv\sigma_{H}\otimes\sigma_{H^{*}}\in S_{n}\otimes S_{n}$ encodes the permutation acting on the indices appearing on the right-hand side of the graph.
-The element of a $D^{2n}$-dimensional matrix $\mathsf{F}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}$ would be:
+Since $\mathsf{w}=\mathsf{E}-nJ\mathbf{1}$, the factor $\mathsf{E}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n}}$ also commutes with all terms in $\mathcal{L}_{n}$.
+2.1 Notation for a graph
+In this work, we employ the concise notation $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$ to represent an arbitrary graph.
+Here, $0\leq p\leq n$ denotes the number of paired vertices, and $\sigma\equiv\sigma_{1}\otimes\sigma_{2}\in S_{n}\otimes S_{n}$ encodes the permutation acting on the indices appearing on the right-hand side of the graph (or equivalently, a $D^{2n}$-dimensional matrix $\mathsf{F}_{I_{1}J_{1}I_{2}J_{2}\ldots I_{n}J_{n};I_{1}^{\prime}J_{1}^{\prime}I_{2}^{\prime}J_{2}^{\prime}\ldots I_{n}^{\prime}J_{n}^{\prime}}$:
 $\left(\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]\right)_{\{IJ\};\{I^{\prime}J^{\prime}\}}=\prod_{i\not\in\{a\},\bar{j}\not\in\{\bar{b}\}}\delta_{I_{i}I^{\prime}_{\sigma_{1}\left(i\right)}}\delta_{J_{\bar{j}}J^{\prime}_{\sigma\left(\bar{j}\right)}}\prod_{i=1}^{p}\delta_{I_{a_{i}}J_{\bar{b}_{i}}}\delta_{I^{\prime}_{\sigma_{1}\left(a_{i}\right)}J^{\prime}_{\sigma_{2}\left(\bar{b}_{i}\right)}}\ .$ (15)
-Here, $c[\sigma]$ is the permutation on the contour labels on the right-hand side, $c_{i\bar{j}}$ represents the pair between the left sites $i,\bar{j}$ and the right sites $\sigma(i)^{\prime},\sigma(\bar{j})^{\prime}$.
-To make the complicated notations clear, we can take a simple example $c_{1\bar{1}}c[X_{12}\otimes X_{\bar{1}\bar{2}}]=\delta_{I_{1}J_{1}}\delta_{I_{2}^{\prime}J_{2}^{\prime}}\delta_{I_{2}I_{1}^{\prime}}\delta_{J_{2}J_{1}^{\prime}}$ and show how the graphs work as follows:
+Here, $c[\sigma]$ means to act the permutation on the contour labels on the right-hand side, $c_{i\bar{j}}$ means to connect the left points $i,\bar{j}$ and right points $\sigma(i)^{\prime},\sigma(\bar{j})^{\prime}$; finally, we connect the remaining points $i,\sigma(i)^{\prime}$ and $j,\sigma(\bar{j})^{\prime}$.
+Taking $c_{1\bar{1}}c[X_{12}\otimes X_{\bar{1}\bar{2}}]=\delta_{I_{1}J_{1}}\delta_{I_{2}^{\prime}J_{2}^{\prime}}\delta_{I_{2}I_{1}^{\prime}}\delta_{J_{2}J_{1}^{\prime}}$ as an example:
 […]
-The filled dots are the sites on the contour $H$ and the hollow circles are those on the dual contour $H^{*}$.
-And we can directly find how the swapping and connecting operations work in the first two steps.
-The graph is finally denoted as the R.H.S.
-of (17).
-The notations presented in (16) for the graphs make it much easier to calculate the contraction (product) of two graphs and study the action of $\mathcal{L}_{n}$ on a certain graph, as we will see later in this section.
-Symmetries of a Graph
-Before examining the action of $P_{i\bar{j}}$ and $X_{ij},X_{\bar{i}\bar{j}}$ on a graph $\mathsf{F}$, we note that the notations presented above have some redundancy.
+Since we use $I$ to indicate the indices of the contours $1,2$ and $J$ for $\bar{1},\bar{2}$, we do not distinguish $J_{1}$ and $J_{\bar{1}}$.
+In this paper, we use black disks to label contour vertices and circles to label the dual vertices.
+For simplicity, we will not draw the channel labels $i,\bar{i}$ or the indices $I_{i},J_{i}$ in the graphs.
+Moreover, the denotation given in Eq.(15) for the graphs makes it much easier to calculate the contraction (product) of two graphs and study the action of $\mathcal{L}_{n}$ on a certain graph.
+Symmetry of a graph
+Before examining the action of $P_{i\bar{j}}$ and $X_{ij},X_{\bar{i}\bar{j}}$ on a graph $\mathsf{F}$, we note that the notation presented here contains some redundancy.
 For a graph $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$, the vertices for each connected pair $(a_{i},\bar{b}_{i})$ are identical.
-To simplify it, we define the action of a permutation on the graph as
+To simplify matters, we define the action of a permutation on the graph
 $\sigma^{\prime}*\mathsf{F}$ $\equiv\sigma^{\prime}(\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}})c[\sigma\sigma^{\prime}]=\prod_{i=1}^{p}c_{\sigma^{\prime}_{1}(a_{i})\sigma^{\prime}_{2}(\bar{b}_{i})}c[\sigma\sigma^{\prime}]\ .$ (17)
-If a permutation $\sigma^{\prime}=\sigma^{\prime}_{H}\otimes\sigma^{\prime}_{H^{*}}$ is an element of the group $G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})$ generated by $\{X_{a_{i}a_{j}}\otimes X_{\bar{b}_{i}\bar{b}_{j}}\}_{i,j=1,2,\ldots,p}$, then its action is trivial: $\sigma^{\prime}*\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b_{i}}}c[\sigma\sigma^{\prime}]=\mathsf{F}$.
+If a permutation $\sigma^{\prime}=\sigma^{\prime}_{1}\otimes\sigma^{\prime}_{2}$ is an element of the group $G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})$ generated by $\{X_{a_{i}a_{j}}\otimes X_{\bar{b}_{i}\bar{b}_{j}}\}_{i,j=1,2,\ldots,p}$, then its action is trivial: $\sigma^{\prime}*\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b_{i}}}c[\sigma\sigma^{\prime}]=\mathsf{F}$.
 It can be observed that $G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})$ resembles the gauge group in quantum field theory and is isomorphic to the permutation group $S_{p}$.
 Number of individual graphs
-$n$ $N_{n}$#\## Graph Categories
+$n$ $N_{n}$ #\# Graph Categories
 1 2 2
@@
 For general $n$, using $|G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})|=p!$ and $|S_{n}\otimes S_{n}|=(n!)^{2}$, we have
-$N_{n}$ $=\sum_{\boldsymbol{a},\bar{\boldsymbol{b}}}{|S_{n}\otimes S_{n}|\over|G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})|}=\sum_{p=0}^{n}\binom{n}{p}^{2}p!\frac{\left(n!\right)^{2}}{p!}=\sum_{p=0}^{n}\left(n!\binom{n}{p}\right)^{2}\equiv\sum_{p=0}^{n}N_{n,p}\ .$ (19)
+$N_{n}$ $=\sum_{\boldsymbol{a},\bar{\boldsymbol{b}}}{|S_{n}\otimes S_{n}|\over|G_{n}(\boldsymbol{a},\bar{\boldsymbol{b}})|}=\sum_{p=0}^{n}\binom{n}{p}^{2}p!\frac{\left(n!\right)^{2}}{p!}=\sum_{p=0}^{n}\left(n!\binom{n}{p}\right)^{2}\ .$ (18)
 One can utilize the individual graphs as a basis, then $\mathcal{L}_{n}\mathsf{F}_{i}=\sum_{j=1}^{N_{n}}\widetilde{M}_{ji}\mathsf{F}_{j}$.
-However, as indicated in (19), the dimension of $\widetilde{M}$ grows too quickly with $n$.
+However, as indicated in Eq.
+(18), the dimension of $\widetilde{M}$ grows too quickly with $n$.
 A clever approach is to group these individual graphs into categories $\mathsf{F}_{\alpha}=\sum_{i=1}^{N_{n;\alpha}}\mathsf{F}_{\alpha,i}$ such that
-$\mathcal{L}_{n}\mathsf{F}_{\alpha}=\sum_{\beta=1}^{N_{n}^{\text{ca}}}M_{\beta\alpha}\mathsf{F}_{\beta}\ .$ (20)
+$\mathcal{L}_{n}\mathsf{F}_{\alpha}=\sum_{\beta=1}^{N_{n}^{\text{ca}}}M_{\beta\alpha}F_{\beta}\ .$ (19)
 In this paper, we further require that $\mathsf{F}_{\alpha}\cap\mathsf{F}_{\beta}=\emptyset$ for all $\alpha\neq\beta$.
-The dimension of the transfer matrix $M$, denoted as $N_{n}^{\text{ca}}$, can be significantly smaller than $N_{n}$, as illustrated in Table 1.
-2.2 Operating $\mathcal{L}_{n}$ on an Individual Graph
-Figure 1: Four non-trivial cases of $P_{i\bar{j}}\mathsf{F}$.
+The dimension of $M_{\beta\alpha}$, denoted as $N_{n}^{\text{ca}}$, can be significantly smaller than $N_{n}$, as illustrated in Table 1.
+2.2 Action of $\mathcal{L}_{n}$ on an individual graph
+Figure 1: The four non-trivial cases of $P_{i\bar{j}}\mathsf{F}$.
 Since $\mathcal{L}_{n}$ consists of $P_{i\bar{j}}$ and $X_{ij},X_{\bar{i}\bar{j}}$, we first evaluate their action on an individual graph $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$.
-We denote the set of paired sites as $C(\mathsf{F})=\{(a_{i},\bar{b}_{i})\}_{i=1}^{p}$.
-For later reference, we also define the set of all possible paired sites between different contours as $\mathcal{C}$ and those on the same contour as $\mathcal{E}$ and $\bar{\mathcal{E}}$, respectively:
+We denote the set of pair contours as $C(\mathsf{F})=\{(a_{i},\bar{b}_{i})\}_{i=1}^{p}$.
+For later reference, we define the set of all possible pair contours as $\mathcal{C}$ and the exchange contours as $\mathcal{E}$ and $\bar{\mathcal{E}}$:
 […]
-Operation of $P_{i\bar{j}}$
+Action of $P_{i\bar{j}}$
 There are five cases:
@@
 Since the first case is trivial, we illustrate the other four cases in Figure 1.
-Operation of $X_{ij},{X}_{\bar{i}\bar{j}}$
+Action of $X_{ij},{X}_{\bar{i}\bar{j}}$
 There are five cases depending on the result of $(i,j)\cap C(\mathsf{F})$. We depict two nontrivial cases in Figure 2. It is observed that we always have
@@
 $X_{\bar{i}\bar{j}}\mathsf{F}=X_{\bar{i}\bar{j}}*\mathsf{F}.$
-Operation of $\mathcal{L}_{n}$
-Next, we can study the operation of $\mathcal{L}_{n}$.
-Notice that we have the condition $\bar{u}\ \cap\ C[\mathsf{F}]=\emptyset$.
-Hence the permutations $X_{u}$ and $X_{\bar{u}}$ act trivially on $C(\mathsf{F})$.
+Action of $\mathcal{L}_{n}$
+Next, we can study the action of $\mathcal{L}_{n}$.
+Noting the condition $\bar{u}\cap C[\mathsf{F}]=\emptyset$, the permutations $X_{u}$ and $X_{\bar{u}}$ act trivially on $C(\mathsf{F})$.
 Thus, we ultimately find a simple relation
@@
 where $|C[\mathsf{F}]|$ counts the number of pairs, and $\mathsf{F}[\sigma\sigma^{\prime}]\equiv\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma\sigma^{\prime}]$.
-The definitions of $u$ and $\bar{u}$ are encoded in the summation expressions.
+The definitions of $u$ and $e$ are encoded in the summation expressions.
 For example, in the term $\sum_{u\cap C[\mathsf{F}]=\emptyset}c_{u}\mathsf{F}$, $u$ can be considered as a pair in $\mathcal{C}$ that is excluded by $C(\mathsf{F})$.
 Meanwhile, in the term $\sum_{u\cap C[\mathsf{F}]=\emptyset}\mathsf{F}[\sigma X_{u}]$, $u$ can be regarded as an element of $\mathcal{E}$ that is excluded by $C(\mathsf{F})$.
-Here we choose the set of $u$ to meet those conditions of $u\cap C[\mathsf{F}]=\emptyset$ from all the summations, which requires flattening $C(\mathsf{F})$ into a list of site labels and find those $u$ that are available.
+There is a commitment to taking the common set $u\cap C[\mathsf{F}]$, which requires flattening $C(\mathsf{F})$ into a list of contour labels.
 For example, taking $u=\{(1,\bar{1})\},C(\mathsf{F})=\{(1,\bar{2}),(2,\bar{3})\}$, then
-$u\cap C(\mathsf{F})=\{1,\bar{1}\}\cap\{1,\bar{2},2,\bar{3}\}=\{1\}\not=\emptyset\ ,$ (24)
-which shows that $u=\{(1,\bar{1})\}$ is not available.
+$u\cap C(\mathsf{F})=\{1,\bar{1}\}\cap\{1,\bar{2},2,\bar{3}\}=\{1\}\not$$\emptyset\$. (23)
 We always define three graph categories, as they play a crucial role in the definition of $\mathcal{L}_{n}$:
-$\mathsf{F}_{0}\equiv c[I],\mathsf{F}_{1}\equiv\sum_{u}c_{u}c[I],\mathsf{F}_{0,X+\bar{X}}\equiv\sum_{e}c[X_{e}]+\sum_{\bar{e}}c[X_{\bar{e}}]\ .$ (25)
+$\mathsf{F}_{0}\equiv c[I],\mathsf{F}_{1}\equiv\sum_{u}c_{u}c[I],\mathsf{F}_{0,X+\bar{X}}\equiv\sum_{e}c[X_{e}]+\sum_{\bar{e}}c[X_{\bar{e}}]\ ,$ (24)
 then we apply $\mathcal{L}_{n}$ to the three initial categories and their descendants until a complete basis is obtained.
-After that, we obtain the expression for the transfer matrix $M$ in (20), which allows us to solve $\mathcal{U}_{n}(t)$ using (5).
-We will illustrate the basic idea with examples for $n=2$ and $n=3$ in the next two sections.
-Figure 2: Two non-trivial cases of $X_{ij}\mathsf{F}$.
+After that, we obtain the expression for $M$ in Eq.
+(19), which allows us to solve $\mathcal{U}_{n}(t)$ using Eq.
+(4).
+We illustrate the basic idea with examples for $n=2$ and $n=3$.
+Figure 2: The two non-trivial cases of $X_{ij}\mathsf{F}$.
 3 Time evolution for $n=2$
-As a warm-up, we consider $n=2$, which gives us four sites: $\{1,\bar{1},2,\bar{2}\}$. The permutation group is $S_{2}\otimes S_{2}$, where $S_{2}=\{I,X_{12}\}$. Then
-$\mathcal{C}$ =$=\left\{\left(1,\bar{1}\right),\left(1,\bar{2}\right),\left(2,\bar{1}\right),\left(2,\bar{2}\right)\right\}\ ,$ (26)
+As a warm-up, we consider $n=2$, which gives us four contours: $\{1,\bar{1},2,\bar{2}\}$. The permutation group is $S_{2}\otimes S_{2}$, where $S_{2}=\{I,X_{12}\}$. Then
+$\mathcal{C}$ $=\left\{\left(1,\bar{1}\right),\left(1,\bar{2}\right),\left(2,\bar{1}\right),\left(2,\bar{2}\right)\right\}\ ,$ (25)
 […]
@@
 The number of individual graphs is $24$, while the number of operators ($P_{i\bar{j}},X_{ij},X_{\bar{i}\bar{j}}$) is $6$.
-Therefore, in principle, one needs to calculate $6\times 24=168$ graph contractions when calculating $M$ as defined earlier.
-Fortunately, (23) enables us to perform a simple algebraic calculation without drawing any graphs.
+Therefore, in principle, one needs to calculate $6\times 24=168$ graph contractions when evaluating $M$ as defined earlier.
+Fortunately, Eq.
+(22) enables us to perform a simple algebraic calculation without the need to draw or recognize any graphs.
 3.1 Counting graph categories
-As we have discussed earlier, we will find the appropriate graph groups by applying $\mathcal{L}_{n}$ to the three initial categories $\mathsf{F}_{0},\mathsf{F}_{1},\mathsf{F}_{0,X+\bar{X}}$ and their descendants until no new graph category is generated.
-In the case of $n=2$, we will perform $\mathcal{L}_{2}$ on the initial sets as follows.
+As discussed earlier, we aim to find the appropriate graph groups by applying $\mathcal{L}_{n}$ to the three initial sets $\mathsf{F}_{0},\mathsf{F}_{1},\mathsf{F}_{0,X+\bar{X}}$ and their descendants until no new graph sets are generated.
 $\mathcal{L}_{2}$ acts on $\mathsf{F}_{0},\mathsf{F}_{1},\mathsf{F}_{0,X+\bar{X}}$
-  • $\mathcal{L}_{2}\mathsf{F}_{0}$: The operation of $\mathcal{L}_{2}$ on $\mathsf{F}_{0}$ is trivial, since $\mathsf{F}_{0}$ is a trivial set
+  • $\mathcal{L}_{2}\mathsf{F}_{0}$: The action of $\mathcal{L}_{2}$ on $\mathsf{F}_{0}$ is trivial
     $\mathcal{L}_{2}\mathsf{F}_{0}=w\mathsf{F}_{0}+{J\over D}\mathsf{F}_{1}-{J\over D}\mathsf{F}_{0,X+\bar{X}}\ .$ (27)
@@
 Therefore, we have
-    $\mathcal{L}_{2}\mathsf{F}_{1}=\left(w+J\right)\mathsf{F}_{1}+\frac{2J}{D}\mathsf{F}_{2}\ ,$ (30)
+    $\mathcal{L}_{2}\mathsf{F}_{1}=\left(w+J\right)\mathsf{F}_{1}+\frac{2J}{D}\mathsf{F}_{2}$ (29)
     where we have defined
@@
     $\hskip 15.0pt-\frac{J}{D}\sum_{e,u}\left(c[X_{e}X_{u}]+c[X_{\bar{e}}X_{u}]+c[X_{e}X_{\bar{u}}]+c[X_{\bar{e}}X_{\bar{u}}]\right)\ ,$ (31)
-    then we can define $\mathsf{F}_{1,X+\bar{X}}\equiv\sum_{u}c_{u}\mathsf{F}_{0,X+\bar{X}}$. Notice that
+    then we can define $\mathsf{F}_{1,X+\bar{X}}\equiv\sum_{u}c_{u}\mathsf{F}_{0,X+\bar{X}}$. Notice
     $\sum_{e,u}\left(c[X_{e}X_{u}]+c[X_{\bar{e}}X_{u}]+c[X_{e}X_{\bar{u}}]+c[X_{\bar{e}}X_{\bar{u}}]\right)=2c[I]+2c[X_{12}X_{\bar{1}\bar{2}}]\ .$ (32)
@@
     $\mathcal{L}_{2}\mathsf{F}_{0,X+\bar{X}}=w\mathsf{F}_{0,X+\bar{X}}+\frac{J}{D}\mathsf{F}_{1,X+\bar{X}}-\frac{2J}{D}\left(\mathsf{F}_{0}+\mathsf{F}_{0,X\bar{X}}\right)\ .$ (33)
-We find that three new graph categories, $\mathsf{F}_{2},\mathsf{F}_{1,X+\bar{X}},\mathsf{F}_{0,X\bar{X}}$, emerge after applying $\mathcal{L}_{2}$ on initial sets, which is the first order calculations.
-Next, we consider the second order, which means the operation of $\mathcal{L}_{2}$ on these new categories.
-For clarity, we restate the definitions here
-$\mathsf{F}_{2}=c_{1\bar{1}}c_{2\bar{2}}c[I]+c_{1\bar{2}}c_{2\bar{1}}c[I]\ ,$ (35)
-$\mathsf{F}_{1,X+\bar{X}}=\sum_{u}c_{u}\mathsf{F}_{0,X+\bar{X}}\ ,$ (36)
-$\mathsf{F}_{0,X\bar{X}}=c[X_{12}X_{\bar{1}\bar{2}}]\ .$ (37)
+We find that three new graph categories, $\mathsf{F}_{2},\mathsf{F}_{1,X+\bar{X}},\mathsf{F}_{0,X\bar{X}}$, emerge when calculating $\mathcal{L}_{2}^{2}$.
+Next, we consider the action of $\mathcal{L}_{2}$ on these new categories.
 $\mathcal{L}_{2}$ acts on $\mathsf{F}_{2},\mathsf{F}_{1,X+\bar{X}},\mathsf{F}_{0,X\bar{X}}$
-  • $\mathcal{L}_{2}\mathsf{F}_{2}$: Notice that only the first term in (23) survives. The operation of $\mathcal{L}_{2}$ on $\mathsf{F}_{2}$ is straightforward to obtain
+  • $\mathcal{L}_{2}\mathsf{F}_{2}$: Notice that only the first term in Eq. (22) survives. The action of $\mathcal{L}_{2}$ on $\mathsf{F}_{2}$ is straightforward to obtain
     $\mathcal{L}_{2}\mathsf{F}_{2}$ $=\left(w+2J\right)\mathsf{F}_{2}\ .$ (34)
-  • $\mathcal{L}_{2}\mathsf{F}_{1,X+\bar{X}}$: Using (23), we have
+  • $\mathcal{L}_{2}\mathsf{F}_{1,X+\bar{X}}$: Using Eq.(22), we have
     $\mathcal{L}_{2}\mathsf{F}_{1,X+\bar{X}}$ $=\left(w+J\right)\mathsf{F}_{1,X+\bar{X}}+\frac{J}{D}\left(\sum_{e}\sum_{g\cap u=\emptyset}c_{g}c_{u}c[X_{e}]+\left(e\leftrightarrow\bar{e}\right)\right)\ ,$ (35)
@@
     $\mathcal{L}_{2}\mathsf{F}_{1,X+\bar{X}}=\left(w+J\right)\mathsf{F}_{1,X+\bar{X}}+{4J\over D}\mathsf{F}_{2,X+\bar{X}}\ .$ (37)
-  • $\mathcal{L}_{2}\mathsf{F}_{0,X\bar{X}}$: Using (23), we have
+  • $\mathcal{L}_{2}\mathsf{F}_{0,X\bar{X}}$: Using Eq.(22), we have
     $\mathcal{L}_{2}\mathsf{F}_{0,X\bar{X}}$ $=w\mathsf{F}_{0,X\bar{X}}+\frac{J}{D}\sum_{u}c_{u}c[X_{12}X_{\bar{1}\bar{2}}]-\frac{J}{D}\left(\sum_{u}c[X_{12}X_{\bar{1}\bar{2}}X_{u}]+\sum_{u}c[X_{12}X_{\bar{1}\bar{2}}X_{\bar{u}}]\right)$
@@
     where we have used $X_{u}=\{X_{12}\}$, so that $X_{12}X_{\bar{1}\bar{2}}X_{u}=X_{\bar{1}\bar{2}},X_{12}X_{\bar{1}\bar{2}}X_{\bar{u}}=X_{12}$.
-Here, we find two new graph categories, $\mathsf{F}_{2,X+\bar{X}}$ and $\mathsf{F}_{1,X\bar{X}}$. We will then study the third order calculations by operating $\mathcal{L}_{2}$ on these categories
-$\mathsf{F}_{2,X+\bar{X}}=c_{1\bar{1}}c_{2\bar{2}}c[X_{12}]+c_{1\bar{2}}c_{2\bar{1}}c[X_{12}]\ ,$ (43)
-$\mathsf{F}_{1,X\bar{X}}=\sum_{u}c_{u}c[X_{12}X_{\bar{1}\bar{2}}]\ .$ (44)
+Here, we find two new graph categories, $\mathsf{F}_{2,X+\bar{X}}$ and $\mathsf{F}_{1,X\bar{X}}$. We will then study the action of $\mathcal{L}_{2}$ on these categories.
 $\mathcal{L}_{2}$ acts on $\mathsf{F}_{2,X+\bar{X}},\mathsf{F}_{1,X\bar{X}}$
-The calculations are straightforward by using (23)
-  • $\mathcal{L}_{2}\mathsf{F}_{2,X+\bar{X}}$:
+  • $\mathcal{L}_{2}\mathsf{F}_{2,X+\bar{X}}$: Using Eq.(22), we have
     $\mathcal{L}_{2}\mathsf{F}_{2,X+\bar{X}}=\left(w+2J\right)\mathsf{F}_{2,X+\bar{X}}\ .$ (39)
-  • $\mathcal{L}_{2}\mathsf{F}_{1,X\bar{X}}$:
+  • $\mathcal{L}_{2}\mathsf{F}_{1,X\bar{X}}$: Using Eq.(22), we have
     $\mathcal{L}_{2}\mathsf{F}_{1,X\bar{X}}$ $=\left(w+J\right)\mathsf{F}_{1,X\bar{X}}+\frac{J}{D}\sum_{v\cap u=\emptyset}\sum_{u}c_{v}c_{u}c[X_{12}X_{\bar{1}\bar{2}}]$ (40)
     $=\left(w+J\right)\mathsf{F}_{1,X\bar{X}}+\frac{J}{D}\sum_{v\cap u=\emptyset}\sum_{u}c_{v}c_{u}c[I]$
-    $=\left(w+J\right)\mathsf{F}_{1,X\bar{X}}+\frac{2J}{D}\mathsf{F}_{2}\ ,$
+    $=\left(w+J\right)\mathsf{F}_{1,X\bar{X}}+\frac{2J}{D}\mathsf{F}_{2}$
     where we have used the gauge redundancy: $c_{u}c_{v}c[\sigma X_{12}X_{\bar{1}\bar{2}}]=c_{u}c_{v}c[\sigma]$.
-Rank $r$ New Categories #$\#\mathsf{F}$
+Rank $r$ New Categories $\mathsf{F}$
 1 $\mathsf{F}_{0},\mathsf{F}_{1},\mathsf{F}_{0,X+\bar{X}}$ 1+4+2=7
@@
 Table 2: 8 categories for $n=2$.
-So far, we find that no new categories emerge. In summary, we list the total of 8 categories in Table 2.
+So far, we have found that no new categories have appeared. In summary, we list the total of 8 categories in Table 2.
 3.2 Observables
@@
 where $f_{a}(t)$ are given by
-$\begin{bmatrix}f_{1}(t)\\ f_{2}(t)\\ f_{3}(t)\\ f_{4}(t)\\ f_{5}(t)\\ f_{6}(t)\\ f_{7}(t)\\ f_{8}(t)\end{bmatrix}=\begin{bmatrix}0&0&\frac{1}{4}&\frac{1}{2}&\frac{1}{4}\\ 0&\frac{D^{2}-2}{D\left(D^{2}-4\right)}&\frac{-1}{4(D-2)}&\frac{-1}{2D}&\frac{-1}{4(D+2)}\\ 0&0&-\frac{1}{4}&0&\frac{1}{4}\\ 0&\frac{-1}{D^{2}-4}&\frac{1}{4(D-2)}&0&\frac{-1}{4(D+2)}\\ \frac{1}{D^{2}-1}&\frac{-2}{D^{2}-4}&\frac{1}{2(D-1)(D-2)}&0&\frac{1}{2(D+1)(D+2)}\\ 0&0&\frac{1}{4}&-\frac{1}{2}&\frac{1}{4}\\ \frac{-1}{D^{3}-D}&\frac{4}{D(D^{2}-4)}&\frac{-1}{2(D-1)(D-2)}&0&\frac{1}{2(D+1)(D+2)}\\ 0&\frac{2}{D(D^{2}-4)}&\frac{-1}{4(D-2)}&\frac{1}{2D}&\frac{-1}{4(D+2)}\end{bmatrix}\times\begin{bmatrix}1\\ e^{-Jt}\\ e^{-(2-2D)^{-1}Jt}\\ e^{-2Jt}\\ e^{-(2+2D)^{-1}Jt}\end{bmatrix}\ .$ (58)
-Generally, we can calculate OTOC by using the expression of (57)
-$\text{Tr}\left(O_{1}(t)O_{2}O_{3}(t)O_{4}\right)=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\equiv\sum_{a=1}^{8}f_{a}(t)\times\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\right)_{O_{1},O_{2},O_{3},O_{4}}\ .$ (59)
-A special case would be the noise-free case, which means that the noise strength $J$ vanishes.
-We expect that the time evolution will be simplified into a unitary one.
-Note that from (58) we can directly find that all $f_{a}(t)$ vanishes except $f_{1}(t)=1$.
-Hence
+$\begin{bmatrix}f_{1}(t)\\ f_{2}(t)\\ f_{3}(t)\\ f_{4}(t)\\ f_{5}(t)\\ f_{6}(t)\\ f_{7}(t)\\ f_{8}(t)\end{bmatrix}=\begin{bmatrix}0&0&\frac{1}{4}&\frac{1}{2}&\frac{1}{4}\\ 0&\frac{D^{2}-2}{D\left(D^{2}-4\right)}&\frac{-1}{4(D-2)}&\frac{-1}{2D}&\frac{-1}{4(D+2)}\\ 0&0&-\frac{1}{4}&0&\frac{1}{4}\\ 0&\frac{-1}{D^{2}-4}&\frac{1}{4(D-2)}&0&\frac{-1}{4(D+2)}\\ \frac{1}{D^{2}-1}&\frac{-2}{D^{2}-4}&\frac{1}{2(D-1)(D-2)}&0&\frac{1}{2(D+1)(D+2)}\\ 0&0&\frac{1}{4}&-\frac{1}{2}&\frac{1}{4}\\ \frac{-1}{D^{3}-D}&\frac{4}{D(D^{2}-4)}&\frac{-1}{2(D-1)(D-2)}&0&\frac{1}{2(D+1)(D+2)}\\ 0&\frac{2}{D(D^{2}-4)}&\frac{-1}{4(D-2)}&\frac{1}{2D}&\frac{-1}{4(D+2)}\end{bmatrix}\times\begin{bmatrix}1\\ e^{-Jt}\\ e^{-(2-2D)^{-1}Jt}\\ e^{-2Jt}\\ e^{-(2+2D)^{-1}Jt}\end{bmatrix}$ (52)
+Generally, we can calculate
+$\text{Tr}\left(O_{1}(t)O_{2}O_{3}(t)O_{4}\right)=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\equiv\sum_{a=1}^{8}f_{a}(t)\times\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\right)_{O_{1},O_{2},O_{3},O_{4}}\ .$ (53)
+The noise-free case $J=0$, all $f_{a}(t)$ vanishes except $f_{1}(t)=1$, one can find
 $\text{Tr}\left(O_{1}(t)O_{2}O_{3}(t)O_{4}\right)=\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{0}\right)_{O_{1},O_{2},O_{3},O_{4}}\ .$ (54)
-which is just the unitary time evolution for a closed system and finishes our consistency check.
-Thus, we will be more confident to use (59) to calculate the fluctuations of two-point functions, spectral form factor (SFF) and other observables.
-To end this section, we would mention that there is an alternative method Guo:2024zmr to evaluate $\mathcal{U}_{2}(t)$ by observing that
+which is just the unitary time evolution for a closed system. Using Eq.(53), we can calculate the fluctuations of two-point functions, spectral form factor (SFF) and other obervables.
+There is an alternative method Guo:2024zmr to evaluate $\mathcal{U}_{2}(t)$ by observing that
 $[\mathcal{L}_{n},V^{\otimes n}\otimes V^{*\otimes n}]=0,~\forall V\in U(D)\ .$ (55)
-We can always decompose a wavefunction $\psi=\psi_{i_{1}i_{2}}^{j_{1}j_{2}}\ket{i_{1},j_{1},i_{2},j_{2}}_{1\bar{1}2\bar{2}}$ into two singlets $\mathbf{1}$, four adjoints $\mathbf{D}^{2}-\mathbf{1}$, an antisymmetric traceless $\mathbf{A}$, two mixed $\mathbf{M}$ and a symmetric traceless $\mathbf{S}$ rank $(2,2)$ irreducible representations of $\text{SU}(D)$, with all the coefficients.
-Thus, we can also do the exponentiation on this basis.
-However, notice that the dimension of the representation is $10$, which is larger than the number of independent graph categories $8$, indicating that there are still some redundancies in the Young tableaux method.
-This might cause a longer processing time in numerical calculations.
+[…]
 4 Time evolution for $n=3$
-Then we consider $n=3$, so we have six sites: $\{1,\bar{1},2,\bar{2},3,\bar{3}\}$. The permutation group is $S_{3}\otimes S_{3}$, $S_{3}=\{I,X_{12},X_{13},X_{23},X_{123},X_{132}\}$, then
+Then we consider $n=3$, so we have six contours: $\{1,\bar{1},2,\bar{2},3,\bar{3}\}$. The permutation group is $S_{3}\otimes S_{3}$, $S_{3}=\{I,X_{12},X_{13},X_{23},X_{123},X_{132}\}$, then
 […]
 […]
-$p$ $\mathsf{F}_{a}$$\#\mathsf{F}$ $I$ $X+\bar{X}$ $X\bar{X}$ $XX+\bar{X}\bar{X}$ $XX\bar{X}+\bar{X}\bar{X}X$ $XX\bar{X}\bar{X}$
+$p$ $\mathsf{F}_{a}$ $\mathsf{F}$ II $\bar{X} XX$$\bar{X}$ $\bar{X}\bar{X} XXX¯+X$$\bar{X}+\bar{X}\bar{X}X$ $XX\bar{X}\bar{X}$
 0 1 6 9 4 12 4
@@
 4.1 Grouping graphs into categories
-Here we first define some notations:
+Here we first define some denotations:
 $X\equiv\{X_{12},X_{13},X_{23}\},~[XX]\equiv\{X_{123},X_{132}\}\ .$ (57)
 As discussed earlier, the number of individual graphs with $p$ connected pairs is given by $N_{n,p}=(n!C_{n}^{p})^{2}$. For $n=3$, it is
-$N_{3,p}=\{36,324,324,36\}\ ,\quad\text{for }p=\{0,1,2,3\}\ \ .$ (64)
-Next, we discuss the graph categories $\mathsf{F}_{p,\alpha}$ with $p$ connected pairs.
-For a graph with $p$ pairs, notice that the second term in (23) ($\sum_{u\cap C[\mathsf{F}]=\emptyset}c_{u}\mathsf{F}$) represents graphs with $p+1$ pairs, while the last two terms indicate applying additional 2 swaps on $\mathsf{F}$, which are excluded by $C(\mathsf{F})$.
+$N_{3,p}=\{36,324,324,36\}\ .$ (58)
+Next, we discuss the categories $\mathsf{F}_{p,\alpha}$ for $p=\{0,1,2,3\}$.
+For a graph with $p$ pairs, observe that the second term in Eq.(22) ($\sum_{u\cap C[\mathsf{F}]=\emptyset}c_{u}\mathsf{F}$) represents graphs with $p+1$ pairs, while the last two terms indicate applying additional 2-cycles, which are excluded by $C(\mathsf{F})$.
   • For $p=0$, there are no restrictions on $X$.
-  • For $p=1$, with $C[\mathsf{F}]=c_{u}$, we should require $X\cap u=\emptyset$.
-We can define the allowed set $X^{(0)}=X^{(ex)}[u]\equiv\{X|X\cap u\neq\emptyset\}$ and its complement as $X^{(1)}$.
-For example, if $u=(1\bar{1})$, then we have $X^{(0)}=\{X_{23}\}$ and $X^{(1)}=X-X^{(0)}=\{X_{12},X_{13}\}$.
+  • For $p=1$, with $C[\mathsf{F}]=c_{u}$, we require $X\cap u=\emptyset$.
+We denote this as $X^{(ex)}[u]$, or simply $X^{(0)}$.
+Similarly, we can define $X^{(0)}=X^{(ex)}[u]\equiv\{X|X\cap u\neq\emptyset\}$.
+For example, if $u=(1\bar{1})$, then $X^{(0)}=\{X_{23}\}$ and $X^{(1)}=X-X^{(0)}=\{X_{12},X_{13}\}$.
   • For $p>1$, the condition $X\cap u=\emptyset$ cannot be satisfied when $n=3$.
-While the discussion of graph categories by connected pairs is complete for $n=3$, we could generalize the definitions of $X^{(0)}$ and $X^{(1)}$ to $X^{(\alpha)}$ for generic $n$ and $p$.
-We define the conditions set as follows
+We can generalize the definitions of $X^{(0)}$ and $X^{(1)}$ for a general $p$. We define the conditions set as follows
 $\mathsf{Cons}\equiv\otimes^{p}\{ex,in\}\ .$ (59)
 For simplicity, we just denote $ex=0,in=1$, and define
-$\text{Con}(0,A,u)\equiv(A\cap u=\emptyset),~{}\text{Con}(1,A,u)\equiv(A\cap u\not=\emptyset)\ ,$ (66)
-showing that whether $u$ intersects with $A$ or not. Hence for a series of certain connected pairs $u_{1},u_{2},\ldots,u_{p}$, $X^{(\alpha)}$ is a set of 2-swap operators:
+[…]
+and for a certain connected pair $c(u_{1},u_{2},\ldots,u_{p})$, $X^{(\alpha)}$ is a set of two-cycles:
 $X^{(\alpha)}=\{X_{e}|\cap_{j=1}^{p}\text{Con}(\alpha[j],e,u_{j})\}\ .$ (61)
-In our case, for instance, we take $n=3,p=2$, then the condition set is $\textsf{Cons}=\{00,01,10,11\}$. And if we further choose $u_{1}=(1\bar{1}),u_{2}=(2\bar{2})$, we have
+Here we take $n=3,p=2$ and $u_{1}=(1\bar{1}),u_{2}=(2\bar{2})$, then the condition set is $\textsf{Cons}=\{00,01,10,11\}$, so
 $X^{(00)}=\emptyset,X^{(01)}=\{X_{23}\},X^{(10)}=\{X_{13}\},X^{(11)}=\{X_{12}\}\ .$ (62)
-Equipped with the new categories $\mathsf{F}_{0,I}$, we could calculate the operations of $\mathcal{L}_{n}^{r}$ on those categories.
-We will find that $\mathcal{L}_{n}^{r}\mathsf{F}_{0,I}$ can be expressed as a linear combination of graphs represented by permutations
+Starting with $\mathsf{F}_{0,I}$, we will find that $\mathcal{L}_{n}^{r}\mathsf{F}_{0,I}$ can be expressed as a linear combination of
 $\sum_{u_{1},u_{2},\ldots,u_{p}}c(u_{1},u_{2},\ldots,u_{p})c\left[\prod_{j=1}^{m}(X+\bar{X})^{(\alpha_{j})}\right]$ (63)
-where the product should be understood as the product between sets, which means the repeated terms should be eliminated.
-And the summation runs over all possible permutations.
-We will primarily discuss how to assign these blocks (69) into the appropriate graph categories in the followings.
-Zero Pair: $C[\mathsf{F}]=0$, #$\#\mathsf{F}=36$
-For $p=0$, (69) becomes
+where the product in square brackets should be understood as the product between permutation sets, which we then sum over all permutations.
+The condition indices $\alpha_{j}$ belong to the conditions set.
+Next, we will primarily discuss how to classify these blocks (Eq.(63)) into the appropriate graph categories.
+Zero pair: $C[\mathsf{F}]=0$, $\mathsf{F}=36$
+For $p=0$, Eq.(63) becomes
 $c\left[\prod_{j=1}^{m}(X+\bar{X})\right],m=0,1,2,3,\ldots\ .$ (64)
-Notice that for $n=3$, we have the following relation for permutation sets
-$XX=3I+3[XX],~{}[XX]X=2X\ .$ (71)
-Hence all terms appearing in $(X+\bar{X})^{m}$ are
+Notice that there is a relation for $n=3$
+$XX=3I+3[XX],~[XX]X=2X\ ,$ (65)
+so all terms appearing in $(X+\bar{X})^{m}$ are
 $I,X+\bar{X},X\bar{X},[XX]+[\bar{X}\bar{X}],[XX]\bar{X}+[\bar{X}\bar{X}]X,[XX][\bar{X}\bar{X}]\ .$ (66)
-Since the gauge group vanishes, there is no redundancy. Hence we obtain $6$ categories
-$\mathsf{F}_{0,I},\mathsf{F}_{0,X+\bar{X}},\mathsf{F}_{0,X\bar{X}},\mathsf{F}_{0,[XX]+[\bar{X}\bar{X}]},\mathsf{F}_{0,[XX]\bar{X}+[\bar{X}\bar{X}]X},\mathsf{F}_{0,[XX][\bar{X}\bar{X}]}\ ,$ (73)
+Since the gauge group vanish here, there is no redundancy, then we obtain $6$ categories
+$\mathsf{F}_{0,I},\mathsf{F}_{0,X+\bar{X}},\mathsf{F}_{0,X\bar{X}},\mathsf{F}_{0,[XX]+[\bar{X}\bar{X}]},\mathsf{F}_{0,[XX]\bar{X}+[\bar{X}\bar{X}]X},\mathsf{F}_{0,[XX][\bar{X}\bar{X}]}$ (67)
 where $F_{0,A}\equiv c[A]=\cup_{x\in A}c[x]$, or explicitly:
-$\mathsf{F}_{0,I}\equiv c[I],\quad\mathsf{F}_{0,X+\bar{X}}\equiv\sum_{e}c[X_{e}]+\sum_{\bar{e}}c[X_{\bar{e}}]+(X\leftrightarrow\bar{X})\ ,$ (74)
-$\mathsf{F}_{0,X\bar{X}}\equiv\sum_{u,\bar{e}}c[X_{\bar{e}}X_{u}]+(X\leftrightarrow\bar{X})\ ,$
-$\mathsf{F}_{0,[XX]+[\bar{X}\bar{X}]}\equiv c[X_{123}]+c[X_{132}]+c[X_{\bar{1}\bar{2}\bar{3}}]+c[X_{\bar{1}\bar{3}\bar{2}}]+(X\leftrightarrow\bar{X})\ ,$
-$\mathsf{F}_{0,[XX]\bar{X}+[\bar{X}\bar{X}]X}\equiv\sum_{e}c\left[X_{\bar{1}\bar{2}\bar{3}}X_{e}+X_{\bar{1}\bar{3}\bar{2}}X_{e}\right]+\sum_{\bar{e}}c\left[X_{123}X_{\bar{e}}+X_{132}X_{\bar{e}}\right]+(X\leftrightarrow\bar{X})\ ,$
-$\mathsf{F}_{0,[XX][\bar{X}\bar{X}]}\equiv c\left[X_{123}X_{\bar{1}\bar{2}\bar{3}}\right]+c\left[X_{123}X_{\bar{1}\bar{3}\bar{2}}\right]+c\left[X_{132}X_{\bar{1}\bar{2}\bar{3}}\right]+c\left[X_{132}X_{\bar{1}\bar{3}\bar{2}}\right]+(X\leftrightarrow\bar{X})\ .$
-It is straightforward to count their sizes, which are $\{1,6,9,4,12,4\}$ respectively. Thus, we have $1+6+9+4+12+4=36$ graphs, as we have calculated in (64).
-One Pair: $C[\mathsf{F}]=1$, #$\#\mathsf{F}=324$
-For $p=1$, (69) becomes
+$\mathsf{F}_{0,I}$ $\equiv c[I],\mathsf{F}_{0,X}\equiv\sum_{e}c[X_{e}]+\sum_{\bar{e}}c[X_{\bar{e}}]\ ,$ (68)
+$\mathsf{F}_{0,X\bar{X}}$ $\equiv\sum_{u,\bar{e}}c[X_{\bar{e}}X_{u}],\mathsf{F}_{0,XX}\equiv c[X_{123}]+c[X_{132}]+c[X_{\bar{1}\bar{2}\bar{3}}]+c[X_{\bar{1}\bar{3}\bar{2}}]\ ,$
+$\mathsf{F}_{0,XX\bar{X}}$ $\equiv\sum_{e}c\left[X_{\bar{1}\bar{2}\bar{3}}X_{e}+X_{\bar{1}\bar{3}\bar{2}}X_{e}\right]+\sum_{\bar{e}}c\left[X_{123}X_{\bar{e}}+X_{132}X_{\bar{e}}\right]\ ,$
+$\mathsf{F}_{0,XX\bar{X}\bar{X}}$ $\equiv c\left[X_{123}X_{\bar{1}\bar{2}\bar{3}}\right]+c\left[X_{123}X_{\bar{1}\bar{3}\bar{2}}\right]+c\left[X_{132}X_{\bar{1}\bar{2}\bar{3}}\right]+c\left[X_{132}X_{\bar{1}\bar{3}\bar{2}}\right]\ .$
+Without introducing complexities, we neglect the dual part obtained by $X\leftrightarrow\bar{X}$ and simply denote the categories containing $[XX]$ and $\bar{X}\bar{X}]$ as $\mathsf{F}_{0,XX}$, $\mathsf{F}_{0,XX\bar{X}}$, and $\mathsf{F}_{0,XX\bar{X}\bar{X}}$.
+It is straightforward to count their sizes: $\{1,6,9,4,12,4\}$.
+Thus, we have $1+6+9+4+12+4=36$, as required.
+One Pair: $C[\mathsf{F}]=1$$\mathsf{F}=324$
+For $p=1$, Eq.(63) becomes
 $\sum_{u}c_{u}\left[\prod_{j=1}^{m}(X+\bar{X})^{(\alpha_{j})}\right],m=0,1,2,3,\ldots,\alpha_{j}\in\{(0),(1)\}\ .$ (69)
-The size of the gauge group is just $p!=1$, so there is no redundancy. For arbitrary $u=(i\bar{j})$, we find that
-#$\#X^{(0)}=1,~{}\#X^{(1)}=2,~{}X=X^{(0)}+X^{(1)}\ .$ (76)
-Hence we have
+The size of gauge group is just $p!=1$, so there is no redundancy. For arbitrary $u=(i\bar{j})$, we find
+$\#X^{(0)}=1,~\#X^{(1)}=2,~X=X^{(0)}+X^{(1)}\ ,$ (70)
+so we have
 $X^{(0)}X^{(0)}=\{I\},~X^{(0)}X^{(1)}=X^{(1)}X^{(0)}=[XX],~X^{(1)}X^{(1)}=2I+2[XX]\ .$ (71)
@@
 $\mathsf{F}_{1,I},\mathsf{F}_{1,X}^{(0)},\mathsf{F}_{1,X}^{(1)},\mathsf{F}_{1,X\bar{X}}^{(00)},\mathsf{F}_{1,X\bar{X}}^{(01)},\mathsf{F}_{1,X\bar{X}}^{(11)},\mathsf{F}_{1,XX},\mathsf{F}_{1,XX\bar{X}}^{(0)},\mathsf{F}_{1,XX\bar{X}}^{(1)},\mathsf{F}_{1,XX\bar{X}\bar{X}}$ (73)
-with there sizes to be $\{1,18,36,9,36,36,36,36,72,36\}$.
-Two Pairs: $C[\mathsf{F}]=2$,#$\#\mathsf{F}=324$
-For $p=2$, (69) becomes
+with sizes $\{1,18,36,9,36,36,36,36,72,36\}$.
+Two pairs: $C[\mathsf{F}]=2$$\mathsf{F}=324$
+For $p=2$, Eq.(63) becomes
 $\sum_{u,v}c_{u}c_{v}\left[\prod_{j=1}^{m}(X+\bar{X})^{(\alpha_{j})}\right],m=0,1,2,3,\ldots,\alpha_{j}\in\{(00),(01),(10),(11)\}\ .$ (74)
-$\forall u,v,\ u\cap v=\emptyset$, we have
+For any two pairs $u\cap v=\emptyset$, we have
 $X^{(00)}=\emptyset,X=\{X^{(01)},X^{(10)},X^{(11)}\}\ .$ (75)
-So all terms which can appear are
+So all terms appearing are
 $I,X^{(01)+(10)}+\bar{X}^{(01)+(10)},X^{(11)}+\bar{X}^{(11)},X^{(01)}\bar{X}^{(01)}+X^{(10)}\bar{X}^{(10)},$ (76)
@@
 But for $p=2$, the gauge group generated by $X^{(11)}\otimes\bar{X}^{(11)}$ has size $p!=2$. We have the equivalent relation
-$c_{u}c_{v}c[\sigma]=c_{u}c_{v}c\left[\sigma\cdot X^{(11)}\bar{X}^{(11)}\right]\ ,$ (83)
+$c_{u}c_{v}c[\sigma]=c_{u}c_{v}c\left[\sigma\cdot X^{(11)}\bar{X}^{(11)}\right]\ .$ (77)
 which means $X^{(11)}\bar{X}^{(11)}\sim I$, and furthermore
@@
 $\mathsf{F}_{2,I},\mathsf{F}_{2,X}^{(01)},\mathsf{F}_{2,X}^{(11)},\mathsf{F}_{2,X\bar{X}}^{(01,01)},\mathsf{F}_{2,X\bar{X}}^{(01,10)},\mathsf{F}_{2,X\bar{X}}^{(01,11)},\mathsf{F}_{2,XX\bar{X}}\ .$ (79)
-Their sizes are $\{18,72,18,36,36,72,72\}$.
-Three Pairs: $C[\mathsf{F}]=3$,#$\#\mathsf{F}=36$
-For $p=2$, (69) becomes
+Their size are $\{18,72,18,36,36,72,72\}$.
+Three pairs: $C[\mathsf{F}]=3$$\mathsf{F}=36$
+For $p=2$, Eq.(63) becomes
 $\sum_{u_{1},u_{2},u_{3}}c_{u_{1}}c_{u_{2}}c_{u_{3}}\left[\prod_{j=1}^{m}(X+\bar{X})^{(\alpha_{j})}\right],m=0,1,2,3,\ldots$ (80)
-For three pairs $u_{1},u_{2},u_{3}$, we can only take three conditions
+For three pairs $u_{1},u_{2},u_{2}$, we can only take three conditions
 $\alpha_{j}\in\{(011),(101),(110)\}\ .$ (81)
-Any permutation $X_{e}X_{\bar{e}}$ is considered as a redundancy, allowing us to fix the order of $\bar{1},\bar{2},\bar{3}$.
+Any permutation $X_{e}X_{\bar{e}}$ is considered a redundancy, allowing us to fix the order of $\bar{1},\bar{2},\bar{3}$.
 This means we only need to consider permutations of $1,2,3$.
-Utilizing the fact that $X=X^{(011)}+X^{(101)}+X^{(110)}$, we identify three building blocks: $I,X,[XX]$.
+Utilizing the fact that $X=X^{(011)}+X^{(101)}+X^{(110)}$, we identify the three building blocks: $I,X,[XX]$.
 The respective graph categories are:
-$\mathsf{F}_{3,I}=\cup_{u_{1},u_{2},u_{3}}c[I],~{}\mathsf{F}_{3,X}=\cup_{u}c_{u_{1}}c_{u_{2}}c_{u_{3}}c[X],~{}\mathsf{F}_{3,XX}=\cup_{u_{1},u_{2},u_{3}}c_{u_{1}}c_{u_{2}}c_{u_{3}}c[[XX]]\ .$ (88)
-They have sizes of $\{6,18,12\}$, respectively.
-4.2 Operation of $\mathcal{L}_{3}$
+$\mathsf{F}_{3,I}=\cup_{u_{1},u_{2},u_{3}}c[I],\mathsf{F}_{3,X}=\cup_{u}c_{u_{1}}c_{u_{2}}c_{u_{3}}c[X],\mathsf{F}_{3,XX}=\cup_{u_{1},u_{2},u_{3}}c_{u_{1}}c_{u_{2}}c_{u_{3}}c[[XX]]\ .$ (82)
+They have sizes $\{6,18,12\}$.
+4.2 Action of $\mathcal{L}_{3}$
 Using the relations for $n=3$
 $XX$ $=3I+3[XX],~[XX]X=2X,$ (83)
-$XX_{e}$ $=X_{e}+[XX],~{}[XX]X_{e}=X_{e}[XX]=X-X_{e}\ ,$
-and the gauge symmetry discussed earlier, it is straightforward to calculate the operation of $\mathcal{L}_{3}$ on the graph categories defined in the previous section.
-For simplicity, we define $\mathcal{L}_{n}^{\prime}$ as the non-trivial part of $\mathcal{L}_{n}$:
+$XX_{e}$ $=X_{e}+[XX],~[XX]X_{e}=X_{e}[XX]=X-X_{e}\ .$
+and the gauge symmetry discussed earlier, it is straightforward to calculate the action of $\mathcal{L}_{3}$ on the graph categories defined in the previous section.
+For simplicity, we define $\mathcal{L}_{n}^{\prime}$ as the non-trivial action of $\mathcal{L}_{n}$:
 $\mathcal{L}_{n}\mathsf{F}_{p,\alpha}=(w+pJ)\mathsf{F}_{p,\alpha}+\mathcal{L}_{n}^{\prime}\mathsf{F}_{p,\alpha}\ .$ (84)
-With some calculations one can get:
+It is not hard to obtain:
 Zero pair: $\mathsf{F}_{0,I},\mathsf{F}_{0,X},\mathsf{F}_{0,X\bar{X}},\mathsf{F}_{0,XX},\mathsf{F}_{0,XX\bar{X}},\mathsf{F}_{0,XX\bar{X}\bar{X}}$
@@
 Three pairs: $\mathsf{F}_{3,I},\mathsf{F}_{3,X},\mathsf{F}_{3,XX}$
-$\mathcal{L}_{3}\mathsf{F}_{3,\beta}=(w+3J)\mathsf{F}_{3,\beta},\quad\beta=I,X,XX\ .$ (94)
+$\mathcal{L}_{3}\mathsf{F}_{3,\beta}=(w+3J)\mathsf{F}_{3,\beta},\beta=I,X,XX\ .$ (88)
 4.3 Observables
-Recall the factor $\mathsf{w}=-\text{i}\mathsf{E}-nJ$ from (13), then we have $M=-\text{i}\mathsf{E}\mathbf{1}+M_{J}$. Here $\mathbf{1}$ denotes identity matrix. $M_{J}$ has the explicit form
+We write $\mathsf{w}=-\text{i}\mathsf{E}-nJ$, then $M=-\text{i}\mathsf{E}\mathbf{1}+M_{J}$, $M_{J}$ has the explicit form
 $\left(\begin{array}[]{cccccccccccccccccccccccccc}-3J&-\frac{6J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ -\frac{J}{D}&-3J&-\frac{3J}{D}&-\frac{2J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&-\frac{2J}{D}&-3J&0&-\frac{4J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&-\frac{3J}{D}&0&-3J&-\frac{3J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&-\frac{3J}{D}&-\frac{J}{D}&-3J&-\frac{2J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&-\frac{6J}{D}&-3J&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ \frac{J}{D}&0&0&0&0&0&-2J&-\frac{2J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&\frac{J}{D}&0&0&0&0&-\frac{J}{D}&-2J&0&-\frac{J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&\frac{J}{D}&0&0&0&0&0&0&-2J&0&-\frac{J}{D}&0&-\frac{J}{D}&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&\frac{J}{D}&0&0&0&0&-\frac{2J}{D}&0&-2J&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&\frac{J}{D}&0&0&0&0&0&-\frac{J}{D}&0&-2J&0&0&-\frac{J}{D}&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&\frac{J}{D}&0&0&0&0&0&0&0&0&-2J&0&0&-\frac{2J}{D}&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&\frac{J}{D}&0&0&0&0&-\frac{J}{D}&0&0&0&-2J&-\frac{J}{D}&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&\frac{J}{D}&0&0&0&0&0&-\frac{J}{D}&0&-\frac{J}{D}&-2J&0&0&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&\frac{J}{D}&0&0&0&0&0&0&-\frac{J}{D}&0&0&-2J&-\frac{J}{D}&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&\frac{J}{D}&0&0&0&0&0&0&0&0&-\frac{2J}{D}&-2J&0&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&\frac{2J}{D}&0&0&0&0&\frac{2J}{D}&0&0&0&0&-J&0&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&\frac{J}{D}&\frac{J}{D}&0&0&0&0&0&\frac{2J}{D}&0&0&-J&0&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&\frac{4J}{D}&0&0&0&0&0&0&0&0&0&-J&0&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&\frac{J}{D}&0&\frac{J}{D}&0&0&0&\frac{2J}{D}&0&0&0&-J&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&\frac{2J}{D}&0&0&0&0&\frac{2J}{D}&0&0&0&0&-J&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&\frac{J}{D}&\frac{J}{D}&\frac{2J}{D}&0&0&0&0&0&0&0&0&-J&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&\frac{2J}{D}&\frac{2J}{D}&0&0&0&0&0&0&0&-J&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&\frac{3J}{D}&0&0&\frac{6J}{D}&0&0&0&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&\frac{4J}{D}&\frac{J}{D}&0&0&0&\frac{4J}{D}&0&0&0\\ 0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&0&\frac{3J}{D}&\frac{6J}{D}&0&0&0&0\\ \end{array}\right)\ .$ (115)
-It is straightforward to obtain the spectrum of $M_{J}$:
-[…]
+It is direct to obtain the spectrum of $M_{J}$:
+$\Bigg\{$ $0,0,0,-\frac{3(D-2)J}{D},-\frac{3(D-1)J}{D},-\frac{2(D-1)J}{D},-\frac{2(D-1)J}{D},-\frac{2(D-1)J}{D},$ (116)
 $-3J,-3J,-2J,-2J,-2J,-2J,-J,-J,-J,-J,-J,-J,-J,$
 $-\frac{3(D+1)J}{D},-\frac{2(D+1)J}{D},-\frac{2(D+1)J}{D},-\frac{2(D+1)J}{D},-\frac{3(D+2)J}{D}\Bigg\}\ .$
-Then we can find the analytical expression of $\mathcal{U}_{3}(t)$ as
-$\mathcal{U}_{3}(t)=e^{\mathcal{L}_{3}t}=e^{-\text{i}\mathsf{E}t}\sum_{a=1}^{26}f_{a}(t)\mathsf{F}_{a}\ .$ (123)
+Then we can find the analytical expression of $\mathcal{U}_{3}(t)$
+$e^{\mathcal{L}_{3}t}=e^{-\text{i}\mathsf{E}t}\sum_{a=1}^{26}f_{a}(t)\mathsf{F}_{a}\ .$ (117)
 We plot $f_{a}(t)$ in Fig.3, while the explicit expressions are given in Appendix A.
-Figure 3: The pattern coefficient $f_{a}(t)$ and its log-plot, where we set $D=10$ and $J=1$.
-We depict the categories of $p=0,1,2,3$ with red, orange, green, and blue colors, respectively.
-One can see that the three categories with $p=3$ have energy 00, so they (the blue lines) remain finite at time $t=\infty$.
-Generally, we can calculate the correlation functions as
-$\text{Tr}\left[O_{1}(t)O_{2}O_{3}(t)O_{4}O_{5}(t)O_{6}\right]=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\equiv\sum_{a=1}^{26}f_{a}(t)\times\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\right)_{O_{1},O_{2},O_{3},O_{4},O_{5},O_{6}}$ (124)
-where we have used $O(t)\equiv U^{\dagger}OU$ and the graph representation for an operator $A$
-$A_{ij}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\ .$ (125)
-To calculate $\mathsf{F}_{a}[O_{1},O_{2},O_{3},O_{4},O_{5},O_{6}]$, in principle, we need to contract each individual graph in the category with the operator tensor in the way as shown in (124).
-The expression $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$ has told us the connection between points in the graph.
-We treat the connections in the graph as free propagators and the operators as bridges between points $i,\bar{i}$ and $\bar{i}^{\prime},\overline{(i+1)}^{\prime}$, where the prime means that the point is on the right-hand side of the graph.
+Figure 3: The pattern coefficient $f_{a}(t)$ or its log-plot, where we set $D=10,J=1$.
+We depict the categories of $p=0,1,2,3$ with red, orange, green and blue color respectively.
+One can see the three categories with $p=3$ has the energy 0, so they (blue lines) reminds finite at time $t=\infty$.
+Generally, we can calculate the observables
+$\text{Tr}\left[O_{1}(t)O_{2}O_{3}(t)O_{4}O_{5}(t)O_{6}\right]=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\equiv\sum_{a=1}^{26}f_{a}(t)\times\left(e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}\right)_{O_{1},O_{2},O_{3},O_{4},O_{5},O_{6}}$ (118)
+where we have used $O(t)\equiv U^{\dagger}OU$ and the graph representation for an operator AA
+$A_{ij}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\ .$ (119)
+To calculate $\mathsf{F}_{a}[O_{1},O_{2},O_{3},O_{4},O_{5},O_{6}]$, in principle, we need to contract each individual graph in the category with the operator tensor in the way as shown in Eq.(118).
+The denotation $\mathsf{F}=\prod_{i=1}^{p}c_{a_{i}\bar{b}_{i}}c[\sigma]$ has told us the connection between points in the graph, we treat the connections on the graph as free path and the operators as bridges which connect points $i,\bar{i}$ and $\bar{i}^{\prime},\overline{(i+1)}^{\prime}$, here prime means the point on the right hand of the graph.
 So there is no need to draw the graph.
-We can just start with any points following the connections in the graph and form a closed loop.
-Hence we will have a trace of operators
+We can just start with any points following the connections in the graph then finally we will return back, then we will have an operator trace
 $\text{Tr}(O_{i}O_{j}O^{T}_{k}\cdots)\ .$ (120)
-If there are still points that we have not passed, we then start with those points, finding another closed loop until all points in the graph have been visited.
-The process may be a bit tedious, but it can be accomplished with some code.
-Alternatively, one can simply write down the explicit expression of a graph using the Kronecker delta function as defined in (16).
+If there remind points we do not pass, we then do the same thing until all points in the graph are passed.
+The process may be a little boring, but it can be achieved by some code.
+Or one can just write the explicit expression of a graph with Kronecker delta function as defined in Eq.(15).
 Moreover, with the expression of $\mathcal{U}_{1,2,3}(t)$, one can calculate the correlation function with general time order $\text{Tr}\left[O_{1}(t_{1})O_{2}O_{3}(t_{2})O_{4}O_{5}(t_{3})O_{6}\right]$.
 For example, we consider the case $t_{1}<t_{2}<t_{3}$, then the ensemble-averaged time-evolution $\mathcal{U}_{3}(t_{1},t_{2},t_{3})$ is
-$\mathbb{E}\left(\mathbb{I}_{D}^{\otimes 2}\otimes\mathbb{I}_{D}^{\otimes 2}\otimes U(t_{32})\otimes U^{*}(t_{32})\right)\ \mathbb{E}\left(\mathbb{I}_{D}\otimes\mathbb{I}_{D}\otimes U(t_{21})^{\otimes 2}\otimes U^{*}(t_{21})^{\otimes 2}\right)\ \mathbb{E}\left(U(t_{1})^{\otimes 3}\otimes U^{*}(t_{1})^{\otimes 3}\right)$
-$=$ $\left[\mathbb{I}_{D}^{\otimes 2}\otimes\mathbb{I}_{D}^{\otimes 2}\otimes\mathcal{U}_{1}(t_{32})\right]\times\left[\mathbb{I}_{D}\otimes\mathbb{I}_{D}\otimes\mathcal{U}_{2}(t_{21})\right]\times\mathcal{U}_{3}(t_{1})\ ,$ (127)
-where $\mathbb{I}_{D}$ denotes the $D$-dimensional identity matrix.
-Finally, we simply need to contract $\mathcal{U}_{3}(t_{1},t_{2},t_{3})$ with the operator tensor as demonstrated in (124).
+$\mathbb{E}\left(1_{D}^{\otimes 2}\otimes 1_{D}^{\otimes 2}\otimes U(t_{32})\otimes U^{*}(t_{32})\right).\mathbb{E}\left(1_{D}\otimes 1_{D}\otimes U(t_{21})^{\otimes 2}\otimes U^{*}(t_{21})^{\otimes 2}\right).\mathbb{E}\left(U(t_{1})^{\otimes 3}\otimes U^{*}(t_{1})^{\otimes 3}\right)$
+$=$ $\left[1_{D}^{\otimes 2}\otimes 1_{D}^{\otimes 2}\otimes\mathcal{U}_{1}(t_{32})\right].\left[1_{D}\otimes 1_{D}\otimes\mathcal{U}_{2}(t_{21})\right].\mathcal{U}_{3}(t_{1})\ .$ (121)
+where $1_{D}$ denotes the $D$-dimensional identity matrix.
+Finally, we simply need to contract $\mathcal{U}_{3}(t_{1},t_{2},t_{3})$ with the operator tensor as demonstrated in Eq.(118).
 In this paper, we do not provide explicit expressions for the correlation functions.
-5 General Cases
-For general $n$, we have $2n$ sites $\{1,\bar{1},2,\bar{2},3,\bar{3},\ldots,n,\bar{n}\}$. The permutation group is $S_{n}\otimes S_{n}$. As discussed in $n=3$ case, we should define
+5 General case
+For general $n$, so we have $2n$ contours $\{1,\bar{1},2,\bar{2},3,\bar{3},\ldots,n,\bar{n}\}$. The permutation group is $S_{n}\otimes S_{n}$. As discussed before, we will meet with terms
+$\sum_{u_{1},u_{2},\ldots,u_{p}}c(u_{1},u_{2},\ldots,u_{p})c\left[\prod_{j=1}^{m}(X+\bar{X})^{(\boldsymbol{\alpha}_{j})}\right]$ (122)
+where recall $X^{(\boldsymbol{\alpha}_{j})}$ is a set of two-cycles submitting to the condition $\boldsymbol{\alpha}_{j}$, i.e.,
 $X^{(\boldsymbol{\alpha}_{j})}\equiv\left\{X_{e}|\cap_{k=1}^{p}\text{Con}(\boldsymbol{\alpha}_{j}[k],e,u_{k})\right\}\ .$ (123)
-$X^{(\boldsymbol{\alpha}_{j})}$ is a set of 2-swap operators submitting to the condition $\boldsymbol{\alpha}_{j}$. Our primary goal is to deal with the following graph expression
-$\sum_{u_{1},u_{2},\ldots,u_{p}}c(u_{1},u_{2},\ldots,u_{p})c\left[\prod_{j=1}^{m}(X+\bar{X})^{(\boldsymbol{\alpha}_{j})}\right]\ ,$ (129)
-and assign them into appropriate categories.
 Since it is free to change the order of $u_{1},u_{2},\ldots,u_{p}$, there is a $S_{p}$ permutation redundancy for the conditions:
 $\{\boldsymbol{\alpha}_{j}\}_{j=1,2,\ldots,p}\sim\{\sigma\boldsymbol{\alpha}_{j}\}_{j=1,2,\ldots,p},~\forall\sigma\in S_{p}\ .$ (124)
-In practice, we can always find a permutation $\sigma\in S_{n}$ as a product of cycles $\{\sigma_{i}\}_{i=1,2,\ldots,g}$ with length $1<l_{1}\leq l_{2}\ldots\leq l_{g}$ such that:
+In practice, we can always write a permutation $\sigma\in S_{n}$ as a product of cycles $\{\sigma_{i}\}_{i=1,2,\ldots,g}$ with length $1<l_{1}\leq l_{2}\ldots\leq l_{g}$:
 $X_{[l_{1},l_{2},\ldots,l_{g}]}\equiv\left\{X_{\sigma_{1}\sigma_{2}\cdots\sigma_{g}}||\sigma_{i}|=l_{i},\sigma_{i}\cap\sigma_{j}=0,i\not=j\right\}\ .$ (125)
-Then for a certain $p$, we first analyze all permutation categories appearing in the products $\prod_{j=1}^{m}X^{(\alpha_{j})},m=0,1,2,\ldots$, which gives us a division of the permutation group $S_{n}$
+Then for certain $p$, we first analyze all permutation categories appearing in the products $\prod_{j=1}^{m}X^{(\alpha_{j})},m=0,1,2,\ldots$, which gives a division of the permutation group $S_{n}$
 $S_{n}=\cup_{\alpha=1}^{N_{X}}X_{\alpha},X_{\alpha}\cap X_{\beta}=\emptyset,\forall\alpha\not=\beta\ .$ (126)
@@
 $\mathsf{F}_{p,X_{\alpha}\bar{X}_{\beta}}\equiv\cup_{x\in X_{\alpha}\bar{X}_{\beta}\cup X_{\beta}\bar{X}_{\alpha}}\cup_{u_{1},u_{2},\ldots,u_{p}}c(u_{1},u_{2},\ldots,u_{p})c[x]\ .$ (127)
-Finally when $p\geq 2$, we need to consider removing the permutation redundancy of the conditions and the gauge symmetry of the graph representation.
-In other words, we need to module some equivalent relations.
+Finally when $p\geq 2$, we need to consider removing the permutation redundancy of the conditions and the gauge symmetry of the graph representation, i.e, we need to module some equivalence relation.
 5.1 Example: $n=4$
-Here we take $n=4$ as an example. For different number of connected pairs, the number of graphs should be $N_{4,p}={\{576,9216,20736,9216,576\}}$. We then define
+Here we take $n=4$ as an example, then $N_{4,p}={\{576,9216,20736,9216,576\}}$. We then define
 $X_{[2]}\equiv X=\{X_{12},X_{13},X_{14},X_{23},X_{24},X_{34}\}\ .$ (128)
@@
 Zero Pair $p=0,\#\mathsf{F}=576$
-For $p=0$, there is no restriction on $X$, so we have the division of $S_{4}$ as
+Due to $p=0$, there is no constricts on $X$, so we have the division of $S_{4}$
 $S_{4}=I+X_{[2]}+X_{[3]}+X_{[2,2]}+X_{[4]},~\#S_{4}=1+6+8+3+6=24\ .$ (129)
-Since there is no permutation redundancy or gauge redundancy, we have ${N_{X}(N_{X}+1)\over 2}=15$ independent graph categories:
+Since there is no gauge redundancy and condition permutation redundancy, we have ${N_{X}(N_{X}+1)\over 2}=15$ graph categories:
 $\mathsf{F}_{0,I},\mathsf{F}_{0,X},\mathsf{F}_{0,X_{[3]}},\mathsf{F}_{0,X_{[2,2]}},\mathsf{F}_{0,X_{[4]}},\mathsf{F}_{0,X\bar{X}},\mathsf{F}_{0,X\bar{X}_{[3]}},\mathsf{F}_{0,X\bar{X}_{[2,2]}},\mathsf{F}_{0,X\bar{X}_{[4]}},\mathsf{F}_{0,X_{[3]}\bar{X}_{[3]}},$
@@
 For an arbitrary pair $u$, we have two conditions for $X$ and $X_{[3]}$, so that we have the division
-$S_{4}=I+X^{(0)}+X^{(1)}+X_{[3]}^{(0)}+X_{[3]}^{(1)}+X_{[2,2]}+X_{[4]}\equiv\cup_{\alpha=1}^{7}X_{\alpha}\ .$ (137)
-Since $p<2$, there is no permutation or gauge redundancy, so we have ${7\times 8\over 2}=28$ categories:
-$F_{1,X_{\alpha}\bar{X}_{\beta}}\equiv\sum_{u}\sum_{x\in X_{\alpha}\bar{X}_{\beta}\cup X_{\beta}\bar{X}_{\alpha}}c_{u}c[x].$ (138)
+$S_{4}=I+X^{(0)}+X^{(1)}+X_{[3]}^{(0)}+X_{[3]}^{(1)}+X_{[2,2]}+X_{[4]}\equiv\cup_{\alpha}X_{\alpha}\ .$ (131)
+Due to $p<2$, there is no redundancy, so we have ${7\times 8\over 2}=28$ categories: $F_{1,X_{\alpha}\bar{X}_{\beta}}\equiv\sum_{u,x\in X_{\alpha}\bar{X}_{\beta}}c_{u}c[x]$.
 Two Pairs $p=2,\#\mathsf{F}=20736$
 For any two pairs $u_{1},u_{2}$, we find
-$X=X^{(00)}+X^{(01)}+X^{(10)}+X^{(11)},~{}X_{[3]}^{(00)}=\emptyset,~{}X_{[3]}=X_{[3]}^{(01)}+X^{(10)}_{[3]}+X_{[3]}^{(11)}\ ,$ (139)
-so we obtain the division to be
+$X=X^{(00)}+X^{(01)}+X^{(10)}+X^{(11)},X_{[3]}^{(00)}=\emptyset,X_{[3]}=X_{[3]}^{(01)}+X^{(10)}_{[3]}+X_{[3]}^{(11)}\ ,$ (132)
+so we obtain the division
 $S_{4}=I+X^{(00)}+X^{(01)}+X^{(10)}+X^{(11)}+X_{[3]}^{(01)}+X_{[3]}^{(10)}+X_{[3]}^{(11)}+X_{[2,2]}+X_{[4]}\ .$ (133)
-Naively, we have ${10\times 11\over 2}=55$ categories. But here we need to remove the permutation redundancy. For an example, changing the order of conditions $\alpha_{j}$, we have
-$\mathsf{F}_{2,X}^{(01)}=\mathsf{F}_{2,X}^{(10)},~{}\mathsf{F}_{2,X\bar{X}}^{(00,01)}=\mathsf{F}_{2,X\bar{X}}^{(00,10)}\ .$ (141)
+Naively, we have ${10\times 11\over 2}=55$ categories. But here we need to remove the condition redundancy and gauge redundancy. For an example, changing the order of conditions, we have
+$\mathsf{F}_{2,X}^{(01)}=\mathsf{F}_{2,X}^{(10)},\mathsf{F}_{2,X\bar{X}}^{(00,01)}=\mathsf{F}_{2,X\bar{X}}^{(00,10)}\ .$ (134)
 And notice the symmetry between $X$ and $\bar{X}$, we have
 $\mathsf{F}_{2,X\bar{X}}^{(00,01)}=\mathsf{F}_{2,X\bar{X}}^{(01,00)}\ .$ (135)
-As for the gauge redundancy, here we impose the equivalence $X^{(11)}\bar{X}^{(11)}\sim I$, so immediately we have
+As for the gauge redundancy, here we impose the equivalence $X^{(11)}\bar{X}^{(11)}\sim I$, so directly we have
 $\mathsf{F}_{2,X\bar{X}}^{(11,11)}=\mathsf{F}_{2,I}\ .$ (136)
@@
 $X_{[3]}$ $=X_{[3]}^{(011)}+X_{[3]}^{(101)}+X_{[3]}^{(110)}\ ,$ (137)
-so we obtain the division as
+so we obtain the division
 $S_{4}$ $=I+X^{(001)}+X^{(010)}+X^{(100)}+X^{(011)}+X^{(101)}+X^{(110)}$
 $\hskip 15.0pt+X_{[3]}^{(011)}+X_{[3]}^{(101)}+X_{[3]}^{(110)}+X_{[2,2]}+X_{[4]}\ .$ (138)
-Naively, there are ${12\times 13\over 2}=96$ categories. The permutation redundancy is similar to $p=2$ case and the gauge group is generated by
+Naively, there are ${12\times 13\over 2}=96$ categories. The condition permutation redundancy is similar to $p=2$ case and the gauge group is generated by
 $X^{(011)}\bar{X}^{(011)},X^{(101)}\bar{X}^{(101)},X^{(110)}\bar{X}^{(110)}\ .$ (139)
@@
 $X_{[3]}$ $=X_{[3]}^{(0111)}+X_{[3]}^{(1011)}+X_{[3]}^{(1101)}+X_{[3]}^{(1110)}\ ,$ (140)
-which gives the division as
+which gives the division
 $S_{4}$ $=I+X^{(0011)}+X^{(0101)}+X^{(1001)}+X^{(0110)}+X^{(1010)}+X^{(1100)}$
 $\hskip 15.0pt+X_{[3]}^{(011)}+X_{[3]}^{(101)}+X_{[3]}^{(110)}+X_{[2,2]}+X_{[4]}\ .$ (141)
-Notice the gauge group is generated by each permutation like $X_{e}\bar{X}_{e}$. So we can fix the order of $\bar{i}$ to be $\bar{1}\bar{2}\bar{3}\bar{4}$, and we just need to consider the categories
+Notice the gauge group is generated by any permutation like $X_{e}\bar{X}_{e},\forall e$.
+So we can fix the order of $\bar{i}$ to be $\bar{1}\bar{2}\bar{3}\bar{4}$, then we just need to consider the categories
 $\mathsf{F}_{\alpha}\equiv\cup_{u_{1},u_{2},u_{3},u_{4}}c(u_{1},u_{2},u_{3},u_{4})c[X_{\alpha}]\ .$ (142)
-Notice the permutation redundancy, we have
+Notice the permutation redundancy of the condition, we have
 $X^{(0011)}\sim X^{(0101)}\sim X^{(1001)}\sim X^{(0110)}\sim X^{(1010)}\sim X^{(1100)},X_{[3]}^{(011)}\sim X_{[3]}^{(101)}\sim X_{[3]}^{(110)}\ ,$ (143)
-so we only have five categories left
+so we only have five categories
 $\mathsf{F}_{4,I},\mathsf{F}_{4,X},\mathsf{F}_{4,X_{[3]}},\mathsf{F}_{4,X_{[2,2]}},\mathsf{F}_{4,X_{[4]}}\ .$ (144)
-6 Discussions and Outlook
-In this paper we analyze the time evolution of the $n$-replica physics of the BGUE model in a graph-based framework.
-This approach proves to be particularly effective for lower values of $n$, as explicitly demonstrated in the $n=2$ and $n=3$ cases.
-We also provide a systematic strategy that extends to generic $n$.
-Our construction highlights how the increasing complexity of graph structures can be managed by organizing them into well-defined categories, which will simplify the calculations of two-point functions, OTOC, and other relevant observables.
-The general formulation outlined in Sec.
-5 suggests that this method can, in principle, be extended to arbitrary $n$.
-However, as $n$ increases, the growing number of graph structures presents significant challenges.
-Addressing these challenges will require the development of more efficient techniques for classifying, encoding, and manipulating graphs at large $n$.
-An important direction for future research is to explore algorithmic approaches to systematically handle this complexity.
-Moreover, further investigation into the implications of this framework for Brownian disordered systems and quantum information theory may uncover deeper structural insights into the role of replica dynamics in random matrix models.
+6 Discussion and Outlook
+The graph-based framework developed in this work provides a powerful approach for analyzing the $n$-replica physics of the BGUE model.
+The explicit constructions for the $n=2$ and $n=3$ cases demonstrate how the increasing complexity of the graphs can be managed by organizing them into distinct categories.
+The general formulation outlined in Section 5 suggests that this approach can be extended to arbitrary $n$, though the proliferation of graph structures poses a significant challenge.
+Developing efficient methods for packing and manipulating the graphs will be an important direction for future research.
 A key insight is that the factor $w$, which encodes the dependence on the system spectrum, commutes with the graph-based operators.
-This allows us to analyze the system with infinite-temperature GUE noise without introducing additional complexity.
-It is worth noting that the techniques developed in this paper can also be applied to other models, such as the Brownian Gaussian Orthogonal Ensemble (BGOE) and the Brownian Gaussian Symplectic Ensemble (BGSE).
-In particular, for BGOE, we have:
+This allows for the analysis of the system with infinite-temperature GUE noise without introducing additional complexity.
+The techniques discussed in this paper can be applied to other models, such as the Brownian Gaussian Orthogonal Ensemble (BGOE) and Brownian Gaussian Symplectic Ensemble (BGSE). For BGOE, we have
 $\mathcal{L}_{n}=\frac{-JD}{D+1}\left[n\mathbb{I}+\sum_{i<j}\left(P_{ij}+P_{\bar{i}\bar{j}}\right)-\sum_{i,\bar{j}}P_{i\bar{j}}\right]+\frac{-J}{D+1}\left[n+\sum_{i<j}X_{ij}+X_{\bar{i}\bar{j}}-\sum_{i\bar{j}}X_{i\bar{j}}\right]\ ,$ (145)
@@
 $\mathcal{L}_{n}=\frac{-J}{D-1}\left[2n(D-1)\mathbb{I}-D\sum_{i<j}\left(P_{ij}+P_{\bar{i}\bar{j}}\right)-2D\sum_{i,\bar{j}}P_{i\bar{j}}+4\sum_{i<j}\left(X_{ij}+X_{\bar{i}\bar{j}}\right)+2\sum_{i,\bar{j}}X_{i\bar{j}}\right]\ .$ (146)
-We find there are new operators such as $P_{ij},P_{\bar{i}\bar{j}},X_{i\bar{j}}$. Hence an individual graph appearing in $\mathcal{L}_{n}^{r}$ can be written as
+We find there are new operators $P_{ij},P_{\bar{i}\bar{j}},X_{i\bar{j}}$, so an individual graph appearing in $\mathcal{L}_{n}^{r}$ can be written as
 $\mathsf{F}=\prod_{i=1}^{p_{1}}c_{a_{i}a^{\prime}_{i}}\prod_{i=1}^{p_{2}}c_{\bar{b}_{i}\bar{b}^{\prime}_{i}}\prod_{i=1}^{p_{3}}c_{r_{i}\bar{s}_{i}}c\left[\sigma\right]\ .$ (147)
-As a result, when $n$ increases, the introduction of new operators in BGOE and BGSE further complicates the graph structures, leading to a greater number of individual graphs that become increasingly challenging to systematically organize them into appropriate categories.
-While this classification seems to be complicated, it is, at its core, a well-structured problem.
-As discussed in the introduction, the method presented in this paper is fundamentally equivalent to the approach outlined in Guo:2024zmr .
-Moreover, the inherent symmetry of $\mathcal{L}_{n}$ in the BGUE, BGOE, and BGSE cases simplifies its diagonalization, which is closely related to a well-studied mathematical problem: the decomposition of the direct product of fundamental representations.
-Given this, we expect that the techniques developed in this work can also be successfully extended to models without obvious symmetry, further demonstrating the power and versatility of the proposed framework.
+So there will be more individual graphs, which can be challenging to organize into appropriate categories.
+While this task may seem intricate, it is fundamentally a straightforward endeavor.
+However, as discussed in the introduction, the method outlined in this paper is essentially equivalent to the approach detailed in Guo:2024zmr .
+The inherent symmetry of $\mathcal{L}_{n}$ for BGUE, BGOE and BGSE simplifies its diagonalization, a well-studied problem in the field of mathematics: the decomposition of the direct product of the fundamental representations.
+We therefore anticipate that the techniques employed in this paper can be successfully applied to models lacking obvious symmetry, thereby highlighting the power and versatility of the proposed method.
 Acknowledgment
-The authors would like to thank Haifeng Tang, Yingyu Yang, and Yanyuan Li for their valuable discussions and comments during the preparation of this work.
-CP, JY and TL are supported by NSFC NO.
-12175237, CP is also supported in part by NSFC NO.
-12447108, NSFC NO.
-12247103, the Fundamental Research Funds for the Central Universities, and funds from the Chinese Academy of Sciences.
+The author would like to thank Haifeng Tang, Cheng Peng, Yingyu Yang, and Yanyuan Li for their valuable discussions and comments during the preparation of this work.
 Appendix A Expression of $\mathcal{U}_{3}(t)$
-As discussed in Sec.4, we write $\mathcal{U}_{3}=\sum_{a=1}^{26}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}$. Here we list the explicit expressions for $f_{a}$. For simplicity, we denote
-$f_{a}(t)=\boldsymbol{v}_{a}\cdot\boldsymbol{\xi}(t)\ ,$ (155)
+As discussed in Sec.4, we write $\mathcal{U}_{3}=\sum_{a=1}^{26}f_{a}(t)e^{-\text{i}\mathsf{E}t}\mathsf{F}_{a}$. Here we list the explicit expressions for $f_{a}$’s. For simplicity, we denote
+$f_{a}(t)=\boldsymbol{v}_{a}\cdot\boldsymbol{\xi}(t)$ (148)
 where $\boldsymbol{\xi}(t)$ are $10$ wavefunctions.
-$\boldsymbol{\xi}(t)=\left\{e^{-Jt},e^{-2Jt},e^{-\frac{2(D+1)Jt}{D}},e^{-\frac{2(D-1)Jt}{D}},e^{-3Jt},e^{-\frac{3(D+1)Jt}{D}},e^{-\frac{3(D-1)Jt}{D}},e^{-\frac{3(D+2)Jt}{D}},e^{-\frac{3(D-2)Jt}{D}},1\right\}\ ,$ (156)
+$\boldsymbol{\xi}(t)=\left\{e^{-Jt},e^{-2Jt},e^{-\frac{2(D+1)Jt}{D}},e^{-\frac{2(D-1)Jt}{D}},e^{-3Jt},e^{-\frac{3(D+1)Jt}{D}},e^{-\frac{3(D-1)Jt}{D}},e^{-\frac{3(D+2)Jt}{D}},e^{-\frac{3(D-2)Jt}{D}},1\right\}$ (149)
 and the $26$ vectors are
-$\boldsymbol{v}_{1}$ =$=\left\{0,0,0,0,\frac{1}{2},\frac{2}{9},\frac{2}{9},\frac{1}{36},\frac{1}{36},0\right\},$
-$\boldsymbol{v}_{2}$ =$=\left\{0,0,0,0,0,\frac{1}{9},-\frac{1}{9},\frac{1}{36},-\frac{1}{36},0\right\},$
-$\boldsymbol{v}_{3}$ =$=\left\{0,0,0,0,-\frac{1}{18},0,0,\frac{1}{36},\frac{1}{36},0\right\},$
-$\boldsymbol{v}_{4}$ =$=\left\{0,0,0,0,-\frac{1}{6},\frac{1}{18},\frac{1}{18},\frac{1}{36},\frac{1}{36},0\right\},$
-$\boldsymbol{v}_{5}$ =$=\left\{0,0,0,0,0,-\frac{1}{18},\frac{1}{18},\frac{1}{36},-\frac{1}{36},0\right\},$
-$\boldsymbol{v}_{6}$ =$=\left\{0,0,0,0,\frac{1}{6},-\frac{1}{9},-\frac{1}{9},\frac{1}{36},\frac{1}{36},0\right\},$
-$\boldsymbol{v}_{7}$ =$=\Bigg{\{}0,\frac{5-D^{2}}{18D-2D^{3}},\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{D^{2}-3D-2}{4(D-4)(D-1)(D+2)},$
+$\boldsymbol{v}_{1}$ $=\left\{0,0,0,0,\frac{1}{2},\frac{2}{9},\frac{2}{9},\frac{1}{36},\frac{1}{36},0\right\},$
+$\boldsymbol{v}_{2}$ $=\left\{0,0,0,0,0,\frac{1}{9},-\frac{1}{9},\frac{1}{36},-\frac{1}{36},0\right\},$
+$\boldsymbol{v}_{3}$ $=\left\{0,0,0,0,-\frac{1}{18},0,0,\frac{1}{36},\frac{1}{36},0\right\},$
+$\boldsymbol{v}_{4}$ $=\left\{0,0,0,0,-\frac{1}{6},\frac{1}{18},\frac{1}{18},\frac{1}{36},\frac{1}{36},0\right\},$
+$\boldsymbol{v}_{5}$ $=\left\{0,0,0,0,0,-\frac{1}{18},\frac{1}{18},\frac{1}{36},-\frac{1}{36},0\right\},$
+$\boldsymbol{v}_{6}$ $=\left\{0,0,0,0,\frac{1}{6},-\frac{1}{9},-\frac{1}{9},\frac{1}{36},\frac{1}{36},0\right\},$
+$\boldsymbol{v}_{7}$ $=\Bigg\{0,\frac{5-D^{2}}{18D-2D^{3}},\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{D^{2}-3D-2}{4(D-4)(D-1)(D+2)},$
 $\hskip 15.0pt-\frac{20-9D^{2}}{72D-18D^{3}},-\frac{2(D+2)}{9(D+1)(D+3)},-\frac{2(D-2)}{9\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg\},$
-$\boldsymbol{v}_{8}$ =$=\Bigg{\{}0,0,\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{-D^{2}+3D+2}{4\left(D^{3}-3D^{2}-6D+8\right)},\frac{4}{36-9D^{2}},$
+$\boldsymbol{v}_{8}$ $=\Bigg\{0,0,\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{-D^{2}+3D+2}{4\left(D^{3}-3D^{2}-6D+8\right)},\frac{4}{36-9D^{2}},$
 $\hskip 15.0pt-\frac{1}{9D+9},\frac{1}{9(D-1)},-\frac{1}{36(D+4)},\frac{1}{36(D-4)},0\Bigg\},$
-$\boldsymbol{v}_{9}$ =$=\Bigg{\{}0,\frac{1}{18-2D^{2}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},-\frac{D-2}{4(D-4)(D-1)(D+2)},$
+$\boldsymbol{v}_{9}$ $=\Bigg\{0,\frac{1}{18-2D^{2}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},-\frac{D-2}{4(D-4)(D-1)(D+2)},$
 $\hskip 15.0pt\frac{2}{9\left(D^{2}-4\right)},-\frac{2D+3}{18(D+1)(D+3)},\frac{2D-3}{18(D-3)(D-1)},-\frac{1}{36(D+4)},\frac{1}{36(D-4)},0\Bigg\},$
-$\boldsymbol{v}_{10}$ =$=\Bigg{\{}0,-\frac{D^{2}-5}{2D\left(D^{2}-9\right)},\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{D^{2}-3D-2}{4(D-4)(D-1)(D+2)},\frac{20-D^{2}}{72D-18D^{3}},$
+$\boldsymbol{v}_{10}$ $=\Bigg\{0,-\frac{D^{2}-5}{2D\left(D^{2}-9\right)},\frac{D^{2}+3D-2}{4(D-2)(D+1)(D+4)},\frac{D^{2}-3D-2}{4(D-4)(D-1)(D+2)},\frac{20-D^{2}}{72D-18D^{3}},$
 $\hskip 15.0pt-\frac{2}{9(D+1)(D+3)},\frac{2}{9\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg\},$
-$\boldsymbol{v}_{11}$ =$=\Bigg{\{}0,\frac{1}{18D-2D^{3}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D-2}{4(D-4)(D-1)(D+2)},-\frac{D^{2}+4}{72D-18D^{3}},$
+$\boldsymbol{v}_{11}$ $=\Bigg\{0,\frac{1}{18D-2D^{3}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D-2}{4(D-4)(D-1)(D+2)},-\frac{D^{2}+4}{72D-18D^{3}},$
 $\hskip 15.0pt-\frac{1}{18(D+1)(D+3)},\frac{1}{18\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg\},$
-$\boldsymbol{v}_{12}$ =$=\Bigg{\{}0,\frac{1}{D\left(D^{2}-9\right)},\frac{1}{2(D-2)(D+1)(D+4)},\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{8-D^{2}}{72D-18D^{3}},$
+$\boldsymbol{v}_{12}$ $=\Bigg\{0,\frac{1}{D\left(D^{2}-9\right)},\frac{1}{2(D-2)(D+1)(D+4)},\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{8-D^{2}}{72D-18D^{3}},$
 $\hskip 15.0pt\frac{1}{9\left(D^{2}+4D+3\right)},-\frac{1}{9\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg\},$
-$\boldsymbol{v}_{13}$ =$=\Bigg{\{}0,-\frac{1}{18D-2D^{3}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D-2}{4(D-4)(D-1)(D+2)},\frac{4-3D^{2}}{72D-18D^{3}},$
+$\boldsymbol{v}_{13}$ $=\Bigg\{0,-\frac{1}{18D-2D^{3}},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D-2}{4(D-4)(D-1)(D+2)},\frac{4-3D^{2}}{72D-18D^{3}},$
 $\hskip 15.0pt-\frac{D+2}{18(D+1)(D+3)},-\frac{D-2}{18\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg\},$
-$\boldsymbol{v}_{14}$ =$=\Bigg{\{}0,\frac{1}{2\left(D^{2}-9\right)},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},-\frac{D-2}{4(D-4)(D-1)(D+2)},\frac{2}{9\left(D^{2}-4\right)},$
+$\boldsymbol{v}_{14}$ $=\Bigg\{0,\frac{1}{2\left(D^{2}-9\right)},-\frac{D+2}{4\left(D^{3}+3D^{2}-6D-8\right)},-\frac{D-2}{4(D-4)(D-1)(D+2)},\frac{2}{9\left(D^{2}-4\right)},$
 $\hskip 15.0pt\frac{D}{18\left(D^{2}+4D+3\right)},-\frac{D}{18\left(D^{2}-4D+3\right)},-\frac{1}{36(D+4)},\frac{1}{36(D-4)},0\Bigg\},$
-$\boldsymbol{v}_{15}$ =$=\Bigg{\{}0,0,\frac{1}{2(D-2)(D+1)(D+4)},-\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{1}{36-9D^{2}},\frac{1}{18D+18},$
+$\boldsymbol{v}_{15}$ $=\Bigg\{0,0,\frac{1}{2(D-2)(D+1)(D+4)},-\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{1}{36-9D^{2}},\frac{1}{18D+18},$
 $\hskip 15.0pt\frac{1}{18-18D},-\frac{1}{36(D+4)},\frac{1}{36(D-4)},0\Bigg\},$
-$\boldsymbol{v}_{16}$ =$=\Bigg{\{}0,\frac{1}{9D-D^{3}},\frac{1}{2(D-2)(D+1)(D+4)},\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},-\frac{8-3D^{2}}{72D-18D^{3}},$
+$\boldsymbol{v}_{16}$ $=\Bigg\{0,\frac{1}{9D-D^{3}},\frac{1}{2(D-2)(D+1)(D+4)},\frac{1}{2\left(D^{3}-3D^{2}-6D+8\right)},-\frac{8-3D^{2}}{72D-18D^{3}},$
 $\hskip 15.0pt\frac{D+2}{9\left(D^{2}+4D+3\right)},\frac{D-2}{9(D-3)(D-1)},-\frac{1}{36(D+4)},\frac{1}{144-36D},0\Bigg\},$
-$\boldsymbol{v}_{17}$ =$=\Bigg{\{}\frac{D^{4}-8D^{2}+6}{D^{6}-13D^{4}+36D^{2}},\frac{3-D^{2}}{D^{2}\left(D^{2}-9\right)},-\frac{D(D+3)}{2\left(D^{4}+5D^{3}-20D-16\right)},-\frac{(D-3)D}{2\left(D^{4}-5D^{3}+20D-16\right)},$
+$\boldsymbol{v}_{17}$ $=\Bigg\{\frac{D^{4}-8D^{2}+6}{D^{6}-13D^{4}+36D^{2}},\frac{3-D^{2}}{D^{2}\left(D^{2}-9\right)},-\frac{D(D+3)}{2\left(D^{4}+5D^{3}-20D-16\right)},-\frac{(D-3)D}{2\left(D^{4}-5D^{3}+20D-16\right)},$
 $\hskip 15.0pt\frac{6-4D^{2}}{36D^{2}-9D^{4}},\frac{2}{9(D+1)(D+3)},\frac{2}{9\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},\frac{1}{18\left(D^{2}-7D+12\right)},0\Bigg\},$
-$\boldsymbol{v}_{18}$ =$=\Bigg{\{}\frac{1}{9D-D^{3}},-\frac{1}{18D-2D^{3}},-\frac{D}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D}{4\left(D^{3}-3D^{2}-6D+8\right)},-\frac{2}{36D-9D^{3}},$
+$\boldsymbol{v}_{18}$ $=\Bigg\{\frac{1}{9D-D^{3}},-\frac{1}{18D-2D^{3}},-\frac{D}{4\left(D^{3}+3D^{2}-6D-8\right)},\frac{D}{4\left(D^{3}-3D^{2}-6D+8\right)},-\frac{2}{36D-9D^{3}},$
 $\hskip 15.0pt\frac{1}{18\left(D^{2}+4D+3\right)},-\frac{1}{18\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},-\frac{1}{18(D-4)(D-3)},0\Bigg\},$
-$\boldsymbol{v}_{19}$ =$=\Bigg{\{}\frac{1}{9D-D^{3}},\frac{2}{D\left(D^{2}-9\right)},\frac{1}{(D-2)(D+1)(D+4)},\frac{1}{D^{3}-3D^{2}-6D+8},\frac{4}{36D-9D^{3}},$
+$\boldsymbol{v}_{19}$ $=\Bigg\{\frac{1}{9D-D^{3}},\frac{2}{D\left(D^{2}-9\right)},\frac{1}{(D-2)(D+1)(D+4)},\frac{1}{D^{3}-3D^{2}-6D+8},\frac{4}{36D-9D^{3}},$
 $\hskip 15.0pt\frac{2}{9(D+1)(D+3)},-\frac{2}{9\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},-\frac{1}{18(D-4)(D-3)},0\Bigg\},$
-$\boldsymbol{v}_{20}$ =$=\Bigg{\{}\frac{2D^{2}-3}{D^{2}\left(D^{4}-13D^{2}+36\right)},\frac{D^{2}-3}{2D^{2}\left(D^{2}-9\right)},\frac{D^{2}+3D+4}{-4D^{4}-20D^{3}+80D+64},$
+$\boldsymbol{v}_{20}$ $=\Bigg\{\frac{2D^{2}-3}{D^{2}\left(D^{4}-13D^{2}+36\right)},\frac{D^{2}-3}{2D^{2}\left(D^{2}-9\right)},\frac{D^{2}+3D+4}{-4D^{4}-20D^{3}+80D+64},$
 $\hskip 15.0pt\frac{D^{2}-3D+4}{-4D^{4}+20D^{3}-80D+64},\frac{D^{2}+3}{9D^{2}\left(D^{2}-4\right)},-\frac{1}{9(D+1)(D+3)},-\frac{1}{9\left(D^{2}-4D+3\right)},$
 $\hskip 15.0pt\frac{1}{18(D+3)(D+4)},\frac{1}{18\left(D^{2}-7D+12\right)},0\Bigg\},$
-$\boldsymbol{v}_{21}$ =$=\Bigg{\{}\frac{D^{2}+6}{D^{6}-13D^{4}+36D^{2}},\frac{3}{D^{2}\left(D^{2}-9\right)},\frac{D}{2\left(D^{4}+5D^{3}-20D-16\right)},\frac{D}{-2D^{4}+10D^{3}-40D+32},$
+$\boldsymbol{v}_{21}$ $=\Bigg\{\frac{D^{2}+6}{D^{6}-13D^{4}+36D^{2}},\frac{3}{D^{2}\left(D^{2}-9\right)},\frac{D}{2\left(D^{4}+5D^{3}-20D-16\right)},\frac{D}{-2D^{4}+10D^{3}-40D+32},$
 $\hskip 15.0pt\frac{D^{2}-6}{9D^{2}\left(D^{2}-4\right)},-\frac{1}{9(D+1)(D+3)},-\frac{1}{9\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},\frac{1}{18\left(D^{2}-7D+12\right)},0\Bigg\},$
-$\boldsymbol{v}_{22}$ =$=\Bigg{\{}\frac{2D^{2}-3}{D^{2}\left(D^{4}-13D^{2}+36\right)},\frac{3}{18D^{2}-2D^{4}},\frac{3D+4}{4\left(D^{4}+5D^{3}-20D-16\right)},\frac{4-3D}{4\left(D^{4}-5D^{3}+20D-16\right)},$
+$\boldsymbol{v}_{22}$ $=\Bigg\{\frac{2D^{2}-3}{D^{2}\left(D^{4}-13D^{2}+36\right)},\frac{3}{18D^{2}-2D^{4}},\frac{3D+4}{4\left(D^{4}+5D^{3}-20D-16\right)},\frac{4-3D}{4\left(D^{4}-5D^{3}+20D-16\right)},$
 $\hskip 15.0pt\frac{3-2D^{2}}{9D^{2}\left(D^{2}-4\right)},\frac{1}{18\left(D^{2}+4D+3\right)},\frac{1}{18\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},\frac{1}{18\left(D^{2}-7D+12\right)},0\Bigg\},$
-$\boldsymbol{v}_{23}$ =$=\Bigg{\{}-\frac{5}{D^{5}-13D^{3}+36D},\frac{1}{9D-D^{3}},\frac{D}{2\left(D^{4}+5D^{3}-20D-16\right)},\frac{D}{2\left(D^{4}-5D^{3}+20D-16\right)},$
+$\boldsymbol{v}_{23}$ $=\Bigg\{-\frac{5}{D^{5}-13D^{3}+36D},\frac{1}{9D-D^{3}},\frac{D}{2\left(D^{4}+5D^{3}-20D-16\right)},\frac{D}{2\left(D^{4}-5D^{3}+20D-16\right)},$
 $\hskip 15.0pt\frac{1}{36D-9D^{3}},-\frac{1}{9(D+1)(D+3)},\frac{1}{9\left(D^{2}-4D+3\right)},\frac{1}{18(D+3)(D+4)},-\frac{1}{18(D-4)(D-3)},0\Bigg\},$
-$\boldsymbol{v}_{24}$ =$=\Bigg{\{}\frac{3}{9D-D^{3}},0,\frac{3}{2(D-2)(D+1)(D+4)},\frac{3}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{2}{12D-3D^{3}},0,0,$
+$\boldsymbol{v}_{24}$ $=\Bigg\{\frac{3}{9D-D^{3}},0,\frac{3}{2(D-2)(D+1)(D+4)},\frac{3}{2\left(D^{3}-3D^{2}-6D+8\right)},\frac{2}{12D-3D^{3}},0,0,$
 $\hskip 15.0pt-\frac{1}{6(D+2)(D+3)(D+4)},-\frac{1}{6\left(D^{3}-9D^{2}+26D-24\right)},\frac{2-D^{2}}{D^{5}-5D^{3}+4D}\Bigg\},$
-$\boldsymbol{v}_{25}$ =$=\Bigg{\{}\frac{5}{D^{4}-13D^{2}+36},0,\frac{1}{2\left(D^{3}+7D^{2}+14D+8\right)},\frac{1}{-2D^{3}+14D^{2}-28D+16},0,0,0,$
+$\boldsymbol{v}_{25}$ $=\Bigg\{\frac{5}{D^{4}-13D^{2}+36},0,\frac{1}{2\left(D^{3}+7D^{2}+14D+8\right)},\frac{1}{-2D^{3}+14D^{2}-28D+16},0,0,0,$
 $\hskip 15.0pt-\frac{1}{6(D+2)(D+3)(D+4)},\frac{1}{6(D-4)(D-3)(D-2)},\frac{1}{D^{4}-5D^{2}+4}\Bigg\},$
-$\boldsymbol{v}_{26}$ =$=\Bigg{\{}-\frac{15}{D^{5}-13D^{3}+36D},0,-\frac{3}{D^{4}+5D^{3}-20D-16},\frac{3}{D^{4}-5D^{3}+20D-16},-\frac{1}{12D-3D^{3}},$
+$\boldsymbol{v}_{26}$ $=\Bigg\{-\frac{15}{D^{5}-13D^{3}+36D},0,-\frac{3}{D^{4}+5D^{3}-20D-16},\frac{3}{D^{4}-5D^{3}+20D-16},-\frac{1}{12D-3D^{3}},$
 $\hskip 15.0pt0,0,-\frac{1}{6(D+2)(D+3)(D+4)},-\frac{1}{6\left(D^{3}-9D^{2}+26D-24\right)},-\frac{2}{D^{5}-5D^{3}+4D}\Bigg\}\ .$